A Complete Bibliography of the Publications of John R. Rice

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

15 August 2025 Version 1.54

Abstract

This bibliography records publications of John R. Rice.

Title word cross-reference

#4 [Ric84c].

ab^{x} + c [Ric60b, Ric61b]. ADI [LR68a]. a ∏ \frac{x−r_{i}}{x+s_{i}} [dBR63]. Erfc(x) [Ric64d]. Γ(x) [Ric64d]. L_{1} [HR65, Ric64e, Ric64c]. L_{∞} [Ric64d]. O(h^{4}) [HRV86, HVR88]. ∥ [HRP89].

// [HRP88, HR89, HRW^{+}96]. //ELLPACK [HHK^{+}90, HRC^{+}90a, HRC^{+}90b, HHR^{+}91, WCHR92, WHCR95].

1 [LR78c, Ric74a]. 10-14 [MN95]. 12th [MH89]. 13-15 [Cra87]. 13th [BK92, HR92c, VM91]. 14th [Ame94]. 15th [Syd97]. 16th [DO00]. 17-21 [FCC87]. 17th [Ger03, IEE03]. 1965 [Kal65]. 1968 [OR70]. 1969 [Sch69, Tal70]. 1973 [LBCS73]. 1975 [LS76, Tra76]. 1976 [HW76]. 1977 [Han78, Ric77c]. 1978 [ACM79, dBG78]. 1981 [Rei82]. 1982 [Lei83]. 1983 [ESY84]. 1985 [FCC87, MC87]. 1987 [Cra87]. 1988 [HRV90, Wri89]. 1990 [Ano90, SdG90]. 1991 [AvdH91, BK92, DL91, HR92c]. 1993 [Zah94]. 1995 [IEE95b]. 1996 [Ric96d, Sie96]. 1999 [BH02]. 1st [HPP88].

2-Year [Ric78a]. 2/W.G [FCC87]. 21st [ACM66, Ric99b, Ric02]. 24th [ACM69].

32 [MR87d]. 3294OX [Ric75f]. 3rd [Jáj90].

438 [HMR72]. 49 [Ric65c].

5-Point [RM88b]. 525 [Ric78c].

620 [RH84]. 622 [Lev98, RRW84]. 637 [HMR85a]. 638 [HMR85b]. 65th [BDH00, BH02, Ric94d].

77 [Ric78d, Ric78e]. ’78 [Ric78f]. 7th [IEE95b].

8th [Ano90].

’91 [VM91]. ’94 [Ame94, Ano94]. ’95 [IEE95a]. ’98 [ACM98]. 9th [IEE95a].

Abstract [BDH^{+}88, Ric74d, MRV91, Ric92a, Ric92c, Ric00a]. Academic [Ric91a, Ric93d, Ric93a, Ric94b, Ric94e]. Accuracy [LR75b, LR78a]. Accurately [Ric99a]. ACM [ACM66, ACM69, BK93, Cra87, HMR72, Ric87c, Ric90b, Ham85, HM90, RH84]. Adapt [Ric75a, Ric78c]. Adaptation [JDR^{+}95]. Adaptive [ABB^{+}00, ABE^{+}03b, DDH^{+}98, LR75a, LR79, LR80a, MMR95, Ric73b, Ric73c, Ric73g, Ric73f, Ric74b, Ric74c, Ric74i, Ric74h, Ric75a, Ric75e, Ric76d, Ric78c, Ric78b, Ric85a, Ric87a, AAB^{+}02, ABE^{+}03a, Ric73d, Ric75j, Ric76g, Ric76i, dBR79]. ADI [LR68b, dBR63]. Adjoint [RD81]. Advances [RY76, VS79, Zel98, Zel01, dBG78]. Age [HJR95]. Agent [BMR^{+}00, DJR95b, DRJ96, JDR^{+}96, JRD^{+}97, MHC^{+}99, MHC^{+}00b, Ric97f, Ric98, DJR95a, JDR^{+}97, JRD^{+}01, MHC^{+}00a]. Agent-Based [MHC^{+}99, MHC^{+}00b, JRD^{+}01, MHC^{+}00a]. Agents [DHRR99, TBMR99a, DHJ^{+}97, RV98, TBMR99b]. Agents’99 [Ano99]. Agglomerative [JRRH95, JRRH96]. Aided [MN95, Lei83]. air [RV98]. Alberta [LS76]. Algebra [Ric84c]. Algorithm [HMR72, HRM83a, HHRS90, RR96a, RR96b, RJHR97a, Ric73b, Ric73c, Ric74f, Ric74e, Ric74d, Ric75a, Ric75b, Ric76e, Ric84b, RM88a, HHR^{+}88, HHR^{+}89c, MR89a, RR02a, Ric59a, RU68, Ric71b, Ric73d, Ric79a, dBR79, Ric81a, Ham85, HM90, HMR85a, HMR85b, Lev98, Ric78c, RH84, RRW84]. Algorithm/Software [RJHR97a]. Algorithmic [MR88c, Ric75c, Ric76f].

Algorithms [Bor70, CHR94, Ham85, HM90, HLPR74, JGD87, LR75a, LR79, MR86, MR87c, MR87a, MC87, MR90i, Ric61b, Ric73g, Ric73f, Ric74e, Ric74i, Ric74h, RH84, BK92, LR80a, MR87b, RR69, Ric74c, Ric75j, Ric76i, MC87]. allocation [HHR84]. Alternating [KHHR95, LRT65, KHHR96]. America [HJR95]. Amsterdam [SdG90]. Anaheim [Kum91]. Analysis [HHRV99, LHHR94a, LRT65, MR86, MR87c, MR87a, MR89a, MR88e, RR66b, Ric67b, Ric68c, Ric70d, Ric74e, Ric78i, Ric81e, Ric83d, Ric83e, RM88a, RM88b, Ric88b, Ric92e, RVY93, RVY97, dBG78, DL91, LRT64a, MR87b, MR89c, RR66a, Ric67a, Ric68b, Ric71a, Ric76a, Ric79b, Ric79d, Ric83i, Ric83g, Ric93f]. Analytic [Tra76, Tra76]. Announcing [Ric96d]. Answer [Ric96h, Ric97b, Ric00d]. Antipolis [FCC87]. applicability [Ric63a]. Application [ACR02, BCRR83, MR90d, RLR00, dBR63, dBR82, Ric71e]. Application-Specific [BCRR83]. Applications [HHRS90, HRW^{+}96, HRW^{+}98, Hwa84, Ric67b, Ric78b, WHR^{+}94b, AAB^{+}02, HHR^{+}88, HHR^{+}89c, HRW^{+}00, LS76, MTH^{+}03, CCC^{+}94, S^{+}83]. Applications/Architecture [HHRS90, HHR^{+}88, HHR^{+}89c]. Applied [Ame94, Ano90, KR68, OR70, Syd97, VM91, AvdH91, BK92, DO00]. Approach [ABE^{+}03b, JRRH95, ABE^{+}03a, JRRH96]. Approaches [JWR^{+}95, RJHR97b]. Approximate [Ric68c]. approximating [Ric65a]. Approximation [Gar65, HR67, LR75b, Mei67, MR92g, MR93d, MR93e, MR94e, Ric60b, Ric62a, Ric63b, Ric64a, Ric67a, Ric68a, Ric69a, Ric74b, Ric74f, Ric75i, Ric75f, Ric75l, Ric78b, Ric91e, Ste77, dBR63, dBR68b, Han78, HR65, LS76, LR78a, MC87, Ric61b, Ric61a, Ric62b, Ric63d, Ric64b, Ric64c, Ric65d, Ric67d, Ric69c, Ric69e, Ric70c, Ric73e, Ric76d, Ric76g, Ric76h, Ric77b, Ric78h, Ric92d, dBR68a, dBR79, Gar65, LBCS73, Mei64, Ric65c, Tal70, Zah94]. Approximations [HCL^{+}68a, HCL^{+}68b, Ric59b, Ric64e, Ric67c, Ric59a, Ric60a, Ric60c, Ric60d, Ric61c, Ric61d, Ric64f, Ric71b, Sch69]. April [Ger03, HW76, HRV92, IEE93, IEE95a, IEE03, Rob73, Tra76]. Architecture [HHRS90, HR92b, MR92b, Ric97f, Ric98, BT01, HHR^{+}88, HHR^{+}89c, MTH^{+}03]. Architectures [CHH^{+}90, HHR^{+}89a, CHH^{+}91, HHR^{+}89b]. Area [BCRR83, Ric80b, Ric81c]. areas [Ric76b]. Arithmetic [HRR80]. Army [Ano90, Sch69]. Array [Ric81b]. arrays [Ric64d]. Artificial [Ano95, IEE95b, HR92c]. Aspect [Ric85b]. Aspects [Ric87d, Ric88d, Ric96b, Ric96c, Wri89, Ric88c]. Assessment [BBC^{+}96, HD80]. Assignment [MR92h, MR89d]. Association [MH89]. Asynchronous [MR88b, MR94b, RM88a, Wri89, MR89a, MR90a]. ATHENA [HHK^{+}90, HHR^{+}91]. Athens [HPP88]. Atlanta [AGH^{+}95, Ame94]. August [ESY84, LS76, Rei82, Ros74, Sie96, Syd97, Wri89]. Austin [LBCS73]. Authentication [ZJ04]. Authors [Ric76j]. Automated [JWR^{+}96, RTV98a, RTV99, Ard80]. Automatic [CHH^{+}89a, CHH^{+}89b, RR96a, Ric65a, Ric69f, Ric68b, Ric69b]. available [Ric91a, Ric93a]. Averaging [MRCH^{+}92, MRCH^{+}93].

B [Ric75o]. B74 [Ric74a]. B74-1 [Ric74a]. Balance [MR90b, MR91b]. Balanced [CHH^{+}89a, CHH^{+}89b]. balancing [MMR95]. Bandwidth [DCRS81]. Barbara [IEE95a]. Base [HHK^{+}90, HVC^{+}98, HVC^{+}99, HHR^{+}91]. Based [CR92b, DJR95b, DRJ96, DHR95, HRW^{+}96, HRW^{+}98, HCB^{+}01, JRD^{+}97, LHHR94a, MHC^{+}99, MHC^{+}00b, MR92h, Ric97f, WHR^{+}94b, WHR^{+}95, WHR^{+}96c, AAB^{+}02, BMR^{+}00, DJR95a, HRMW97, HRW^{+}00, JRD^{+}01, KDM^{+}92, MHC^{+}00a, MWHR00, MC87, MR89d, Ric60d, Ric92c, Ric92b, CCC^{+}94]. BASIC [LR75f, RR73a]. be [Ard80]. Beach [IEE93]. bearings [Ric63c, Ric65b]. Behandlung [Mei64, Ric65c]. behavior [RR82]. Benchmark [MR92i]. Benchmarking [HRV89a]. Berlin [Ric65c, Syd97]. Best [Ric59b, Ric61c, Ric59a, Ric60a, Ric60c, Ric64f]. Bethlehem [VS79]. Between [Rei82]. Beyond [Ano01]. Bi [DLRH81, HRM83a]. Bi-Cubic [HRM83a]. Bi-Cubics [DLRH81]. Bicubic [DR81, DR84a, HMR85a, HMR85b, DR84b, DR86]. bicubics [DHLR84]. Biographical [Ricxx, Ano00]. Biometric [ZJ04]. Birthday [BH02, BDH00, Ric94d]. Blading [FZHR99]. BLAS [Ric84c]. Blocking [MR92a, MR93a]. Book [Ric65c, Ric67b, Ric74g]. Boulder [Rei82]. Boundary [DRR86, HCR86, KHHR95, RR82, DRR88, HCR88, KHHR96]. Breakthroughs [Ric75h]. Broadening [KWRW93]. Building [Ric84d, WHR^{+}94b, WHR^{+}96a, WHR90, CCC^{+}94, WHR^{+}00]. Business [Ano01].

CA [ACM79, BS84, Wri89, IEE95a, Kum91]. Calgary [LS76]. California [IEE93]. Can [Ric99a, Ard80]. Canada [LS76, MN95, BT01]. Capabilities [Ric83b]. Capacitance [CHR88b, CHR88a]. Carnegie [Tra76]. Case [RR96b]. Celebration [BH02]. Center [ADHR73, BK93, Ric77c, Sch69, dBG78]. Centers [BBC^{+}96]. Century [Ric99b, Ric02]. Challenges [Ric75h, Sie96, Ric97a]. Characteristics [CR92b, JGD87]. Characterization [MR90e, MR94a, Ric59b, Ric67c, MR90f, Ric60a, Ric64f, Ric70b]. Chebyshev [Ric60b, Ric61b, Ric62a, Ric67c, dBR63]. China [ZJ04]. Chinook [NR02]. City [Kal65]. Clustering [JRRH95, JRRH96]. cm [Ric65c]. Collaborating [DHJ^{+}97, MR91d, MR92c, MR93b, MR94c, RV98, MR94d, MR95]. Collaborative [RJHR97a, RJHR97b]. Collected [Ham85, HM90, RH84]. collection [OR70]. College [AGH^{+}95, Jáj90, MC87, VM91]. Collocation [CHR88b, DR81, DLRH81, DR84a, HHRV91b, HHRV91a, HHRV99, HRM83a, HRM83b, HMR85a, HMR85b, HMR85c, HRV86, HCR86, HVR88, HRV88b, LHHR92, LHHR94a, LHHR94b, LHHR95, MR88e, CHR88a, DR84b, DHLR84, DR86, HHRV93, HCR88, LHHR94c, MR89c]. Collocation-Capacitance [CHR88b, CHR88a]. Colorado [Rei82].

Combustion [ZFHR99]. Combustor [FHRZ99a, Ric99a]. Come [Ric03]. Communication [MRV90b, MR90b, MR91a, MR92a, MR91b, MR93a]. Community [Ric97c]. Commutative [dBR64]. compact [Ric62b]. Comparison [CPR69b, FR93, CPR69a]. Compiler [Ric80c, Ric83c]. Complex [JDR^{+}96, JDR^{+}97, VHR99]. Complexity [MR92a, Ric74a, Ric75i, MR93a, Ric73e, Ric76h, Tra76, Ric74g]. Composite [MR97]. Composition [Ric89]. Compressor [FHRZ99a]. Compressor-Combustor [FHRZ99a]. Computation [Ame94, Cra87, DCRS81, GOP^{+}79, GOP^{+}80, Jáj90, MH89, MR87a, MR90d, MSV92, RJW^{+}95b, Rei82, Ric64e, Ric72, Ric75h, Ric75o, Ric75m, Ric79c, Ric81d, Ric84f, Ric84g, Syd97, VM91, WHR94a, Wri89, DO00, Ric71e, Ric76b, RGO^{+}79, Ric81g, Ric82a, Ric82b, Ric96f, Zah94, MH89]. Computational [ABB^{+}00, APRS99, BH02, CGH^{+}97, DDH^{+}98, GHR92, GHR94, HHR^{+}93, HR78a, HGBR97a, HGBR97b, RR96b, Ric75i, Ric93d, Ric94b, Ric94a, Ric94c, Ric95b, Ric95c, Ric96b, Ric96c, Ric96d, Ric97c, Ric99b, SCK^{+}96, CGH^{+}00, HHR^{+}94, HR80a, HRR^{+}98, HGRB00, HR00b, Ric73e, Ric76h, Ric91a, Ric92a, Ric93a, Tra76, AvdH91, BH02, BK92]. Computations [AGH^{+}95, CHH^{+}89a, CHH^{+}90, CR92b, HHR86a, MR87c, MR89b, MRV90b, MR94b, Ric74a, Ric81h, AR01a, AR01b, CHH^{+}89b, CHH^{+}91, HHR84, HHR87, MR91a, Ric74g]. Computer [AFI73, ADHR73, Bor70, BK93, GHR94, HCL^{+}68a, HCL^{+}68b, HW76, HRR80, KWRW93, LHRL93, MN95, RR83, RR93, Ric74a, Ric74g, Ric91c, Ric93c, Ric93e, Ric96e, Ric97c, RR02b, RR04, RR90b, SW86, Tra76, VS79, Ard80, RRH00, RR69, RD94, Lei83, RR90a, RR94a]. Computer-Aided [MN95]. Computers [LR75f, LR77, LR78c, RY76, Zel98, Zel01, RR69, Bor70]. Computing [Ano90, Ano94, BDH^{+}87, BDH^{+}88, BDH^{+}89, DFF^{+}03, DJR95b, DHRR99, HHR86b, HRV89b, HRV90, HR92c, HRV94, HCB^{+}01, IFI95, JDR^{+}95, JDR^{+}96, LHRL93, MR88d, MWHR97a, MWHR97b, MR92d, Nas90, RJHR97a, Ric75p, RVH90, RVH93, Ric95b, Ric95c, Ric95a, Ric96b, Ric96c, RR90b, SKR^{+}93, WCHR92, WJH^{+}95, WHR^{+}95, WHR^{+}96c, WHR^{+}96b, BCG^{+}00, CHCM^{+}94, DJR95a, FCC87, HRV92, HR92a, HJR^{+}97b, JDR^{+}97, KX94, MWHR00, RJW^{+}95a, RJHR97b, RR73a, RR73b, Ric76c, Ric90a, Rod89, RR94b, S^{+}83, WJH^{+}97]. Concentration [Ric68c]. Concrete [Ric74e]. Condition [Ric66b]. Conditioning [Ric65e]. Conditions [DR94, DRR86, Ric94f, DRR88]. Conducted [dBG78, LS76, LBCS73, Ric77c, Sch69]. Conference [ACM79, ACM88, AFI73, AGH^{+}95, Ano90, BT01, BK93, Cra87, ESY84, FCC87, Fos79, GH92, HPP88, HRV90, HRV92, HRV94, IEE95b, IEE96, IEE97, KX94, MC87, Rei82, RVH90, RVH93, SKR^{+}93, SdG90, SW86, Ver94, Wri89, ZJ04, DL91, LS76, Zah94, ACM66, ACM69]. Conferences [Ric96d, Ric96e]. Congress [AvdH91, Ame94, BK92, DO00, HR92c, MH89, Ros74, Syd97, VM91, Kal65]. conjunction [Ger03]. cons [Ric91b]. Constitutive [HR73]. constraints [Ric63b]. Construction [MR93d, MR93e, Ric68b]. contact [ABCR05]. Conte [RD94]. Contributor [Ric78d]. Control [MRV90b, AAB^{+}02, MR91a]. Convention [BK93]. Convergence [HRV86, HVR88, LR68b, LR68a, MR88e, Ric69d, Ric73f, Ric74h, RM88b, RVY93, MR89c, Ric59a, Ric69e, Ric70b, Ric71c, Ric74c, dBR79]. Converges [Ric97c]. convex [Ric63b]. Cooperative [DJR95b, Ric75d, DJR95a]. Cornell [Ano90]. Correctness [Ric73g, Ric74i, Ric75j, Ric76i]. Correspondence [KWRW93, SWM^{+}89]. COSERS [Ard80]. cost [RR99]. Cracks [Ric68c]. Creation [Ric92c, Ric92b]. Criteria [Ric60c]. CS [Ric91b, Ric96d]. CS&E [Ric02]. CSE [Ric96e, Ric97c]. Cubic [HHRV91a, HRM83a, HRV86, HVR88, HRV88b, LHHR94a, dBR68b, HHRV93, dBR68a]. Cubics [DLRH81]. Current [IFI95]. Curvature [Ric67d]. Curve [Ric75a, Ric78c, HR67, Ric70a].

D [RD94]. DAGS’94 [Mak94]. Dallas [SW86]. Data [CR68, CHR94, HVC^{+}98, HCR^{+}99, HVC^{+}99, MC87, MR90h, Ric75o, Ric88b, CR78, DL91, HVC^{+}00, HCR^{+}00, MTH^{+}03, Ric65a, Ric68a, Ric76b]. database [HCR^{+}00]. Databases [HCR^{+}99]. DC [IEE96, KR68]. December [HRV90, Zah94]. decomposer [CHH^{+}91, CHH^{+}90]. Decomposition [GKM^{+}91, MR88c, MR92g, MR94e, RVY93, KX94, RVY97]. deformation [RR80]. Degree [Ric69d, Ric61d, Ric69e, Ric70b, Ric71c]. Density [MRCH^{+}92, MRCH^{+}93]. Department [Ric91c, Tra76, RR90a, Ric93c, Ric93e, RR94a]. Departments [Ric94e]. Derivatives [CR68, CR78]. Design [ABB^{+}00, DDH^{+}98, FHRZ99a, HR78a, Hwa84, MTH^{+}03, Ric74e, Ric75g, Ric86b, TBMR99a, WJH^{+}95, WHR94d, ZFHR99, HR80a, HCT^{+}02b, HCT^{+}02a, Lei83, Ric92c, Ric92b, TBMR99b, WJH^{+}97, WHR94c, FHRZ99b]. Development [ACM79, HLPR75b, HLPR75a, HR92a]. Difference [LR75b, LR75c, LRT64b, LRT64a, LR78a, LR80b]. differences [Ric63a]. Differencing [RD81, DR85]. Differential [BHR78, BRH79, CHR88b, DHR95, GKM^{+}91, HHRV91a, HHRV99, HLPR74, HLPR75b, HLPR75c, HR78a, HR78b, HR80b, HRM83b, HMR85c, HRV86, HVR88, HRP88, LHHR94b, LR75b, LR75e, LR75c, MR90i, MR92h, Ric75c, Ric75d, RTV97, Ric98, RTV98b, RTV00b, VS79, WHCR95, AvdH91, HHRV93, HLPR75a, HLR78, HR80a, HR82, HRP89, HRP90, HR00a, LHHR94c, LR78a, LR78b, LR80b, MR93c, Ric76f, Ric77a, RHD81, Ric87e, RVY97, RTV00a, RTV02]. Dimensional [AR77, HR75, LR91, Ric81a, Ric81d, Ric84g, RR02a, Ric84b]. dimensions [HHRV93]. Direct [LRT64b]. Direction [LRT65]. Directions [GHR92, IFI95, Ric88e, RGO^{+}79, Ric92a, Weg79]. disciplinary [MHC^{+}00a]. Discovery [HCR^{+}99, RR96b, VHR98, VHR00]. Discretization [LHHR94a]. Distributed [DJR95b, Ger03, HHR^{+}89a, HHR86b, IEE03, MRW^{+}90, MRV90a, MR90c, MR91c, MRCH^{+}92, MRCH^{+}93, MR88f, MR91f, MR91h, MR92i, MR92f, DJR95a, HHR^{+}89b, HHR84, MR88a, MRV91, MRV93, MMR95, MR93c].

DM [Ric65c]. Documentation [BRW78]. Doer [GHR94]. Domain [BCR87, CHH^{+}91, GKM^{+}91, KX94, MR86, MR87b, MR88c, MR92g, MR94e, Ric81a, RVY93, Ric84b, RVY97, CHH^{+}90, KX94]. Domains [HR75, HRM83a, HMR85a, HMR85b, LR91, MR88e, RR86, Ric81d, Ric84g, RM88b, MR89c, RHD81]. driven [MTH^{+}03]. Driving [Ric96b, Ric96c]. Dublin [AvdH91, BK92, HR92c, VM91]. Durham [Han78]. Dynamic [MR90b, MR91b]. Dynamics [FHRZ99a, ZFHR99].

E-Business [Ano01]. E/T [MR90d]. E2 [Ric78c]. Editor [RD68, Ric94c]. Editorial [BDH00]. Editors [HGBR97a]. eds [Ric74g]. Education [Ric75p, Ric76c]. Educational [Ric73b, Ric73c, Ric73d]. Effect [Ric75q]. Effectiveness [Ric80d, Ric80e]. Effects [MR88d, MR88c, MR90b, MR90c, MR91c, Ric80c, MR91b, RR82, Ric83c]. Effort [Ric75d, Ric78i, Ric76a, Ric79b, Ric79d]. Eighth [Ger03]. Elastic [HR73]. Electron [MRCH^{+}92, MRCH^{+}93]. Electronic [WHR94d, WHR94c]. Element [Ric85b]. Elimination [MR88f, MR91f, MR92f, MR93c]. Elliptic [CHR88b, DRR86, HHRV91b, HHRV91a, HHRV99, HLPR74, HLPR75b, HLPR75c, HR75, HR78a, HR80b, HRM83b, HMR85c, HRV86, HVR88, HRV88b, HRV89a, KHHR95, LHHR94a, LHHR94b, LR75b, Ric75d, Ric75g, Ric80c, RB81, RD81, Ric83h, RB85, RVY93, RTV97, RTV98b, RTV00b, CHR88a, DRR88, HHRV93, HLPR75a, HLR78, HR80a, HR82, KHHR96, LHHR94c, LR78a, LR78b, Ric77a, RHD81, Ric83c, Ric84d, Ric84h, RVY97, RTV00a, RTV02, BS84]. ELLPACK [BRW78, BBRW78, BDR86, HRP88, HRP89, HR89, HHK^{+}90, HRC^{+}90a, HRC^{+}90b, HRP90, HHR^{+}91, HR92a, HRW^{+}96, MWHR97a, Ric75d, Ric76a, Ric77a, Ric78d, Ric78f, Ric78e, Ric78g, Ric78a, Ric78i, Ric79b, Ric79d, Ric80a, RB81, Ric83b, Ric84d, Ric84e, RB85, RDHR86, Ric87b, Ric88e, WCHR92, WHCR95, WHR^{+}96b, WHR90]. emphasis [Sch69]. Enabling [HGRB00]. Encyclopedia [RR83, RR93, RRH00]. End [ABB^{+}00, DDH^{+}98]. End-to-End [ABB^{+}00, DDH^{+}98]. Engine [FHRZ99a, HCT^{+}02b, HCT^{+}02a]. Engineering [Ano95, LHRL93, MN95, RR83, RR93, Ric93d, Ric94b, Ric94c, Ric94g, Ric96d, Ric97c, Ard80, KX94, Ric91a, Ric93a, Ric94a, SCK^{+}96, Zel98]. Enhanced [ACK^{+}02a]. Entire [Ric69d, Ric70b, Ric71c]. Environment [ACM79, CHR94, DCRS81, DJR95b, DJR^{+}96, DHR95, FRC^{+}00, HRC^{+}90a, HRW^{+}96, HRW^{+}98, MR92b, Ric84e, WHR94a, WHR94d, DJR95a, FHR^{+}00, HRC^{+}90b, HR92a, HVC^{+}00, HRW^{+}00, HR00a, HCT^{+}02b, HCT^{+}02a, MTH^{+}03, Ric87b, WHR94c]. Environments [GHR92, GHR94, GHR95a, GHR95b, Ger03, HRJ^{+}95, HGBR97a, HGBR97b, HR00c, JWR^{+}96, JRD^{+}97, MHC^{+}99, MHC^{+}00b, RJHR97a, Ric89, Ric95a, Ric96a, RB96, Ric96g, WHR^{+}94b, CHCM^{+}94, FCC87, FDA^{+}03, GH92, HJR95, HRMW97, HJR^{+}97a, HRR^{+}98, HGRB00, HR00b, MHC^{+}00a, Ric92a, CCC^{+}94, Ric00c, RB00]. EPPOD [WHR94c, WHR94d]. Equation [LR75c, Ric69g, Ric98, HR80a, Ric71e]. Equations [BHR78, BRH79, CHR88b, DR81, DR84a, DHR95, GKM^{+}91, HHRV91a, HHRV99, HLPR74, HLPR75b, HLPR75c, HR78a, HR78b, HR80b, HRM83b, HMR85c, HRV86, HVR88, HRP88, LHHR92, LHHR94b, LHHR95, LR75b, LR75e, MR90i, MR92h, Ric69f, Ric75c, Ric75d, Ric80d, Ric80e, Ric81e, RTV97, RTV98b, VS79, WHCR95, AvdH91, DR84b, DR86, HHRV93, HLPR75a, HLR78, HR82, HRP89, HRP90, HR00a, LHHR94c, LRT64b, LRT64a, LR78a, LR78b, LR80b, MR93c, Ric60e, Ric69b, Ric76f, Ric77a, RHD81, Ric83g, Ric87e, RVY97, RTV00a, RTV02]. equioscillation [Ric60c]. Equispaced [CR68, CR78]. Errata [Rob73]. Error [LR68b, LR68a]. essay [Ric96h, Ric97b, Ric00d]. estimating [CR78]. Estimation [CR68, RW64, RTV98a, RTV99, Ric67a]. Evaluate [Ric94e]. Evaluating [HR78b, HHRS90, HHR^{+}88, HHR^{+}89c]. Evaluation [BHR78, BRH79, BDR86, DMNR68, HLPR75b, HLPR75c, HR78a, HLR78, HHR86b, HHR86a, Ric76j, Ric83b, Ric88e, Ric90c, Ric91f, Fos79, HLPR75a, HR80a, HHR87, HRV88a, VHR00]. Evolving [Ric84e, Ric87b]. Excellence [ACD^{+}86]. Execution [Ric75q]. existence [Ric60c, Ric64f]. Expansion [Ric83b]. Experiences [WCHR92]. Experiment [Ric91d, HR80a]. Experimental [HR78a, KR68, MR88e, RM88b, MR89c]. Experiments [MRW91, Ric66a]. Expert [HRP88, HRV89b, HRV90, HR92c, HRV92, HRV94, MR92d, RVH90, RVH93, HRP89, HRP90]. Exponentials [Ric62a, Ric64e]. Exponet [FR93]. Exposition [AFI73]. extended [MRV91, Ric92a, Ric92c]. Extensions [Ric84f, RU68]. External [ACR04]. Extremal [dBR82].

Facilities [Ric81b, Ric81g, Ric82a]. Factorization [MR90g]. Fall [OR70, SW86]. Fatigue [FZHR99]. Fe [ACM98]. features [RR80]. February [BK93]. festschrift [Zah94]. Fifth [Kum91]. Final [Ric70d, Ric75f]. Finders [DMNR68]. Findings [GHR95a, GHR95b]. Fine [RTV98b, RTV00b, RTV02]. fingerprint [KAMR04]. Finite [LR75b, Ric85b, LR78a, Ric63a]. First [ACM98, AGH^{+}95, HRV90, Ric91c, Ric96d, Ric97c, ZJ04, Ric96e]. Fitting [Ric75a, Ric78c, Ric70a]. Five [ACD^{+}86, DMNR68, Ric91d]. Fixed [dBR68b]. Flex [Ric86d, MR87d]. FLEX/32 [MR87d]. fonctions [Ric69a]. Force [Ric96b, Ric96c]. Foreword [Ric94d]. Forms [Ric65e, Ric91d]. formulas [Ric68a]. Fortran [LR77, RR73b, Ric84f]. Forum [HJR95]. Foundation [LBCS73, BBC^{+}96]. Fourth [GKM^{+}91]. Framework [CGH^{+}97, MHC^{+}99, MHC^{+}00b, WHR^{+}94b, CGH^{+}00, JRD^{+}01, MHC^{+}00a, CCC^{+}94]. frameworks [HGRB00]. France [ACM88, FCC87, Ger03, IEE03]. Free [RR99]. friction [ABCR05]. Frontiers [Jáj90]. full [Ric95c]. Function [MR92g, MR94e]. Functions [Gar65, MC87, Ric64e, Ric64a, Ric69d, Ric69g, Ric96h, HR67, Mei64, Mei67, Ric61c, Ric61d, Ric64b, Ric65a, Ric69c, Ric70b, Ric71c, Ric97b, Ric00d, Sch69, Ste77]. Funktionen [Mei64, Ric65c]. Future [GHR92, GOP^{+}79, GOP^{+}80, HR00c, HR00b, LHRL93, Ric72, Ric75h, Ric75p, Ric92a, Ric95b, Ric95c, Ric97a, Ric97d, Ric97e, Ric00a, Ric71a, RLC^{+}74].

Fuzzy [JRRH95, JWR^{+}95, JWR^{+}96, JRHR96, RJW^{+}95b, JRRH96, JRHR97, RJW^{+}95a, RJHR97b].

Galerkin [DLRH81, DHLR84, Ric80d, Ric80e, Ric81e, Ric83g]. Galleria [IEE97]. Gas [CFRZ00, FHRZ99a, FRC^{+}00, FRH00, Ric99a, ZFHR99, FHR^{+}00, HCT^{+}02b, HCT^{+}02a, Ric63c, Ric65b]. gas-lubricated [Ric63c, Ric65b]. GasTurbine [HCT^{+}02b]. GasTurbnLab [CFRZ00, FHRZ99b, FRC^{+}00, FHR^{+}00, FHRZ00, HCT^{+}02a]. GAUSS [RR96a, RR02a, HMR72, MR88f, MR90g, MR91f, MR92f, MR93c]. Gautschi [Zah94, Ric94d]. GENCOL [HRM83a, HMR85a]. General [Gar65, HR75, HRM83a, HMR85a, LHHR94b, LHHR94c, LR91, MR88e, Ric70a, Ric81d, Ric84g, RM88b, WHR90, MR89c]. generation [KDM^{+}92]. Geometric [BCR87]. Geometry [CR92b, Ric92c, Ric92b, Ric67d]. Georgia [Ame94]. German [Mei64]. get [RR99]. giving [dBR79]. go [RD68]. gone [Ric03]. GP [Ric75f]. GP-3294OX [Ric75f]. Gram [Ric66a]. Grant [Ric75f]. Graphics [FR93]. Greece [HPP88]. Grid [ABE^{+}03b, CR89, MR92h, CR92a, MR89d, ABE^{+}03a]. Grid-Based [MR92h]. Grids [Ric85a, Ric87a]. Guards [ACR02, ACR04]. Guest [HGBR97a]. Guide [BBRW78, Ric78d, Ric78f, Ric78e, Ric80a, Ric81f, RTV98c, HHR^{+}88]. Guide-Preliminary [Ric78f]. Günter [Ric65c].

h [CR89, CR92a]. h-p [CR89, CR92a]. Hardware [CR92b, HRV89a, MR88a]. hardware-software [MR88a]. Harvard [HW76]. Held [Jáj90, AGH^{+}95, Ger03, MC87, OR70, SKR^{+}93, Tal70, Tra76, VS79]. HERMCOL [HMR85b]. Hermite [DR81, DLRH81, DR84a, DR84b, DHLR84, DR86, HRM83a, HMR85a, HMR85b, LHHR92, LHHR94a, LHHR94b, LHHR94c, LHHR95]. Herndon [IEE95b]. Heterogeneous [JDR^{+}96, JDR^{+}97]. Heuristics [CR92b]. High [AR77, Ano94, DCRS81, Ger03, HR80b, HR82, IFI95, LR75b, LR75c, LR78a, MR90e, MR94a, CHCM^{+}94, GH92, HR92a, LR80b, MR88a]. High-Level [Ger03, GH92]. high-order [LR80b]. History [Cra87, Nas90, RR90a, RR94a]. Hodie [LR75d, LR75e, LR78b]. Homogeneous [MR87c, HCT^{+}02b]. Hong [ZJ04]. honor [BDH00, RD94, Zah94]. Hotel [IEE96, IEE97]. Houston [IEE97]. HSNC’87 [Cra87]. hydrodynamic [Ric65b]. Hypercube [MR88c, MR90b, MR91b]. Hypercubes [MR89e, MR90i, MRW91, MR91g, MR92h, MR89d, MR92e].

ICBA [ZJ04]. ICPP [Sie96]. Ideals [HRR80]. IDL [FR93]. IEEE [IEE97, Ric96e, Ric96d, Ric97c]. IEEE-CS [Ric96d]. IFAC [HRV90, Lei83]. IFIP [ESY84, FCC87, Fos79, GH92, Kal65, Rei82, Wri89, BT01, Ros74]. ihre [Mei64, Ric65c]. II [AvdH91, BS84, DR94, HVC^{+}98, HCR^{+}99, HVC^{+}99, HCR^{+}00, JR90, MR87e, Ric67d, Ric73g, Ric74e, Ric75i, Ric75j, Ric76h, RHD81, dBR68a]. III [Ric77c, Ric74f, Ric74i, Ric76i, VS79]. ILLIAC [LR80a]. IMA [MC87]. IMACS [AvdH91, Ame94, BK92, DO00, HRV90, HR92c, MH89, Ame94, HRV92, Syd97, VS79, VM91]. IMACS/IFAC [HRV90]. Imbalance [MR88c, MR91c, MR90c]. Impact [LR75f, LR77, LR78c]. Implementation [WJH^{+}95, WJH^{+}97]. Implicit [LRT65]. Importance [DR84a, DR86]. IMSL [FR93]. IMSL/IDL [FR93]. Incomplete [MR90g]. indefinite [dBR82]. Independent [Ric68c, Ric86c]. Indiana [BH02, BK93, Lei83, Ric96d, BK93]. Indianapolis [BK93]. Industrial [OR70]. Inelastic [HR73]. INFOMART [SW86]. Information [Ano01, Bor70, Ros74, BCG^{+}00, RR69, HJR95, Kal65]. Initial [LR68b, LR68a]. Innovations [Ano01]. Innovations-2001 [Ano01]. Inserting [ACK^{+}02b]. Instability [CFRZ00, FRH00]. Institute [Ame94, HW76]. INTCOL [HMR85b]. Integral [Ric68c]. Integrated [WCHR92]. Integrating [WHR94a]. Integration [CPR69b, HMR72, CPR69a]. Intelligence [HR92c, IEE95b]. Intelligent [HJR^{+}97b, JWR^{+}95, RJW^{+}95b, WHR^{+}95, WHR^{+}96c, HRV90]. interactions [MR88a]. Interactive [WHR90, KR68]. Interface [BDH^{+}88, DR94, HW76, MR91e, MR93b, MR97, RS83b, RS83a, RS87, RS89, Ric94f, RTV97, RTV98a, RTV98b, RTV99, RTV00b, RTV00a, WHR90, RTV02]. Interfaces [WJH^{+}95, ESY84, WJH^{+}97]. Interior [LHHR94b, LHHR94c]. International [ACM88, ACM98, AGH^{+}95, BH02, Ger03, GKM^{+}91, HPP88, HRV90, HRV92, HRV94, IEE93, IEE95a, IEE95b, IEE96, IEE97, IEE03, KX94, Kum91, LBCS73, MH89, MN95, RVH90, RVH93, SdG90, VS79, ZJ04]. interpolating [Ric61c]. Introduction [Bor70, HGBR97a, Ric61a, RR69, RR73a, RR73b]. IPDPS [Ger03, IEE03]. IPPS [IEE95a]. Ireland [AvdH91, BK92, HR92c, VM91]. irregular [MMR95]. Islander [Big83]. isolated [Ric90a]. issue [BDH00]. Issues [RS83b, RS83a, RS87, RS89]. Iteration [MR88b, MRV90a, Ric80d, Ric80e, dBR63, MR90a, dBR82]. iterations [Ric71e]. Iterative [CHK^{+}92, CHR94, HHR^{+}89a, HHRV91b, HHRV93, HHRV99, LHHR92, LHHR94a, LHHR95, HHR^{+}89b, MRV91, MRV93]. Ithaca [Ano90]. IV [LR80a].

J [Ric67b]. James [Ric74g]. January [LBCS73]. Japan [IFI95]. Jersey [Cra87]. John [BH02, Ano00, BDH00, Bor70, Hai10]. Joint [SW86]. journal [Ric65b]. July [ACM88, AvdH91, Ame94, BK92, Han78, HR92c, MC87, MN95, Tal70, VM91, ZJ04]. June [AFI73, Ano90, FCC87, HPP88, IEE96, IEE97, SdG90, VS79].

Karlin [Ric67b]. Karlsruhe [GH92]. Kernel [WHR^{+}96a, WHR^{+}00, WHR^{+}96a]. Keywords [Ham85, HM90, RH84]. KIVA [Ric99a]. Knots [dBR68b, dBR68a]. Knowledge [HHK^{+}90, HVC^{+}98, HCR^{+}99, HVC^{+}99, KDM^{+}92, RR96b, VHR98, WHR^{+}95, WHR^{+}96c, BCG^{+}00, HHR^{+}91, HCR^{+}00, VHR00].

Knowledge-based [KDM^{+}92]. Knowledge/Data [HVC^{+}98, HVC^{+}99]. knowledge/database [HCR^{+}00]. Kong [ZJ04]. Kutta [Ric60e]. Kyoto [IFI95, IFI95].

Lab [HCT^{+}02b]. Laboratories [Gar65]. Laboratory [CGH^{+}97, HHR^{+}93, CGH^{+}00, HHR^{+}94]. Lafayette [BH02, Lei83]. Lancaster [Tal70]. Language [Ric78i, Ric86c, Ric76a, Ric79b, Ric79d, Ric81g]. Languages [Bor70, JR90, Rei82, Ric75o, Ric81b, Ric85c, RR69, Ric76b, Ric82a]. Large [ABB^{+}00, DDH^{+}98, DCRS81, Ric83f, Ric84a, Zel98]. Latency [MR90b, MR91b]. Laws [HR73]. Lawson [RU68]. Learning [JDR^{+}95, Ric91d, Ric91e, Ric92d, DL91]. Least [Ric83f, dBR68a, dBR68b, Ric71b, Ric84a]. Legendre [HMR72]. Legitimate [Ric80b, Ric81c]. Lehigh [VS79]. Letter [RD68]. Level [Ger03, MR90e, MR94a, RM88a, GH92, MR88a, MR89a, MR90a]. Libraries [Ric89, RB96, Ric96g, RB00]. Library [ADHR73, FR93]. Life [FZHR99]. like [Ric73a]. Line [HHRV91b, HHRV91a, HHRV99, HHRV93]. linéaire [Ric69a]. Linear [HR75, MR89e, Ric84c, RHD81, dBR82, Ric64b]. Lines [RLR00]. linked [Ric95c]. Load [CHH^{+}89a, MR88c, MR90b, MR90c, MR91c, CHH^{+}89b, MR91b, MMR95]. local [KAMR04, Ric67d]. localization [RR80]. London [Han78]. Long [Ric93c]. Lower [NR02]. lubricated [Ric63c, Ric65b]. lubrication [Ric63a]. Lunch [RR99].

Machine [DCRS81, Ric80c, Ric83c]. Machinery [Ric74f]. Machines [HRC^{+}90a, MR91f, MR91h, Ric86d, HRC^{+}90b, HR92a, HCT^{+}02b, HCT^{+}02a, MRV91, MRV93, MR93c]. Macro [Ric81f, RW81]. Macromolecular [MRCH^{+}92, MRCH^{+}93]. Macroprocessor [RRW84, Lev98]. Madison [Ric77c, Sch69, dBG78]. Malo [ACM88]. Managing [HCR^{+}99, HCR^{+}00]. Manual [Ric93e]. Mapper [HHRS90, HHR^{+}88, HHR^{+}89c]. Mapping [CHH^{+}90, CHR94, CHH^{+}91]. March [Ric77c, SKR^{+}93]. Maryland [Jáj90]. Massachusetts [HW76]. Massively [Jáj90, MR92i]. matematicheskoe [Raĭ84]. Math [FR93]. Mathematical [ADHR73, Hai10, Han78, Ric71d, RLC^{+}74, Ric80b, Ric81h, Ric83a, Ric87d, Ric87c, Ric88d, Ric88c, Ric88b, Ric90b, Ric93b, Ric00b, HRV90, KDM^{+}92, Ric81c, Ric77c]. Mathematics [Ame94, Ano90, BH02, KR68, MH89, OR70, Ric77c, Sch69, Syd97, VM91, dBG78, AvdH91, BK92, DO00, S^{+}83]. Matrices [MR89e, dBR64]. Matrichnye [Raĭ84]. Matrix [MR90i, Ric71e, Ric81h]. May [AGH^{+}95, BH02, Cra87, Kal65, Sch69, dBG78]. Measuring [Ric96f]. mechanical [Ric92c, Ric92b]. Mechanisms [ACK^{+}02a, ACK^{+}02b]. meeting [OR70]. Meinardus [Ric65c]. Mellon [Tra76]. Memory [HHR^{+}89a, MRV90a, MR90c, MR91c, MRCH^{+}92, MRCH^{+}93, MR88f, MR91f, MR91h, MR92i, MR92f, HHR^{+}89b, MRV91, MRV93, MMR95, MR93c]. Metalgorithm [Ric73g, Ric75e, Ric75j]. Metalgorithms [HLPR74].

Method [ACK^{+}02b, CHR88b, LR75d, LR75e, LR75c, Ric80d, Ric80e, Ric81e, RVY93, RLR00, CHR88a, LR78b, LR80b, Ric60e, Ric83g, RVY97]. Methodology [FHRZ00, MR90d, Ric79a, VHR98, VHR00]. Methods [CR89, CHK^{+}92, DLRH81, DRR86, GKM^{+}91, HHR^{+}89a, HHRV91a, HHRV99, HLPR75b, HLPR75c, HR78b, HR80b, HRV86, HHR86a, HCR86, HVR88, HRV88b, KHHR95, LHHR94b, LRT65, LR68b, MR88c, MRV90a, MR91e, Ric75d, Ric81e, Ric83d, Ric83e, Ric86e, Ric92e, RTV97, RTV98b, RTV00b, VS79, CR92a, DHLR84, DRR88, HHR^{+}89b, HHRV93, HLPR75a, HLR78, HR82, HHR87, HCR88, KX94, KHHR96, LHHR94c, LRT64b, LR68a, MRV91, MRV93, Mei67, RJW^{+}95a, Ric83i, Ric83g, Ric87e, Ric93f, RTV00a, RTV02]. metric [Ric62b]. Metrics [Ric94e]. Mexico [ACM98]. Michigan [Gar65]. Microprocessors [MSV92]. middleware [HGRB00]. Military [MC87]. Miller [Ric74g]. MIMD [HRV89a, HRC^{+}90a, HRC^{+}90b, MRCH^{+}92, MRCH^{+}93, MMR95]. Minimization [Ric70c]. Mining [VHR99, HVC^{+}00]. Minkowski [Ric67d]. Misalignment [Ric65b]. Mixed [DRR86, DRR88]. Model [AAB^{+}02]. Model-based [AAB^{+}02]. Modeling [HHR86b, HHRS90, MR88a, MR90d, MR94c, MR94d, MR95, NR02, BMR^{+}00, HHR^{+}88, HHR^{+}89c, HVC^{+}00, KDM^{+}92, RV98]. Modelling [Syd97]. Models [Ger03, JDR^{+}96, Ric74d, JDR^{+}97]. Modules [Ric84c, ESY84]. MOL [RLR00]. moment [HR65]. Monterey [BS84]. Morehouse [AGH^{+}95]. Motors [Gar65]. MPSE [HRJ^{+}95, HJR95, MHC^{+}99, MHC^{+}00b]. MPSEs [DJR^{+}96, JRD^{+}01]. Multi [DRJ96, JDR^{+}96, JDR^{+}97, KHHR92, KHHR95, KHHR96, MR90a, MR91c, Ric86d, DHJ^{+}97, MHC^{+}00a]. Multi-Agent [JDR^{+}96, JDR^{+}97]. Multi-Flex [Ric86d]. Multi-level [MR90a]. Multi-Parameterized [KHHR92, KHHR95, KHHR96]. Multi-Physics [DRJ96, DHJ^{+}97]. Multi-processor [MR91c]. Multiagent [JDR^{+}95, DJR^{+}96]. Multicomputer [HRW^{+}96, HRW^{+}98, HRW^{+}00]. Multidimensions [HHRV91a]. Multidisciplinary [HRJ^{+}95, HRR^{+}98, JRD^{+}97, JWH^{+}97, MHC^{+}99, MHC^{+}00b, HJR^{+}97a, HCT^{+}02b, HCT^{+}02a, HJR95]. Multilevel [MR88b]. Multimachine [DCRS81]. multiphysics [MTH^{+}03]. Multiprocessor [HHR^{+}89a, MR90c, HHR^{+}89b]. Multiprocessors [MR88f, MR92i, MR92f]. multivariable [Lei83]. Multivariate [Ric78b, Ric78h, Ric69c, dBR79, Han78]. Mutually [AR99]. MyPYTHIA [HCD^{+}02].

NAPSS [RR66a, RR66b, Ric73a]. NAPSS-like [Ric73a]. National [ACM66, ACM69, LBCS73, AFI73, BBC^{+}96]. nature [RGO^{+}79]. Naval [OR70]. NCUBE [MR90d]. Net [MWHR97a, MWHR97b]. Net//ELLPACK [MWHR97a]. Netcentric [MHC^{+}99, MHC^{+}00b, JRD^{+}01, MHC^{+}00a]. Netherlands [SdG90]. nets [MR88a]. Network [BRW78, BBRW78, DCRS81, DRJ96, HRMW97, HCB^{+}01, JWH^{+}97, MWHR00, TBMR99a, HCT^{+}02b, HCT^{+}02a, TBMR99b].

Network-Based [HCB^{+}01]. Networked [DHRR99, WJH^{+}95, WHR^{+}96b, HJR^{+}97b, WJH^{+}97]. Networks [Ano95, IEE96, IEE97]. Neural [AGH^{+}95, Ano95, IEE96, IEE97, JWR^{+}95]. Neuro [JRRH95, JWR^{+}95, JWR^{+}96, JRHR96, RJW^{+}95b, RJHR97b, JRRH96, JRHR97, RJW^{+}95a]. Neuro-Fuzzy [JRRH95, JWR^{+}95, JWR^{+}96, JRHR96, RJW^{+}95b, RJHR97b, JRRH96, JRHR97, RJW^{+}95a]. Neurobiological [JRHR96, JRHR97]. Newport [Big83, IEE93]. Newsletter [Rob73]. Nice [Ger03, IEE03]. Ninth [HW76]. no [RR99]. Non [MR87c, MR88c, RR86, RVY93, HCT^{+}02b]. Non-Algorithmic [MR88c]. non-homogeneous [HCT^{+}02b]. Non-Overlapping [RVY93]. Non-Rectangular [RR86]. Nonhomogeneous [MR87a, MR89b, HCT^{+}02a]. Nonlinear [OR70, Ric59b, Ric65d, Ric67d, Ric69f, Ric69g, Ric75f, Ric60a, Ric61a, Ric64f, Ric64c, Ric69e, Ric69b, Ric70c, Ric71e, Ric69c]. nonoverlapping [RVY97]. Nonsymmetric [RD81, DR85]. Norfolk [SKR^{+}93]. Norms [RW64]. Notches [Ric68c]. Note [Ric63c, Ric75g, RR80]. Notebook [WJH^{+}95, WJH^{+}97]. Notes [Ric66a, Ano00, Ricxx]. November [IEE95b, SW86]. NSF [Ric75f]. Numeric [Cra87, WCHR92, WHR90, DL91]. Numerical [ACM79, CPR69b, DRR86, GOP^{+}79, GOP^{+}80, HD80, HLPR75c, HRV89b, HRC^{+}90a, HRV90, HRV92, IFI95, LR91, MR87c, MRV90b, OR70, Rei82, Ric67a, Ric70d, Ric74e, Ric75d, Ric75h, Ric75o, Ric75m, RGO^{+}79, Ric79c, Ric81d, Ric83d, Ric83e, Ric83i, Ric84g, RVH90, Ric91f, Ric92e, Ric93f, RVH93, dBG78, CPR69a, DRR88, Fos79, HLR78, HRC^{+}90b, MR91a, Mei64, Mei67, RR02a, Ric63c, RR66a, Ric68b, Ric71a, Ric76b, Ric81g, Ric82a, OR70, RR66b]. Numerics [Ver94]. numerische [Mei64, Ric65c]. NY [Ano90].

obespechenie [Raĭ84]. Object [Ver94, WHR^{+}94b, WHR94d, CCC^{+}94]. Object-Oriented [Ver94, WHR^{+}94b, CCC^{+}94]. Objects [BDH^{+}87, BDH^{+}88, Ric87f, BDH^{+}89, Ric88a, WHR94c]. Observation [Ric75o]. observations [Ric76b]. October [ACM79, ACM98, BT01, IFI95, Jáj90, KX94, OR70, Ric96d]. Office [OR70]. One [Ric96b, Ric96c, RR02a]. one-dimensional [RR02a]. online [RR02a]. Onset [FRH00]. Ontario [MN95]. Open [HRW94]. Operating [Ric74e]. Operators [Ric75i, Ric73e, Ric76h, Ste77]. optimal [dBR79]. Optimization [Ric91e, Ric92c, Ric92b, Ric92d]. Optimizations [MRW91]. Optimizing [ABE^{+}03b, ABE^{+}03a]. Order [HR80b, HRM83b, HMR85c, LHHR94b, LR75c, HR82, LHHR94c, LR80b, RHD81]. Ordering [DR81, DR84b]. Ordinary [LR75e]. Organization [MR90h, MR91f, MR93c]. organized [Han78]. Oriented [MR86, MR88f, MR92f, Ver94, WHR^{+}94b, MR87b, CCC^{+}94]. Origins [RR90b, RR94b]. Orthogonalization [Ric66a, Ric71f]. Other [Ric64e]. Ottawa [BT01]. Outmigrating [NR02]. Outsourcing [AR99, APRS99, AR01a, AR01b]. Overlapping [RVY93]. Overview [Ric96f, AAB^{+}02]. p [CR89, CR92a]. Package [Ric75n]. Pairs [HHRS90, HHR^{+}88, HHR^{+}89c]. panel [RLC^{+}74]. paper [HJR95]. Papers [HR92c, Ric90c, AvdH91, BK92, Cra87, OR70, Ric66a]. Parabolic [ABE^{+}03b, ABE^{+}03a]. paradox [ABCR05]. Parallel [ABB^{+}00, AGH^{+}95, CHH^{+}90, CR92b, CHK^{+}92, CHR94, CHR88b, DDH^{+}98, DFF^{+}03, Ger03, HRP88, HR89, HHRS90, HRC^{+}90a, HRW^{+}96, IEE93, IEE03, Jáj90, JGD87, JR90, Kum91, LR75a, LR79, MR87c, MR87a, MR88d, MR89b, MRV90b, MR94b, MR87d, MR90i, MR91h, MR92i, Ric73g, Ric73f, Ric74i, Ric74h, Ric75j, Ric76i, Ric84c, Ric84f, Ric85c, Ric86e, Ric87e, Rod89, Sie96, SKR^{+}93, WHR90, Wri89, WHR94d, CHH^{+}91, CHR88a, HHR^{+}88, HHR^{+}89c, HRP89, HRC^{+}90b, HRP90, IEE95a, MR88a, MR91a, Ric71e, Ric74c, Ric96f, WHR94c, MR90h, HRP89, HRP90, HR92a]. Parallelism [MR90e, MR94a, Ric86f, Ric86a, MR90f]. parameter [MR90f]. Parameterized [KHHR92, KHHR95, KHHR96]. Parameters [RTV99]. Park [Jáj90]. Partial [BHR78, BRH79, DHR95, GKM^{+}91, HHRV91a, HHRV99, HLPR74, HLPR75b, HLPR75c, HR78a, HR78b, HR80b, HRM83b, HMR85c, HRV86, HVR88, HRP88, LHHR94b, LR75b, MR90i, MR92h, Ric75c, Ric75d, Ric98, VS79, WHCR95, HHRV93, HLPR75a, HLR78, HR80a, HR82, HRP89, HRP90, HR00a, LHHR94c, LRT64b, LRT64a, LR78a, LR78b, MR93c, Ric76f, Ric77a, RHD81, Ric87e, RVY97]. Particular [Ric75q]. Partitioning [CHH^{+}89a, CR92b, HHR84, HHR86a, HHR87, CHH^{+}89b]. Parts [RS83b, Ric83h, Ric89, RS83a, Ric84h, RS87, RS89]. PARVEC [Ric83f]. Pasadena [ACM79]. Past [GOP^{+}79, GOP^{+}80]. Path [Ric68c]. Pattern [JRHR96, JRHR97]. patterns [RLC^{+}74]. PDE [ESY84, ABE^{+}03a, ABE^{+}03b, CHH^{+}89a, CHH^{+}89b, CHH^{+}90, CHH^{+}91, CR92b, CHR94, DR94, ESY84, HHR84, HHR86a, HHR87, HRV89a, HR89, HR92b, HRW94, HRW^{+}96, HRW^{+}98, HRW^{+}00, LHHR94a, MR86, MR87b, MWHR00, MR91d, MR91e, MR92c, MRW91, MR91g, MR91h, MR92i, MR92e, MR93b, MR94c, MR94d, MR95, RR86, Ric75g, Ric80c, Ric83c, Ric84h, Ric86b, CCC^{+}94, Ric94f, TBMR99a, TBMR99b, WHR^{+}94b, WHR94a, WHR90]. PDE-Based [HRW^{+}98]. PDELab [CCC^{+}94, WHR^{+}94b]. PDEPACK [DHR95]. PDES [Ric86e, HHRV91b, HRV88b, JWR^{+}96, MR88b, MR89a, MR90a, MR89d, MR90g, MR91f, Ric86f, Ric86a, RM88a, RVY93, Ric97f]. Pellpack [MWHR97b, HRW^{+}98, HRW^{+}00]. Pennsylvania [KX94, OR70, Tra76, VS79]. perdu [MR92a]. perdue [MR93a]. Performance [ACM98, ABB^{+}00, Ano94, BHR78, BRH79, BDR86, DDH^{+}98, DLRH81, DRR86, Fos79, HHR86b, HHR86a, HRV88a, HRCV88, HCR^{+}99, IFI95, MR90d, MRV90a, MRV91, MRV93, MR88e, MRW91, MR91g, MR92i, MR92e, Ric80c, Ric81e, Ric83b, Ric83g, RM88b, Ric90c, Ric91f, RTV97, VHR98, CHCM^{+}94, DHLR84, DRR88, HHR87, HR92a, HVC^{+}00, HCR^{+}00, LR78b, MR89c, Ric83c, Ric96f, VHR00, VHR99, Fos79]. Perspective [JDR^{+}95, Ric99b]. Perspectives [Ric94c]. Petri [MR88a]. Philadelphia [OR70]. Physical [BDH^{+}87, BDH^{+}88, Ric87f, WHR94d, BDH^{+}89, Ric68a, Ric88a, S^{+}83, WHR94c]. Physics [DRJ96, DHJ^{+}97]. Piecewise [Ric75l, Ric76g, Ric77b, Ric78h]. Pioneer [Hai10]. Pipeline [DCRS81]. Pittsburgh [Tra76]. PL [LR78c]. PL/1 [LR78c]. Plan [ACD^{+}86]. Plans [Ric93c, WCHR92]. Platform [MR91e, WHR94a, MR92d]. Platforms [HRW^{+}96, HRW^{+}98, HRW^{+}00]. POEMS [ABB^{+}00, DDH^{+}98]. Point [CR89, HCR86, RM88b, CR92a, HMR72, HCR88]. Policy [Ric75n, Ric76j, Ric93e]. pollution [RV98]. Polyalgorithm [Ric69f, Ric69b]. polyalgorithms [Ric68b]. Polynomial [DMNR68, Ric65e, Ric75l, Ric76g, Ric77b, Ric78h]. Polynomials [HRM83a, HMR85a, HMR85b, dBR82]. Population [HR78b, Ric86b, RHD81]. portal [HCD^{+}02]. postdocs [Ric91b]. Potential [Ric88e]. Potentials [HR73]. Power [Ric82a]. pp [Ric65c]. PPE [Ric83h]. PPK [WHR^{+}00]. Practice [MR94c, MR94d, MR95]. Preconditioner [MR93d, MR93e]. Preconditioning [MR92g, MR94e]. Prediction [FZHR99]. Preferences [HRR80]. Preliminary [BRW78, MR87e, Ric78f, Ric86d]. Present [GOP^{+}79, GOP^{+}80, Ric72]. presented [Cra87, OR70]. Price [Ric65c]. Princeton [Cra87]. Principles [HRR80]. Private [KAMR04]. Problem [AR83, BS84, DHR95, FRC^{+}00, FCC87, FDA^{+}03, GHR92, GHR94, GHR95a, GHR95b, HRJ^{+}95, HRW^{+}96, HGBR97a, HGBR97b, HRW^{+}98, HR00c, JWR^{+}95, JWR^{+}96, JRD^{+}97, JWH^{+}97, MHC^{+}99, MHC^{+}00b, MR92b, RR66b, Ric74f, Ric74e, Ric74d, Ric75b, Ric76e, Ric84e, Ric85b, Ric89, Ric95a, Ric96a, RB96, Ric96g, Ric00c, WHR^{+}94b, WHR94d, CHCM^{+}94, DHJ^{+}97, FHR^{+}00, GH92, HR65, HR92a, HJR95, HRMW97, HJR^{+}97a, HRR^{+}98, HR00b, HRW^{+}00, HR00a, HCT^{+}02b, HCT^{+}02a, MHC^{+}00a, RR66a, Ric79a, Ric84d, Ric87b, Ric92a, CCC^{+}94, RB00, WHR94c, FCC87]. Problem-Solving [GHR94, GHR95a, GHR95b, HGBR97a, HRW^{+}98, JWR^{+}96, RR66b, RB96, HGBR97b]. Problems [Bor70, DRJ96, DRR86, HR75, HCR86, JR90, KHHR95, MR87d, MR97, OR70, Ric73a, Ric74e, RB81, RD81, Ric83f, Ric85a, Ric85c, RB85, Ric86b, ABCR05, CHR88a, DHJ^{+}97, DRR88, HCR88, KHHR96, MMR95, Ric63a, RR69, Ric84a, Ric87a]. Proc [Ano01]. procedures [Ric65a]. Proceedings [AGH^{+}95, GH92, Gar65, Ger03, Han78, IEE93, IEE96, IEE03, Jáj90, Mak94, Rei82, Ric77c, Ros74, SW86, Tra76, VS79, dBG78, Ame94, BH02, Cra87, DL91, ESY84, FCC87, Fos79, HRV90, HRV92, Kal65, KX94, KR68, LS76, Lei83, LBCS73, MC87, Sch69, Syd97, Tal70, Zah94, ACM66, ACM69, ACM79, ACM98, Big83, BK93, DO00, HW76, IEE95b, Kum91, Sie96, SKR^{+}93, Wri89, ZJ04, HPP88, MN95, SdG90, VM91]. Processing [DR94, Ger03, HRP88, IEE93, IEE03, Kum91, Ric94f, Sie96, SKR^{+}93, HRP89, HRP90, IEE95a, Kal65, Rod89, Ros74]. Processor [BCR87, MR90d, Ric81a, Ric81f, RW81, MR91c, Ric84b]. Processors [MRV90a, Wri89]. Product [HMR72, LRT65, Ric85a, Ric86b, LRT64b, LRT64a, Ric87a]. Production [HD80]. Products [dBR64]. professional [Ano00, Ricxx]. Program [ACK^{+}02b, ACR04, Ric74i, Ric75q, Ric76j, Ric78a, KDM^{+}92, BBC^{+}96, Ric76i]. Programming [ACM79, Big83, DCRS81, GH92, Ger03, HRC^{+}90a, LR75f, LR77, Rei82, Ric75o, Ric75q, Ric76a, Ric78i, Ric79b, Ric81b, Ric82a, HRC^{+}90b, LR78c, Ric76b, Ric81g, GH92, Ric79d]. Programs [CPR69b, Ric76j, Ric93d, Ric94b, CPR69a, Ric91a, Ric93a]. Progress [Ric75c, Ric78a, Ric76f]. project [HWJR95, RDHR86, SCRS80]. Proposal [Ric78a]. Pros [Ric91b]. prospects [Ric73a]. Protection [ACR04]. Prototyping [WHR94d, WHR94c]. PROTRAN [AR83, MR87e, Ric86c]. PSE [FHRZ99b, WHR^{+}96a, MWHR97a, MWHR97b]. PSEs [MWHR00, WHR^{+}96a, WHR^{+}00]. Publication [Ric75o, Ric76b]. Publications [Ric87c, Ric90b]. Purdue [HRV90, HRV92, Ric96d, Zah94, DCRS81, RR90a, Ric94a, RR94a, RR02b, RR04, RR90b, RR94b, WHR^{+}96a]. Purpose [Ric75k, Ric70a]. PYTHIA [HWJR95, HCR^{+}99, HVC^{+}99, HCR^{+}00, WHR^{+}95, WHR^{+}96c, HVC^{+}98]. PYTHIA-II [HCR^{+}99, HVC^{+}99, HCR^{+}00, HVC^{+}98].

Quadratic [HCR86, HCR88, CHR88a]. Quadratic-Spline [HCR86]. Quadrature [LR91, LR79, RR96a, Ric73b, Ric73c, Ric73g, Ric73f, Ric74e, Ric74i, Ric74h, Ric75e, LR80a, RR02a, Ric73d, Ric74c, Ric75j, Ric76i]. Quadratures [LR75a]. QUANTA [SCRS80].

R [Ano00, BH02, Hai10]. Rate [MR88e, RM88b, MR89c]. Rates [LR68b, LR68a, dBR79]. Ratio [Ric85b]. Rational [MR93d, MR93e, Ric65e]. Rationals [Ric64e]. Raymond [Ric74g]. Realistic [RR86]. recherche [MR92a, MR93a]. Recognition [JRHR96, Ric94g, JRHR97]. recommendation [HCD^{+}02]. Recommendations [GHR95a, GHR95b]. recommender [RR02a]. Recommending [HVC^{+}98, HCR^{+}99, HVC^{+}99, HCR^{+}00]. reconsidered [RD68]. Rectangular [HMR85b, RR86, RHD81]. References [Ham85, HM90, RH84]. Refinement [CR89, CR92a]. Regina [LS76]. Relationship [Rei82]. RELAX [MR91e, MR92b, MR92d]. Relaxation [MR91e, MR97, RTV97, RTV98a, RTV98b, RTV99, RTV00b, MR92d, RTV00a, RTV02]. Remark [Ham85, HM90, Lev98]. Remarks [Ric75l, Ric77b]. Report [GOP^{+}79, GOP^{+}80, MR87e, Ric70d, Ric75f, Ric78a, RDHR86, Ric86d]. representations [Ric71e, Ric92c, Ric92b]. Research [DCRS81, GHR92, Gar65, GOP^{+}79, GOP^{+}80, OR70, Ric75o, Ric77c, Ric80b, Ric95b, Ric95c, Ric96b, Ric96c, Sch69, Weg79, dBG78, Ard80, Ric76b, Ric77a, RGO^{+}79, Ric81c, Ric90a, Ric92a]. Reusability [Big83, BP89, Fre87]. Review [Ric65c, Ric67b, Ric74g, RTV97]. Reviews [Bor70]. Revised [HR92c, AvdH91, BK92]. revolution [Ric76c]. RI [Big83]. Rice [BH02, Ano00, BDH00, Bor70, Hai10]. Richardson [dBR82]. Rights [Ric76j]. River [NR02]. Row [MR88f, MR92f]. Royal [MC87]. Runge [Ric60e]. Running [Ric71f].

Sacramento [NR02]. Salmon [NR02]. Samuel [Ric67b, RD94]. Santa [ACM98, IEE95a]. Scalability [MR94b]. Scalable [MSV92, Ric96g, RB00]. Scale [DCRS81]. Scaling [DR84a, DR86]. Scheduling [MR91h, Ric74e]. Scheme [DR81, DR84b]. Schemes [HHRV91b]. Schmidt [Ric66a]. Schwarz [HRV88b, KHHR92, KHHR95, KHHR96]. SciAGents [DJR95a, DJR95b, RTV98c]. Science [BH02, Bor70, BBC^{+}96, BK93, CGH^{+}97, GHR92, GHR94, HW76, HHR^{+}93, HGBR97a, HGBR97b, KWRW93, LHRL93, LBCS73, MC87, RR83, RR93, RR96b, RR90a, Ric91c, Ric93d, Ric94b, Ric94c, RR94a, Ric94g, Ric95b, Ric95c, Ric96b, Ric96c, Ric96d, Ric97c, Ric99b, RR90b, Tra76, Ard80, CGH^{+}00, HHR^{+}94, HRR^{+}98, HGRB00, HR00b, RRH00, RR69, Ric90a, Ric91a, Ric92a, Ric93a, Ric94a, RD94, SCK^{+}96, BBC^{+}96]. Sciences [Ano01, RR02b, RR04, Ric93c, Ric93e, S^{+}83]. Scientific [AGH^{+}95, Cra87, DJR95b, DHRR99, HRV94, HVC^{+}98, HCR^{+}99, HVC^{+}99, HCB^{+}01, JDR^{+}95, JWR^{+}95, JDR^{+}96, MH89, MWHR97a, MWHR97b, MR92d, MSV92, Nas90, RJW^{+}95b, RJHR97a, Ric72, Ric87d, Ric88d, Ric95a, RB96, Ric96g, Ric97d, Ric97e, SKR^{+}93, Syd97, VHR98, WJH^{+}95, WHR^{+}95, WHR^{+}96c, AR01a, AR01b, BT01, BMR^{+}00, BCG^{+}00, DO00, DJR95a, FCC87, GH92, HRV88a, HRCV88, HRV92, HJR^{+}97b, HRMW97, HVC^{+}00, HCR^{+}00, HCD^{+}02, JDR^{+}97, KX94, RJW^{+}95a, RJHR97b, Ric82b, Ric88c, Ric97a, Ric00a, Rod89, VHR00, WJH^{+}97, S^{+}83]. Scope [Ric75k]. Search [AR77]. Second [HRM83b, HMR85c, HRV92, LHHR94b, RVH90, Ver94, LHHR94c, Lei83, RHD81]. Second-Order [HMR85c, LHHR94c]. Secure [AR99, APRS99, AR01b]. securely [AR01a]. Security [ACK^{+}02a, ACK^{+}02b]. Selected [HR92c, AvdH91, BK92]. Selection [HLPR74, HLPR75b, RR96a, RR96b, RJHR97a, Ric74f, Ric74e, Ric74d, Ric75b, Ric76e, Ric91d, HLPR75a, Ric79a]. Self [RD81]. Self-Adjoint [RD81]. Semi [HHR^{+}89a, HHR^{+}89b]. September [DL91, Lei83]. Sequence [Ric60d]. sequential [Ric74c]. Server [MWHR97a, MWHR97b, HR00a]. Servers [JWH^{+}97]. Service [WHR^{+}96b]. services [HCD^{+}02]. Session [Ric96f, RLC^{+}74]. Set [Ric69g]. Sets [AR77]. Seventh [IEE93, KX94, MN95]. several [HHRV93, Ric63d]. shells [Ric92c, Ric92b]. Sheraton [Big83, IEE96]. Sheraton-Islander [Big83]. Short [Ric66a]. Shrivenham [MC87]. SIAM [SKR^{+}93]. SIDEKIC’98 [BCG^{+}00]. Significant [Ric85b]. SIGNUM [ACM79, Rob73]. Simple [Ric81f, RW81, RRW84, Lev98]. Simpson [HMR72]. Simulate [Ric99a]. Simulating [FRC^{+}00, WHCR95, FHR^{+}00]. Simulation [CFRZ00, FHRZ99a, FRH00, HRC^{+}90a, JDR^{+}96, MH89, ZFHR99, BMR^{+}00, HRC^{+}90b, JDR^{+}97, Ric97a, DO00]. simultaneous [Ric60e]. single [MR90f]. Singular [Ric85a, Ric87a]. Singularities [CR89, HR80b, CR92a, HR82]. Site [Ric96a]. Sixth [SKR^{+}93]. sliding [ABCR05]. Smolts [NR02]. Smooth [LR68b, Ric75a, Ric78c, LR68a]. Smoothing [CR68, CR78, MR93b, RW64, RR82]. Society [Han78, OR70, Ric96e, Ric97c]. Söderköping [ESY84]. SOFTLAB [CGH^{+}00, HHR^{+}94, CGH^{+}97, HHR^{+}93]. Software [ACM79, ACM98, ACK^{+}02a, ACR02, ACK^{+}02b, ACR04, BP89, BHR78, BRH79, BH02, CHH^{+}90, CHR94, Hai10, HD80, HR75, HRM83b, HMR85c, HVC^{+}98, HCR^{+}99, IFI95, MR91e, MN95, RJHR97a, Ric71d, Ric75g, Ric75m, Ric75n, Ric79c, Ric80c, Ric80b, Ric81h, Ric82b, RS83b, Ric83d, Ric83e, Ric83h, Ric84h, Ric87d, Ric88d, Ric89, Ric90c, Ric91f, Ric92e, RB96, Ric96g, Ric97d, Ric97e, TBMR99a, VHR98, WHR94a, AR83, BT01, CHH^{+}91, ESY84, Fos79, Fre87, HRV88a, HRCV88, HRV90, HVC^{+}00, HCR^{+}00, HCD^{+}02, MR88a, MR92d, RLC^{+}74, Ric77a, Ric77c, Ric81c, RS83a, Ric83c, Ric83a, Ric83i, RS87, Ric87c, Ric88c, RS89, Ric90b, Ric93b, Ric93f, Ric00a, Ric00b, RB00, TBMR99b, VHR00, Weg79, Fos79]. Solution [HHRV91b, JWR^{+}96, LHHR92, LHHR95, OR70, Ric69f, LRT64b, Ric69b]. Solutions [LR75b, OR70, RR86, LR78a, Ric63c]. Solve [Ric81e, Ric83g]. Solver [MR92i, Ric75g]. Solvers [ABE^{+}03b, BS84, CHR94, LHHR94a, MR91d, MRW91, MR91g, MR91h, MR93b, MR94c, Ric69g, TBMR99a, WHR90, ABE^{+}03a, HR80a, HRV89a, MR92c, MR92e, MR94d, MR95, Ric84d, TBMR99b]. Solving [DHR95, FRC^{+}00, GHR92, GHR94, GHR95a, GHR95b, HR89, HR92b, HRW94, HRJ^{+}95, HRW^{+}96, HGBR97a, HRW^{+}98, HR00c, JWR^{+}95, JWR^{+}96, JRD^{+}97, JWH^{+}97, MHC^{+}99, MHC^{+}00b, MR92b, MR89e, MR90g, MR90i, MR91f, MR92h, MR97, RR66b, Ric75c, RB81, Ric84e, RB85, Ric86f, Ric89, Ric95a, Ric96a, RB96, Ric96g, Ric97f, Ric98, WHR^{+}94b, WHR94a, WHR94d, AR83, CHCM^{+}94, DHJ^{+}97, FHR^{+}00, FCC87, FDA^{+}03, GH92, HR92a, HJR95, HRMW97, HGBR97b, HJR^{+}97a, HRR^{+}98, HR00b, HRW^{+}00, HR00a, HCT^{+}02b, HCT^{+}02a, LR78b, MHC^{+}00a, MMR95, MR89d, MR93c, RR66a, Ric76f, Ric86a, Ric87b, Ric92a, CCC^{+}94, Ric00c, RB00, WHR94c]. Some [CPR69b, Ric75o, Ric76b, CPR69a, RR80]. Sophia [FCC87]. Sourcebook [DFF^{+}03]. space [Ric62b]. Sparse [MR89e, MR90g, MR90i, MR91f, MRW91, MR91g, MR91h, MR92i, MR92e, MR93c, MR90h]. special [BDH00, Sch69]. Specific [BCRR83]. Speedup [LR75a, LR79, MR92a, MR93a]. Spline [CHR88b, HHRV91b, HHRV91a, HHRV99, HRV86, HCR86, HVR88, HRV88b, dBR68b, CHR88a, HHRV93, HCR88, Ric69e, Sch69]. Splines [Ric67c, RR82, dBR68a]. Split [Ric60e]. Splitting [HRV88b, HRV89a, MR86, MR87b]. Splittings [KHHR92]. SPMD [MRV90b, MR91a]. sponsored [OR70]. Spray [ZFHR99]. Springer [Ric65c]. Springer-Verlag [Ric65c]. Squares [Ric83f, dBR68b, Ric71b, Ric84a, dBR68a, Ric71b]. St [ACM88]. Stanford [Wri89]. Star [RM88b]. State [KX94]. Statement [Ric76j, RD68]. States [Sch69]. Statistical [JRHR96, Ric75p, Ric76c, JRHR97]. Statistics [ADHR73, HW76, Ric88b, Ric67b]. Status [RDHR86]. Step [JWR^{+}96]. stochastic [MR88a]. Stockholm [Ros74]. storage [KAMR04]. Strain [Ric68c]. Strategies [CHH^{+}89a, CHH^{+}89b, MMR95]. Strategy [MR92h, MR89d]. Structure [HR73, HRW94, MR90h, MR90i]. Structured [LR77, Ric75q, LR78c]. Structures [BDH^{+}88]. Studden [Ric67b]. Studies [RD94]. Study [RR96b, Ric75d, Ric75q, Ard80]. Subcube [MR92h, MR89d]. Subdomain [HRV89a]. Subroutines [ADHR73]. Subtree [MR92h, MR89d]. Subtree-Subcube [MR92h, MR89d]. Supercomputers [Hwa84, Ric83f, Ric84c, Ric85d, Ric84a]. Supercomputing [ACM88, Hou89, Ric87f, Ric88a, SdG90, MRW^{+}90, HPP88]. Supplement [Ric80a]. Support [JWR^{+}95, JWR^{+}96, JRD^{+}97]. Supportive [Ger03]. Survey [Ric81g]. Survival [NR02]. Sweden [ESY84, Ros74]. Symbolic [HR92c, WCHR92, WHR94a, WHR90, DL91]. Symbolic-Numeric [WCHR92, WHR90, DL91]. Symmetric [DR85, RD81]. Symposia [OR70]. Symposium [BH02, Gar65, Ger03, GKM^{+}91, HW76, IEE93, IEE03, Jáj90, Kum91, LBCS73, Mak94, Tra76, VS79, dBG78, Han78, IEE95a, Lei83, Ric77c, Sch69, Tal70, dBG78]. Synchronization [MR87c, MR88d, MR89b, MR90b, MR90c, MR91c, MR91b, MR87a]. System [ACK^{+}02b, BHR78, BRH79, BDR86, HRP88, HR89, HHK^{+}90, HHRS90, HCR^{+}99, HVC^{+}99, NR02, RR96a, WHCR95, WHR^{+}95, WHR^{+}96c, HHR^{+}88, HHR^{+}89c, HRP89, HRP90, HHR^{+}91, HCR^{+}00, RR02a, RR66a, RR66b]. Systematic [AR77]. Systems [ABB^{+}00, DDH^{+}98, DRJ96, HRV89b, HR92c, HR92b, HRW94, HRV94, JDR^{+}95, JRD^{+}97, KR68, MR90c, MR91c, MRCH^{+}92, MRCH^{+}93, MR92d, MR89e, RJW^{+}95b, Ric67b, Ric74e, Ric75o, RVH90, RVH93, Ric97d, Ric97e, ESY84, HHR84, HRV90, HRV92, Lei83, MR88a, MMR95, Ric73a, Ric76b, Ric00a, VHR99, Zel98, dBR82].

T [MR90d]. Tamperproofing [ACR02]. Task [BCRR83, RJHR97a]. TC [ESY84, FCC87, Fos79]. TC2 [BT01, GH92, Rei82]. TC2/WG [GH92]. TC2/WG2.5 [BT01]. Tchebycheff [Ric59a, Ric59b, Ric60a, Ric60c, Ric60d, Ric61a, Ric61d, Ric62b, Ric63d, Ric64f, Ric67b]. Teaching [Ric91e, Ric92d]. Technical [KWRW93, Ric66a]. Techniques [APRS99, JRHR96, JRHR97, Ric70c]. technological [Lei83]. technologies [HGRB00]. Technology [Ame94, BBC^{+}96, HW76, RS83b, RS83a, RS87, RS89, Weg79]. temps [MR92a, MR93a]. Tensor [LRT64a, LRT65, Ric85a, Ric86b, dBR64, LRT64b, Ric87a]. Term [Ric93c]. terms [Ric70b]. Test [JR90, Ric69g, Ric85c]. Testing [HVC^{+}98, HVC^{+}99]. Texas [IEE97, LBCS73, SW86]. text [Ric95c]. Thatcher [Ric74g]. Their [LR75f, LR77, LR78c, Mei64]. Théorie [Ric69a]. Theory [Mei67, MR94c, MR95, Ric66b, Ric74f, Ric78b, Ric96h, BK92, LBCS73, MR94d, Ric64b, Ric69c, RR80, Ric97b, Ric00d, Tal70, LS76]. Thinker [GHR94]. Thinker/Doer [GHR94]. Third [HRV94, RVH93, VS79]. Time [Ric75q]. Today [Ric85d]. Tomorrow [Ric85d]. TOMS [Ric76j, Ric90c]. Tool [CHH^{+}90, CHH^{+}91, Ric77a, RTV98c]. Tools [IEE95b]. Toronto [MN95]. torques [Ric65b]. Transactions [Ano90]. transformations [Ric60d]. treatment [Mei64]. trends [Ric82a]. Trinity [VM91]. Tuning [RTV98b, RTV00b, RTV02]. Turbine [CFRZ00, FHRZ99a, FRH00, Ric99a, ZFHR99, HCT^{+}02b, HCT^{+}02a].

Turbines [FRC^{+}00, FHR^{+}00]. Turbomachine [FZHR99]. Tutorial [Fre87, Hwa84]. Two [HMR72, HR75, HCR86, LR91, Ric74e, Ric81a, Ric81d, Ric84g, RM88a, HCR88, MR89a, Ric84b]. Two-Dimensional [LR91, Ric84g, Ric84b]. Two-point [HMR72]. Type [HMR72].

uniqueness [Ric67d]. unisolvent [Ric61d]. United [Sch69]. University [Ano90, Han78, HW76, HRV90, HRV92, Jáj90, LS76, LBCS73, Ric77c, Ric96d, Sch69, Tra76, VS79, dBG78, KX94, RR90a, RR94a, RR02b, RR04, RR90b, RR94b]. unobserved [Ric03]. Unstructured [MSV92, MR91h]. Upon [MR90b]. USA [BH02, HRV90, Lei83, Rei82, Wri89, ACM98, AGH^{+}95, Ame94, HRV92, IEE96, IEE97, SKR^{+}93, VS79]. Use [ADHR73, LR75f, LR77, LR78c]. User [BBRW78, Ric78f, Ric78e, Ric80a, Ric81f, RTV98c, HHR^{+}88]. Using [RJW^{+}95a, RB81, RB85, Ric85d, Ric97f, MR88a].

VA [IEE95b, SKR^{+}93]. Validation [FHRZ00]. valuation [VHR98]. Value [HCR86, KHHR95, HCR88, KHHR96]. Vanguard [Ric75p, Ric76c]. Variable [dBR68a, Ric61d]. variables [Ric63d]. Variant [HRV88b]. Vector [JR90, MR87d, Ric84f, Ric85c]. verification [KAMR04]. Verlag [Ric65c]. Versailles [DL91]. Version [BRW78, Ric78f]. Versus [RD81, DR85]. Very [Ric83f, Ric84a]. via [MWHR97a, MWHR97b]. viii [Ric65c]. Virtual [CGH^{+}97, HHR^{+}93, CGH^{+}00, HHR^{+}94]. Vol [Ric69c]. Vol. [Ric64b]. Volumes [Ric90c]. vychisleniya [Raĭ84].

W [Ric74g]. W.G [FCC87]. Walsh [Ric64d]. Walter [Zah94, Ric94d]. Warren [Gar65]. Was [Ric91c]. Washington [IEE96, KR68]. Web [DHR95, MWHR97a, MWHR97b, Ric96a, WHR^{+}96b]. Web//ELLPACK [WHR^{+}96b]. WebPDELab [HR00a]. West [BH02, Lei83]. Westin [IEE97]. WG [GH92, Wri89]. WG2.5 [BT01]. White [HJR95]. Whither [LHRL93, Ric94c]. Wide [MWHR97a, WHR^{+}96b]. William [Ric67b]. Wis [dBG78]. Wisconsin [Ric77c, Sch69, dBG78]. without [KAMR04]. Working [BT01, ESY84, FCC87, Fos79, GH92, Rei82, Wri89]. Workshop [ACM98, Big83, BCG^{+}00, GHR95a, GHR95b, Ger03, IFI95, MN95, Ric78g, Ric78a, Ric83f, Ric84c, Ric96d, Ric96e, Ric97c, Sie96]. Workshops [Ric96d, Ric96e]. World [AvdH91, Ame94, BK92, DO00, HR92c, MH89, Syd97, VM91, MWHR97a, WHR^{+}96b]. WOSP [ACM98]. WWW [DHR95, WJH^{+}95, WJH^{+}97]. WWW//PDEPACK [DHR95].

Year [ACD^{+}86, Ric78a]. York [AFI73, Kal65, AFI73].

Zero [DMNR68].

References

Adve:2002:MBC

[AAB^{+}02] Vikram Adve, Afolami Akinsanmi, James C. Browne, Det Buaklee, Guoshen Deng, Vinh V. Lam, Trevor Morgan, John R. Rice, Gregory J. Rodin, Patricia J. Teller, Gregory Tracy, Mary K. Vernon, and Stephen J. Wright. Model-based control of adaptive applications: an overview. In Parallel and Distributed Processing Symposium., Proceedings International, IPDPS 2002, Abstracts and CD-ROM, pages 172–179. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2002. ISSN 1070-9924 (print), 1558-190X (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.jsp?
tp=&arnumber=1016573.

Adve:2000:PEE

[ABB^{+}00] Vikram S. Adve, Rajive L. Bagrodia, James C. Browne, Ewa Deelman, Aditya Dube, Elias N. Houstis, John R. Rice, Rizos Sakellariou, David J. Sundaram-Stukel, Patricia J. Teller, and Mary K. Vernon. POEMS: End-to-end performance design of large parallel adaptive computational systems. IEEE Transactions on Software Engineering, 26(11):1027–1048, November 2000. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.jsp?

arnumber=881716. Also published in [DDH^{+}98].

Adams:2005:PSC

[ABCR05] G. G. Adams, J. R. Barber, M. Ciavarella, and J. R. Rice. A paradox in sliding contact problems with friction. Journal of Applied Mechanics, 72(3):450–452, 2005. CODEN JAMCAV. ISSN 0021-8936 (print), 1528-9036 (electronic).

Adve:2003:AOAa

[ABE^{+}03a] V. Adve, J. Browne, B. Ensink, J. R. Rice, P. Teller, M. Vernon, and S. Wright. An approach to optimizing adaptive parabolic PDE solvers for the Grid. In Gerndt [Ger03], page 8. ISBN 0-7695-1880-X. ISSN 1070-9924 (print), 1558-190X (electronic). LCCN QA76.642 .I586 2003. URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=1213385.

Adve:2003:AOAb

[ABE^{+}03b] Vikram S. Adve, James C. Browne, Brian Ensink, John R. Rice, Patricia J. Teller, Mary K. Vernon, and Stephen J. Wright. An approach to optimizing adaptive parabolic PDE solvers for the grid. In IEEE [IEE03], page 210. ISBN 0-7695-1926-1. LCCN QA76.58 .M47 2003.

Atallah:1986:FYP

[ACD^{+}86] Mikhail J. Atallah, Douglas E. Comer, H. E. Dunsmore, Greg Frederickson, and John R. Rice. A five year plan for excellence. Technical report TR-651, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, Decem-

ber 1986. 24 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1986/TR%2086-651.pdf.

Atallah:2002:ESS

[ACK^{+}02a] M. J. Atallah, H. Chang, F. Kerschbaum, D. M’Raihi, and J. R. Rice. Enhanced software security mechanisms. U.S. Patent 60/457542, March 24, 2002.

Atallah:2002:SMI

[ACK^{+}02b] M. J. Atallah, H. Chang, F. Kerschbaum, D. M’Raihi, and J. R. Rice. System and method for inserting security mechanisms into a software program. U.S. Patent 60/452541, March 24, 2002.

ACM:1966:PAN

[ACM66] ACM, editor. Proceedings of the 21st ACM National Conference, volume P-66 of ACM publication. Thompson Book Company, Washington, DC, USA, 1966. LCCN ????

ACM:1969:PAN

[ACM69] ACM, editor. Proceedings of 24th ACM National Conference, volume P-69 of ACM publication. ACM Press, New York, NY 10036, USA, 1969. LCCN ????

ACM:1979:PSC

[ACM79] ACM, editor. Proceedings of the SIGNUM Conference on the Programming Environment for Development of Numerical Software, Pasadena, CA, October 1978, ACM SIGNUM Newsletter. ACM Press, New York, NY 10036, USA, 1979. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

ACM:1988:ICS

[ACM88] ACM, editor. 1988 International Conference on Supercomputing, July 4–8, 1988, St. Malo, France. ACM Press, New York, NY 10036, USA, 1988. ISBN 0-89791-272-1. LCCN QA76.5 .I547 1988.

ACM:1998:PFI

[ACM98] ACM, editor. Proceedings of the First International Workshop on Software and Performance: WOSP ’98: October 12–16, 1998, Santa Fe, New Mexico, USA. ACM Press, New York, NY 10036, USA, 1998. ISBN 1-58113-060-0. LCCN QA76.758; QA76.758 .I579 1998; Internet.

Atallah:2002:NGA

[ACR02] M. J. Atallah, H. Chang, and J. R. Rice. New guards for application in software tamperproofing. U.S. Patent 60/396196, July 16, 2002.

Atallah:2004:EGS

[ACR04] M. J. Atallah, H. Chuang, and J. R. Rice. External guards for software program protection. U.S. Patent P00875-US-01, February 9, 2004.

Aird:1973:SUM

[ADHR73] T. Aird, D. Dodson, E. Houstis, and J. Rice. Statistics on the use of mathematical subroutines from a computer center library. ACM SIGNUM Newsletter, 8(4):8–9, October 1973. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

AFIPS:1973:NCC

[AFI73] AFIPS, editor. National Computer Conference and Exposition, June 4–8, 1973, New York, New York, volume 42 of AFIPS conference proceedings. AFIPS Press, Montvale, NJ, USA, 1973. LCCN ????

Aityan:1995:PFI

[AGH^{+}95] S. K. Aityan, L. T. Grujic, R. J. Hathaway, G. S. Ladde, N. Medhin, and M. Sambandham, editors. Proceedings of the First International Conference on Neural, Parallel and Scientific Computations held at Morehouse College, Atlanta, USA, May 28–31, 1995, Proceedings of Neural Parallel and Scientific Computations 1995. Dynamic Publishers, Atlanta, GA, USA, 1995. ISBN 0-9640398-9-3 (hardback) 0-9640398-8-5 (paperback). LCCN QA76.87 .I58

Ames:1994:IPI

[Ame94] William F. Ames, editor. IMACS ’94: proceedings of the 14th IMACS World Congress on Computation and Applied Mathematics: July 11–15, 1994, Georgia Institute of Technology, Atlanta, Georgia, USA. IMACS, Department of Computer Science, Rutgers University, New Brunswick, NJ, USA, 1994. ISBN ???? LCCN ???? Three volumes.

Anonymous:1990:TAC

[Ano90] Anonymous, editor. Transactions of the 8th Army Conference on Applied Mathematics and Computing, Cornell University, Ithaca, NY, June 19–22, 1990. US Army Research Office, Research Triangle Park, NC, USA, 1990. ISBN ???? LCCN ???? ARO Report 91-1.

Anonymous:1994:HPC

[Ano94] Anonymous, editor. High Performance Computing ’94. National Supercomputer Research Center, Singapore, 1994. ISBN ???? LCCN ????

Anonymous:1995:ANN

[Ano95] Anonymous, editor. Artificial Neural Networks in Engineering. ????, ????, 1995. ISBN ???? LCCN ????

Anonymous:1999:A

[Ano99] Anonymous, editor. Agents’99. ACM Press, New York, NY 10036, USA, 1999. ISBN ???? LCCN ????

Anonymous:2000:JRR

[Ano00] Anonymous. John R. Rice: biographical and professional notes. ACM Transactions on Mathematical Software, 26(2):225–226, June 2000. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Anonymous:2001:PIS

[Ano01] Anonymous, editor. Proc. Information Sciences Innovations-2001, (E-Business and Beyond), volume 1621-021. ????, ????, 2001. ISBN ???? LCCN ????

Atallah:1999:SCO

[APRS99] M. J. Atallah, K. N. Pantazopoulis, J. R. Rice, and E. Spafford. Secure computational outsourcing techniques. U.S. Patent P97053, May 1999.

Aird:1977:SSH

[AR77] Thomas J. Aird and John R. Rice. Systematic search in high dimensional sets. SIAM Journal on Numerical Analysis, 14(2): 296–312, April 1977. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Aird:1983:PPS

[AR83] T. J. Aird and J. R. Rice. PROTRAN: Problem solving software. Advances in Engineering Software, 5(??):202–206, ???? 1983. CO-DEN AESODT. ISSN 0141-1195, 0965-9978.

Atallah:1999:MSO

[AR99] M. J. Atallah and J. R. Rice. Mutually secure outsourcing. U.S. Patent P99094, November 1999.

Atallah:2001:OSC

[AR01a] M. J. Atallah and J. R. Rice. Outsourcing scientific computations securely. In Anonymous [Ano01], page ?? ISBN ???? LCCN ????

Atallah:2001:SOS

[AR01b] M. J. Atallah and J. R. Rice. Secure outsourcing of scientific computations. In Zelkowitz [Zel01], pages 215–272. ISBN 0-12-012155-7. LCCN QA76 .A3 vol. 55.

Arden:1980:WCA

[Ard80] Bruce W. Arden, editor. What can be automated?: The computer science and engineering research study (COSERS), volume 3 of The MIT Press series in computer science. MIT Press, Cambridge, MA, USA, 1980. ISBN 0-262-01060-7. 934 pp. LCCN QA76 .W49.

Ames:1991:CAM

[AvdH91] W. F. Ames and P. J. van der Houwen, editors. Computational and applied mathematics, II: differential equations: selected and revised papers from the IMACS World Congress, Dublin, Ireland, July 1991. North-Holland, Amsterdam, The Netherlands, 1991. ISBN 0-444-89702-X. LCCN QA370 .C625 1991.

Brinkman:1996:ANS

[BBC^{+}96] W. F. Brinkman, M. Beasley, R. Cicerone, G. Field, S. Fraser, E. Jaworski, L. Jelinski, A. Mayadar, J. R. Rice, J. Roessner, R. Schmitt, I. Singer, and J. Wright. An Assessment of the National Science Foundation’s Science and Technology Centers Program. National Academy Press, Washington, DC, USA, 1996. ISBN ???? 81 pp. LCCN ????

Boisvert:1978:ENU

[BBRW78] Ronald F. Boisvert, J. Brophy, John R. Rice, and Daniel D. Warner. ELLPACK network user’s guide. Technical report TR-319, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, November 1978. 36

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1978/TR%2078-319.pdf.

Bramley:2000:WSK

[BCG^{+}00] R. Bramley, B. Char, D. Gannon, T. T. Hewett, C. Johnson, and J. R. Rice. Workshop on scientific knowledge, information and computing (SIDEKIC’98). In Houstis et al. [HGRB00], pages 19– 32. ISBN 0-7923-7809-1. LCCN QA76 .E548 2000.

Bajaj:1987:GDP

[BCR87] C. Bajaj, ?. Cui, and John R. Rice. Geometric domain processor. Technical report TR-717, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, Octo-

ber 1987. 16 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1987/TR%2087-717.pdf.

Batz:1983:AST

[BCRR83] Joseph C. Batz, Paul M. Cohen, Samuel T. Redwine, Jr., and John R. Rice. The application-specific task area. Computer, 16(11):78–85, ???? 1983. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic). URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=1654244.

Bajaj:1987:CAP

[BDH^{+}87] C. Bajaj, Wayne R. Dyksen, C. M. Hoffmann, Elias N. Houstis, J. T. Korb, and John R. Rice. Computing about physical objects. Technical report TR-696, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, July 1987.

64 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1987/TR%2087-696.pdf.

Bajaj:1988:ISA

[BDH^{+}88] C. Bajaj, Wayne R. Dyksen, C. M. Hoffmann, E. N. Houstis, J. T. Korb, and John R. Rice. Interface structures I: Abstract structures for computing about physical objects. Technical report TR-792, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, July 1988. 58

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1988/TR%2088-792.pdf.

Bajaj:1989:CAP

[BDH^{+}89] C. Bajaj, W. R. Dyksen, C. M. Hoffmann, E. N. Houstis, J. T. Korb, and J. R. Rice. Computing about physical objects. In Macaraeg and Hussaini [MH89], pages 642–644. CODEN AN-MAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Boisvert:2000:ESI

[BDH00] Ronald F. Boisvert, Wayne R. Dyksen, and Elias N. Houstis. Editorial: special issue in honor of John Rice’s 65th birthday. ACM Transactions on Mathematical Software, 26(2):223, June 2000. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL http://www.acm.org/pubs/citations/journals/
toms/2000-26-2/p223-boisvert/.

Bonomo:1986:EPE

[BDR86] John P. Bonomo, Wayne R. Dyksen, and John R. Rice. The ELLPACK performance evaluation system. Technical report TR-569, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, January 1986. 24

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1986/TR%2086-569.pdf.

Boisvert:2002:CSM

[BH02] Ronald F. Boisvert and Elias Houstis, editors. Computational Science, Mathematics and Software: proceedings of the International Symposium on Computational Science in Celebration of the 65th Birthday of John R. Rice, West Lafayette, Indiana, USA, 22–26 May, 1999. Purdue University Press, West Lafayette, IN, USA, 2002. ISBN 1-55753-250-8 (paperback). xvii + 386 pp. LCCN Q183.9 .I62 1999. Includes DVD-ROM.

Boisvert:1978:SPE

[BHR78] Ronald F. Boisvert, Elias N. Houstis, and John R. Rice. A system for performance evaluation of partial differential equations software. Technical report TR-278, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA,

July 1978. 22 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1978/TR%2078-278.pdf.

Biggerstaff:1983:PWR

[Big83] T. J. Biggerstaff, editor. Proceedings: Workshop on Reusability in Programming 1983, Sheraton-Islander, Newport, RI. ITT Technology, Stratford, CT, USA, 1983. ISBN ???? LCCN ????

Brezinski:1992:CAM

[BK92] Claude Brezinski and Ulrich W. Kulisch, editors. Computational and applied mathematics, I: algorithms and theory: selected and revised papers from the IMACS 13th World Congress, Dublin, Ireland, July 1991. North-Holland, Amsterdam, The Netherlands, 1992. ISBN 0-444-89701-1. LCCN QA76.95 .C62 1991.

Buck:1993:PAC

[BK93] John F. Buck and Stan C. Kwasny, editors. Proceedings, 1993 ACM Computer Science Conference: February 16–18, 1993, Indiana Convention Center, Indianapolis, Indiana. ACM Press, New York, NY 10036, USA, 1993. ISBN 0-89791-558-5, 0-89791-559-3. LCCN QA75.5; QA75.5 .A26 1993; Internet.

Boloni:2000:ABS

[BMR^{+}00] L. Bölöni, D. C. Marinescu, J. R. Rice, P. Tsompanopoulou, and E. A. Vavalis. Agent based scientific simulation and modeling. Concurrency: Practice and Experience, 12(9):845–861, August 10, 2000. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (elec-

tronic). URL http://www3.interscience.wiley.com/cgi-bin/
abstract/72516484/START;   http://www3.interscience.wiley.
com/cgi-bin/fulltext?ID=72516484&PLACEBO=IE.pdf.

Borodin:1970:RIC

[Bor70] Allan Borodin. Reviews: Introduction to Computer Science. Problems, Algorithms, Languages, Information and Computers, by John K. Rice. American Mathematical Monthly, 77(10):1124– 1125, December 1970. CODEN AMMYAE. ISSN 0002-9890 (print), 1930-0972 (electronic).

Biggerstaff:1989:SR

[BP89] Ted J. Biggerstaff and Alan J. Perlis. Software reusability. Addison-Wesley, Reading, MA, USA, 1989. ISBN 0-201-08017-6 (vol.

1), 0-201-50018-3 (vol. 2). xxv + 425 (vol. 1), xiii + 388 (vol. 2) pp. LCCN QA76.76.R47 B543 1989; QA76.76.R47 S64 1989.

Boisvert:1979:SPE

[BRH79] Ronald F. Boisvert, John R. Rice, and Elias N. Houstis. A system for performance evaluation of partial differential equations software. IEEE Transactions on Software Engineering, SE-5(4): 418–425, July/August 1979. CODEN IESEDJ. ISSN 0098-5589 (print), 1939-3520 (electronic). URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?arnumber=1702646.

Boisvert:1978:END

[BRW78] Ronald F. Boisvert, John R. Rice, and Daniel D. Warner. ELL-PACK network documentation preliminary version. Technical report TR-308, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, November 1978.

34 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1978/TR%2078-308.pdf.

Birkhoff:1984:EPS

[BS84] Garrett Birkhoff and Arthur Schoenstadt, editors. Elliptic Problem Solvers II (Monterey, CA, 1983). Academic Press, New York, NY, USA, 1984. ISBN 0-12-100560-7. LCCN QA377 .E533 1983. Proceedings of the Elliptic Problem Solvers Conference, sponsored by the Mathematics and Mechanics Branch, Office of Naval Research, held in Monterey, California, January 10–12, 1983.

Boisvert:2001:ASS

[BT01] Ronald F. Boisvert and Ping Tak Peter Tang, editors. The architecture of scientific software: IFIP TC2/WG2.5 Working Conference on the Architecture of Scientific Software, October 2–4, 2000, Ottawa, Canada, volume 60 of IFIP. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2001. ISBN 0-7923-7339-1. LCCN QA76.758 .I345 2000.

Rice:1994:CCC

[CCC^{+}94] Ann C. Catlin, C. Chui, C. Crabill, E. N. Houstis, S. Markus, J. R. Rice, and S. Weerawarana. PDELab: An object-oriented framework for building problem solving environments for PDE based applications. In Vermeulen [Ver94], pages 79–92. ISBN ???? LCCN ????

Catlin:2000:GSI

[CFRZ00] Ann C. Catlin, S. Fleeter, John R. Rice, and C. Zhou. Gas-TurbnLab simulation of instability in a gas turbine. Technical report TR-00-020, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, November 2000.

8 pp. URL http://www.cs.purdue.edu/research/technical_
reports/2000/TR%2000-020.pdf.

Catlin:1997:SVL

[CGH^{+}97] Ann C. Catlin, M. G. Gaitatzes, Elias N. Houstis, Z. Ma, S. Markus, John R. Rice, N-H Wang, and S. Weerawarana. Softlab: a virtual laboratory framework for computational science. Technical report TR-97-014, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, Febru-

ary 1997. 26 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1997/TR%2097-014.pdf.

Catlin:2000:SVL

[CGH^{+}00] Ann C. Catlin, M. Gaitatzes, E. N. Houstis, Z. Ma, S. Markus, J. R. Rice, N.-H. Wang, and S. Weerawarana. SOFTLAB: a virtual laboratory framework for computational science. In Houstis et al. [HGRB00], chapter 24, pages 301–313. ISBN 0-7923-7809-1. LCCN QA76 .E548 2000.

Cornea-Hasegan:1994:TPS

[CHCM^{+}94] M. A. Cornea-Hasegan, C. Costian, D. C. Marinescu, I. Martin, and J. R. Rice. Towards problem solving environments for high performance computing. In Anonymous [Ano94], pages 354–366. ISBN ???? LCCN ????

Chrisochoides:1989:ALBa

[CHH^{+}89a] N. Chrisochoides, C. Houstis, E. Houstis, S. K. Kortesis, and John R. Rice. Automatic load balanced partitioning strategies for PDE computations. Technical report TR-862, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, February 1989. 11 pp. URL http://www.cs.purdue.

edu/research/technical_reports/1989/TR%2089-862.pdf.

Chrisochoides:1989:ALBb

[CHH^{+}89b] N. Chrisochoides, C. E. Houstis, E. N. Houstis, S. K. Kortesis, and J. R. Rice. Automatic load balanced partitioning strategies for PDE computations. In Houstis [Hou89], pages 99–107. ISBN ???? LCCN ????

Chrisochoides:1990:DDS

[CHH^{+}90] N. P. Chrisochoides, C. E. Houstis, E. N. Houstis, P. N. Papachiou, S. K. Kortesis, and John R. Rice. DOMAIN DECOM-POSER: a software tool for mapping PDE computations to parallel architectures. Technical report TR-1025, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, September 1990. 19 pp. URL http://www.cs.purdue.edu/

research/technical_reports/1990/TR%2090-1025.pdf.

Chrisochoides:1991:DDS

[CHH^{+}91] N. P. Chrisochoides, C. E. Houstis, E. N. Houstis, S. K. Kortesis, P. Papachiou, and J. R. Rice. Domain decomposer: a software tool for mapping PDE computations to parallel architectures. In Glowinski et al. [GKM^{+}91], pages 341–357. ISBN 0-89871-278-5. LCCN QA402.2 .I57 199.

Chrisochoides:1992:PIM

[CHK^{+}92] N. P. Chrisochoides, Elias N. Houstis, S. B. Kim, M. K. Samartzis, and John R. Rice. Parallel iterative methods. Technical report TR-92-035, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, June 1992. 9

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1992/TR%2092-035.pdf.

Christara:1988:PQS

[CHR88a] C. C. Christara, E. N. Houstis, and J. R. Rice. A parallel quadratic spline collocation-capacitance method for elliptic problems. In ACM [ACM88], pages 375–385. ISBN 0-89791-272-1. LCCN QA76.5 .I547 1988.

Christara:1988:PSC

[CHR88b] C. C. Christara, Elias N. Houstis, and John R. Rice. A parallel spline collocation-capacitance method for elliptic parallel differential equations. Technical report TR-735, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107,

USA, January 1988. 27 pp. URL http://www.cs.purdue.edu/
research/technical_reports/1988/TR%2088-735.pdf.

Chrisochoides:1994:MAS

[CHR94] Nikos Chrisochoides, Elias Houstis, and John Rice. Mapping algorithms and software environment for data parallel PDE iterative solvers. Journal of Parallel and Distributed Computing, 21(1):

75–95, April 1994. CODEN JPDCER. ISSN 0743-7315 (print),

1096-0848 (electronic). URL http://www.idealibrary.com/
links/doi/10.1006/jpdc.1994.1043/production;   http:/
/www.idealibrary.com/links/doi/10.1006/jpdc.1994.1043/
production/pdf.

Casaletto:1969:CSNb

[CPR69a] J. J. Casaletto, M. Pickett, and J. R. Rice. A comparison of some numerical integration programs. ACM SIGNUM Newsletter, 4(3): 30–40, October 1969. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Casaletto:1969:CSNa

[CPR69b] J. J. Casaletto, M. Pickett, and John R. Rice. A comparison of some numerical integration programs. Technical report TR-37, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, June 1969. 11

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1969/TR%2069-37.pdf.

Casaletto:1968:SED

[CR68] J. J. Casaletto and John R. Rice. Smoothing and estimation derivatives of equispaced data. Technical report TR-35, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, May 1968. 28 pp. URL http://www.cs.

purdue.edu/research/technical_reports/1968/TR%2068-35.
pdf.

Casaletto:1978:SED

[CR78] J. J. Casaletto and J. R. Rice. Smoothing and estimating derivatives of equispaced data. Applicable Analysis, 8(2):143–152, ???? 1978. CODEN APANCC. ISSN 0003-6811.

Chen:1989:GRP

[CR89] A. T. Chen and John R. Rice. On grid refinement at point singularities for h-p methods. Technical report TR-848, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, January 1989. 118 pp. URL http://www.cs.purdue.

edu/research/technical_reports/1989/TR%2089-848.pdf.

Chen:1992:GRP

[CR92a] A. Chen and J. R. Rice. On grid refinement at point singularities for h-p methods. International Journal for Numerical Methods in Engineering, 33(??):39–57, ???? 1992. CODEN IJNMBH. ISSN 0029-5981 (print), 1097-0207 (electronic).

Chrisochoides:1992:PHP

[CR92b] Nikos P. Chrisochoides and John R. Rice. Partitioning heuristics for PDE computations based on parallel hardware and geometry characteristics, June 1992.

Crane:1987:HAC

[Cra87] G. E. Crane, editor. HSNC’87: ACM Conference on the History of Scientific and Numeric Computation, conference proceedings: papers presented at the Conference, Princeton, New Jersey, May 13-15, 1987. ACM Press, New York, NY 10036, USA, October 1987. ISBN 0-89791-229-2. LCCN QA76 .A25 1987.

deBoor:1978:SRA

[dBG78] Carl de Boor and Gene H. Golub, editors. Symposium on Recent Advances in Numerical Analysis (1978: Madison, Wis.). Recent Advances in Numerical Analysis: Proceedings of a Symposium Conducted by the Mathematics Research Center, the University of Wisconsin, Madison, May 22–24, 1978, volume 41 of Publication of the Mathematics Research Center, the University of Wisconsin, Madison. Army Mathematics Research Center, Academic Press, New York, NY, USA, 1978. ISBN 0-12-208360-1. LCCN QA3 .U45 no. 41; QA297 S994 1978. URL http://catalog.

hathitrust.org/api/volumes/oclc/65991765.html;   http:
//www.gbv.de/dms/hbz/toc/ht001227727.pdf.

deBoor:1963:CAA

[dBR63] Carl de Boor and John R. Rice. Chebyshev approximation by a ∏ \frac{x−r_{i}}{x+s_{i}} and application to ADI iteration. Journal of the Society for Industrial and Applied Mathematics, 11(1):159–169, March 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 1095-712X (electronic).

deBoor:1964:TPC

[dBR64] Carl de Boor and John R. Rice. Tensor products and commutative matrices. Journal of the Society for Industrial and Applied Mathematics, 12(4):892–896, December 1964. CODEN JSIMAV. ISSN 0368-4245 (print), 1095-712X (electronic).

deBoor:1968:LSAb

[dBR68a] Carl de Boor and John R. Rice. Least squares approximation by cubic splines. II: Variable knots. Technical report TR-21, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, April 1968. 39

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1968/TR%2068-21.pdf. IMSL library routine BSVLS.

deBoor:1968:LSCa

[dBR68b] Carl de Boor and John R. Rice. Least squares cubic spline approximation. I: Fixed knots. Technical report TR-20, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, April 1968. 30 pp. URL http://www.cs.purdue.edu/

research/technical_reports/1968/TR%2068-20.pdf. IMSL li-

brary routine BSLSQ.

deBoor:1979:AAM

[dBR79] Carl de Boor and John R. Rice. An adaptive algorithm for multivariate approximation giving optimal convergence rates. Journal of Approximation Theory, 25(4):337–359, 1979. CODEN JAX-TAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

deBoor:1982:EPA

[dBR82] Carl de Boor and John R. Rice. Extremal polynomials with application to Richardson iteration for indefinite linear systems. SIAM Journal on Scientific and Statistical Computing, 3(1):47– 57, March 1982. CODEN SIJCD4. ISSN 0196-5204.

Denning:1981:PMP

[DCRS81] Peter Denning, Douglas E. Comer, John R. Rice, and Lawrence Snyder. The Purdue Multimachine Pipeline: a high bandwidth machine network and programming environment for research in large scale computation. Technical report TR-378, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, January 1981. 59 pp. URL http://www.cs.purdue.

edu/research/technical_reports/1981/TR%2081-378.pdf.

Deelman:1998:PEE

[DDH^{+}98] Ewa Deelman, Aditya Dube, Adolfy Hoisie, Yong Luo, Richard L. Oliver, David Sundaram-Stukel, Harvey Wasserman, Vikram S. Adve, Rajive Bagrodia, James C. Browne, Elias Houstis, Olaf Lubeck, John Rice, Patricia J. Teller, and Mary K. Vernon. PO-EMS: End-to-end performance design of large parallel adaptive computational systems. In ACM [ACM98], pages 18–30. ISBN 1-58113-060-0. LCCN QA76.758; QA76.758 .I579 1998; Internet. Also published in [ABB^{+}00].

Dongarra:2003:SPC

[DFF^{+}03] Jack Dongarra, Ian Foster, Geoffrey Fox, William Gropp, Ken Kennedy, Linda Torczon, and Andy White, editors. The Sourcebook of Parallel Computing. Morgan Kaufmann Publishers, San Francisco, CA, USA, 2003. ISBN 1-55860-871-0. xvi + 842 + 8 pp. LCCN QA76.58 S638 2003. US$59.95.

Drashansky:1997:CPS

[DHJ^{+}97] T. T. Drashansky, E. N. Houstis, A. Joshi, J. R. Rice, and S. Weerawarana. Collaborating problem solving agents for multi-physics problems. In Sydow [Syd97], pages 541–536. ISBN 3-89685-550-6 (set), 3-89685-551-4 (vol. 1), 3-89685-552-2 (vol. 2), 3-89685-553-0 (vol. 3),3-89685-554-9 (vol. 4), 3-89685-555-7 (vol. 5), 3-89685-556-5 (vol. 6). LCCN Q183.9 .I46 1997. In cooperation with R.-P. Schafer, W. Rufeger, and H. Lehmann.

Dyksen:1984:PCG

[DHLR84] W. R. Dyksen, E. N. Houstis, R. E. Lynch, and J. R. Rice. The performance of the collocation and Galerkin methods with Hermite bicubics. SIAM Journal on Numerical Analysis, 21(4):695– 715, August 1984. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Dyksen:1995:WPW

[DHR95] Wayne R. Dyksen, Elias N. Houstis, and John R. Rice. WWW//PDEPACK: a Web based problem solving environment for partial differential equations. Technical report TR-95-063, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, October 1995. 12

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1995/TR%2095-063.pdf.

Drashansky:1999:NAS

[DHRR99] Tzvetan T. Drashansky, Elias N. Houstis, Naren Ramakrishnan, and John R. Rice. Networked agents for scientific computing. Communications of the ACM, 42(3):48–52, 54, March 1999. CODEN CACMA2. ISSN 0001-0782 (print), 1557-

7317 (electronic). URL http://www.acm.org/pubs/articles/
journals/cacm/1999-42-3/p48-drashansky/p48-drashansky.
pdf;   http://www.acm.org/pubs/citations/journals/cacm/
1999-42-3/p48-drashansky/.

Drashansky:1995:SABb

[DJR95a] T. T. Drashansky, A. Joshi, and J. R. Rice. SciAGents: An agent based environment for distributed, cooperative scientific computing. In IEEE [IEE95b], pages 452–459. CODEN ???? ISBN 0-8186-7312-5, 0-8186-7313-3, 0-8186-7314-1 (microfiche). ISSN 1063-6730. LCCN Q334; Q334 .I574, 1995; Q334 .I574 1995;

Q334 .I574, 1995eb. URL http://ieeexplore.ieee.org/stamp/
stamp.jsp?tp=&arnumber=479840.

Drashansky:1995:SABa

[DJR95b] T. T. Drashansky, A. Joshi, and John R. Rice. SciAgents—an agent based environment for distributed, cooperative scientific computing. Technical report TR-95-029, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA,

April 1995. 21 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1995/TR%2095-029.pdf.

Drashansky:1996:MEM

[DJR^{+}96] T. T. Drashansky, A. Joshi, John R. Rice, E. N. Houstis, and S. Weerawarana. A MultiAgent environment for MPSEs. Technical report TR-96-013, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, February 1996.

27 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1996/TR%2096-013.pdf.

Diday:1991:SND

[DL91] Edwin Diday and Y. Lechevallier, editors. Symbolic-numeric data analysis and learning: proceedings of the conference, Versailles, September 18–20, 1991. Nova Science Publishers, New York, NY, USA, 1991. ISBN 1-56072-042-5. LCCN QA278 .S96 1991.

Dyksen:1981:PCG

[DLRH81] Wayne R. Dyksen, Robert E. Lynch, John R. Rice, and Elias N. Houstis. The performance of the collocation and Galerkin methods with Hermite bi-cubics. Technical report TR-413, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, January 1981. 35 pp. URL http://www.cs.purdue.

edu/research/technical_reports/1981/TR%2081-413.pdf.

Dodson:1968:EFP

[DMNR68] David S. Dodson, P. A. Miller, William C. Nylin, Jr., and John R. Rice. An evaluation of five polynomial zero finders. Technical report TR-24, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, June 1968. 13

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1968/TR%2068-24.pdf.

Deville:2000:PIW

[DO00] Michael Deville and Robert Owens, editors. Proceedings: 16th IMACS World Congress on scientific computation, applied mathematics and simulation, volume 104-3. IMACS, Department of Computer Science, Rutgers University, New Brunswick, NJ, USA, 2000. ISBN 3-9522075-1-9. LCCN QA76.4 .P75 2000; QA76.4.P75. 1 computer optical disc.

Dyksen:1981:NOS

[DR81] Wayne R. Dyksen and John R. Rice. A new ordering scheme for the Hermite bicubic collocation equations. Technical report TR-425, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, January 1981. 22

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1981/TR%2081-425.pdf.

Dyksen:1984:ISH

[DR84a] Wayne R. Dyksen and John R. Rice. The importance of scaling for the Hermite bicubic collocation equations. Technical report TR-477, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, March 1984. 20

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1984/TR%2084-477.pdf.

Dyksen:1984:NOS

[DR84b] Wayne R. Dyksen and John R. Rice. A new ordering scheme for the Hermite bicubic collocation equations. In Birkhoff and Schoenstadt [BS84], pages 467–480. ISBN 0-12-100560-7. LCCN QA377 .E533 1983. Proceedings of the Elliptic Problem Solvers Conference, sponsored by the Mathematics and Mechanics Branch, Office of Naval Research, held in Monterey, California, January 10–12,

Dyksen:1985:SVN

[DR85] Wayne R. Dyksen and John R. Rice. Symmetric versus nonsymmetric differencing. SIAM Journal on Scientific and Statistical Computing, 6(1):45–48, January 1985. CODEN SIJCD4. ISSN 0196-5204.

Dyksen:1986:ISH

[DR86] Wayne R. Dyksen and John R. Rice. The importance of scaling for the Hermite bicubic collocation equations. SIAM Journal on Scientific and Statistical Computing, 7(3):707–719, July 1986. CODEN SIJCD4. ISSN 0196-5204.

Drashansky:1994:PPI

[DR94] T. Drashansky and John R. Rice. Processing PDE interface conditions II. Technical report TR-94-066, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, September 1994. 23 pp. URL http://www.cs.purdue.edu/

research/technical_reports/1994/TR%2094-066.pdf.

Drashansky:1996:ABN

[DRJ96] T. Drashansky, John R. Rice, and A. Joshi. Agent based network systems for multi-physics problems. Technical report TR-96-068, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, November 1996. 13

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1996/TR%2096-068.pdf.

Dyksen:1986:PNM

[DRR86] Wayne R. Dyksen, Calvin J. Ribbens, and John R. Rice. The performance of numerical methods for elliptic problems with mixed boundary conditions. Technical report TR-590, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, March 1986. 14 pp. URL http://www.cs.purdue.

edu/research/technical_reports/1986/TR%2086-590.pdf.

Dyksen:1988:PNM

[DRR88] Wayne R. Dyksen, Calvin J. Ribbens, and John R. Rice. The performance of numerical methods for elliptic problems with mixed boundary conditions. Numerical Methods for Partial Differential Equations, 4(4):347–361, ???? 1988. CODEN NMPDEB. ISSN 0749-159x (print), 1098-2426 (electronic).

Engquist:1984:PSM

[ESY84] Björn Engquist, Tom Smedsaas, and N. N. (Nikolai Nikolaevich) Yanenko, editors. PDE software: modules, interfaces, and systems: proceedings of the IFIP TC 2 Working Conference on PDE Software—Modules, Interfaces, and Systems, Söderköping, Sweden, 22–26 August, 1983. North-Holland, Amsterdam, The Netherlands, 1984. ISBN 0-444-87620-0. LCCN QA377 .I44 1983; QA377.I44.

Ford:1987:PSE

[FCC87] Brian J. Ford and Françoise Chaitin-Chatelin, editors. Problem solving environments for scientific computing: proceedings of the IFIP TC 2/W.G. 2.5 Working Conference on Problem Solving Environments for Scientific Computing, Sophia Antipolis, France, 17-21 June, 1985. North-Holland, Amsterdam, The Netherlands, 1987. ISBN 0-444-70254-7. LCCN Q183.9 .I35 1985.

Fox:2003:PSE

[FDA^{+}03] G. Fox, J. J. Dongarra, D. Arnold, H. Casanova, Ann C. Catlin, T. Haupt, E. N. Houstis, and J. R. Rice. Problem solving environments. In Dongarra et al. [DFF^{+}03], pages 409–442. ISBN 1-55860-871-0. LCCN QA76.58 S638 2003. US$59.95.

Fleeter:2000:GPSb

[FHR^{+}00] S. Fleeter, E. N. Houstis, J. R. Rice, C. Zhou, and Ann C. Catlin. GasTurbnLab: a problem solving environment for simulating gas turbines. In Deville and Owens [DO00], page ?? ISBN 3-9522075-1-9. LCCN QA76.4 .P75 2000; QA76.4.P75. 1 computer optical disc.

Fleeter:1999:GTE

[FHRZ99a] S. Fleeter, Elias N. Houstis, John R. Rice, and C. Zhou. Gas turbine engine compressor-combustor dynamics simulation design. Technical report TR-99-006, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, Febru-

ary 1999. 9 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1999/TR%2099-006.pdf.

Fleeter:1999:GPD

[FHRZ99b] S. Fleeter, Elias N. Houstis, John R. Rice, and Chenn Zhou. Gas-TurbnLab PSE DESIGN. Technical report TR-99-002, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, January 1999. 17 pp. URL http://www.

cs.purdue.edu/research/technical_reports/1999/TR%2099-
002.pdf.

Fleeter:2000:VMG

[FHRZ00] S. Fleeter, Elias N. Houstis, John R. Rice, and C. Zhou. Validation methodology for GasTurbnLab. Technical report TR-00-002, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, February 2000. 8

pp. URL http://www.cs.purdue.edu/research/technical_
reports/2000/TR%2000-002.pdf.

Fosdick:1979:PEN

[Fos79] Lloyd Dudley Fosdick, editor. Performance evaluation of numerical software: proceedings of the IFIP TC 2.5 Working Conference on Performance Evaluation of Numerical Software. North-Hol-land, Amsterdam, The Netherlands, 1979. ISBN 0-444-85330-8. LCCN QA297 .I18 1978.

Fu:1993:CII

[FR93] X. Fu and John R. Rice. Comparison of IMSL/IDL with the IMSL Math Library and Exponet Graphics. Technical report TR-93-025, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, April 1993. 63

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1993/TR%2093-025.pdf.

Fleeter:2000:GPSa

[FRC^{+}00] S. Fleeter, John R. Rice, Ann C. Catlin, E. N. Houstis, and Chen Zhou. GasTurbnLab: a problem solving environment for simulating gas turbines. Technical report TR-00-005, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, February 2000. 6 pp. URL http://www.cs.purdue.

edu/research/technical_reports/2000/TR%2000-005.pdf.

Freeman:1987:TSR

[Fre87] Peter Freeman, editor. Tutorial, software reusability. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1987. ISBN 0-8186-0750-5 (paperback), 0-8186-4750-7 (microfiche). vi + 297 pp. LCCN QA76.76.R47 T87 1987.

Fleeter:2000:SIO

[FRH00] S. Fleeter, John R. Rice, and Elias N. Houstis. Simulation of instability onset in a gas turbine. Technical report TR-00-019, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, December 2000. 5

pp. URL http://www.cs.purdue.edu/research/technical_
reports/2000/TR%2000-019.pdf.

Fleeter:1999:FLP

[FZHR99] S. Fleeter, C. Zhou, Elias N. Houstis, and J. R. Rice. Fatigue life prediction of turbomachine blading. Technical report TR-99-030, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, September 1999. 9

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1999/TR%2099-030.pdf.

Garabedian:1965:AFP

[Gar65] Henry L. Garabedian, editor. Approximation of Functions (Proceedings of the Symposium on Approximation of Functions, General Motors Research Laboratories, Warren, Michigan, 1964). Elsevier, Amsterdam, The Netherlands, 1965. LCCN ????

Gerndt:2003:PEI

[Ger03] Michael Gerndt, editor. Proceedings / Eighth International Workshop on High-Level Parallel Programming Models and Supportive Environments, 22 April 2003, Nice, France: held in conjunction with 17th International Parallel and Distributed Processing Symposium (IPDPS). IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2003. ISBN 0-7695-1880-X. ISSN 1070-9924 (print), 1558-190X (electronic). LCCN QA76.642 .I586 2003.

Gaffney:1992:PEH

[GH92] P. W. Gaffney and E. N. Houstis, editors. Programming environments for high-level scientific problem solving: Proceedings of the IFIP TC2/WG 2.5 Working Conference on Programming Environments for High-Level Scientific Problem Solving, Karlsruhe. North-Holland, Amsterdam, The Netherlands, 1992. ISBN 0-444-89176-5. LCCN QA76.6 .I1782 1992.

Gallopoulos:1992:FRD

[GHR92] S. Gallopoulos, E. Houstis, and John R. Rice. Future research directions in problem solving environments for computational science. Technical report TR-92-032, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA,

May 1992. 65 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1992/TR%2092-032.pdf.

Gallopoulos:1994:CTD

[GHR94] Efstratios Gallopoulos, Elias Houstis, and John R. Rice. Computer as thinker/doer: Problem-solving environments for computational science. IEEE Computational Science & Engineering, 1(2):11–23, Summer 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=326669.

Gallopoulos:1995:WPE

[GHR95a] Efstratios Gallopoulos, Elias N. Houstis, and John R. Rice. Workshop on problem-solving environments: Findings and recommendations. ACM Computing Surveys, 27(2):277–279, June 1995. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (elec-

tronic). URL http://www.acm.org/pubs/toc/Abstracts/0360-
0300/210396.html.

Gallopoulos:1995:WPS

[GHR95b] Efstratios Gallopoulos, Elias N. Houstis, and John R. Rice. Workshop on problem-solving environments: Findings and recommendations. ACM Computing Surveys, 27(2):277–279, ???? 1995. CO-DEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic).

Glowinski:1991:FIS

[GKM^{+}91] Roland Glowinski, Yuri A. Kuznetsov, Gérard A. Meurant, Jacques Périaux, and Olof Widlund, editors. Fourth International Symposium on Domain Decomposition Methods for Partial Differential Equations. SIAM Press, Philadelphia, PA, USA, 1991. ISBN 0-89871-278-5. LCCN QA402.2 .I57 199.

Gear:1979:NCR

[GOP^{+}79] C. W. Gear, J. M. Ortega, B. Parlett, J. R. Rice, M. Schultz, L. F. Shampine, and P. Wolfe. Numerical computation: a report on past, present, and future research. ACM SIGNUM Newsletter, 14(si-1):1–48, February 1979. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic). Also published in [GOP^{+}80].

Gear:1980:NCR

[GOP^{+}80] C. W. Gear, J. M. Ortega, B. Parlett, J. R. Rice, M. Schultz, L. F. Shampine, and P. Wolfe. Numerical computation: a report on past, present, and future research. In Arden [Ard80], pages 51–136. ISBN 0-262-01060-7. LCCN QA76 .W49. Also published in [GOP^{+}79].

Haigh:2010:JRR

[Hai10] Thomas Haigh. John R. Rice: Mathematical software pioneer. IEEE Annals of the History of Computing, 32:72–81, 2010. CO-DEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic).

Hamilton:1985:RRK

[Ham85] Dennis E. Hamilton. Remark on “Algorithm 620: References and keywords for Collected Algorithms of the ACM”. ACM Transactions on Mathematical Software, 11(3):305–307, September 1985. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [RH84, HM90].

Handscomb:1978:MAP

[Han78] D. C. (David Christopher) Handscomb, editor. Multivariate approximation: Proceedings of a symposium organized by the London Mathematical Society at the University of Durham, July 21–30, 1977. Academic Press, New York, NY, USA, 1978. ISBN 0-12-323350-X. LCCN QA297.5 .M84.

Houstis:2001:NBS

[HCB^{+}01] Elias N. Houstis, Ann Christine Catlin, Ganesh Balakrishnan, Nitesh Dhanjani, GaHyun Park, John R. Rice, Spyros Lalis, Manolis Stamatogiannakis, and Catherine E. Houstis. Networkbased scientific computing. In Boisvert and Tang [BT01], pages 3–28. ISBN 0-7923-7339-1. LCCN QA76.758 .I345 2000.

Houstis:2002:MRP

[HCD^{+}02] E. N. Houstis, Ann C. Catlin, N. Dhanjani, J. R. Rice, N. Ramakrishnan, and V. Verykios. MyPYTHIA: a recommendation portal for scientific software and services. Concurrency and Computation: Practice and Experience, 14(13–15):1481–1505, November/December 2002. CODEN CCPEBO. ISSN 1532-0626 (print), 1532-0634 (electronic).

Hart:1968:CAa

[HCL^{+}68a] John F. Hart, E. W. Cheney, Charles L. Lawson, Hans J. Maehly, Charles K. Mesztenyi, John R. Rice, Henry G. Thatcher, Jr., and Christoph Witzgall. Computer Approximations. Robert E. Krieger Publishing Company, Huntington, NY, USA, 1968. ISBN 0-88275-642-7. x + 343 pp. LCCN QA 297 C64 1978. Reprinted 1978 with corrections.

Hart:1968:CAb

[HCL^{+}68b] John F. Hart, E. W. Cheney, Charles L. Lawson, Hans J. Maehly, Charles K. Mesztenyi, John R. Rice, Henry G. Thatcher, Jr., and Christoph Witzgall. Computer Approximations. The SIAM series in applied mathematics. Wiley, New York, NY, USA, 1968. ISBN 0-471-35630-1. x + 343 pp. LCCN QA297 .C64.

Houstis:1986:QSC

[HCR86] Elias N. Houstis, C. C. Christara, and John R. Rice. Quadraticspline collocation methods for two point boundary value problems. Technical report TR-584, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, March 1986.

23 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1986/TR%2086-584.pdf.

Houstis:1988:QSC

[HCR88] E. N. Houstis, C. C. Christara, and J. R. Rice. Quadratic spline collocation methods for two point boundary value problems. International Journal for Numerical Methods in Engineering, 26(4): 935–952, ???? 1988. CODEN IJNMBH. ISSN 0029-5981 (print), 1097-0207 (electronic).

Houstis:1999:PIKa

[HCR^{+}99] Elias N. Houstis, Ann C. Catlin, John R. Rice, V. Verykios, N. Ramakrishnan, and C. E. Houstis. PYTHIA-II: a knowledge discovery in databases system for managing performance data and recommending scientific software. Technical report TR-99-031, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, October 1999. 36

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1999/TR%2099-031.pdf.

Houstis:2000:PIK

[HCR^{+}00] Elias N. Houstis, Ann C. Catlin, John R. Rice, Vassilios S. Verykios, Naren Ramakrishnan, and Catherine E. Houstis.

PYTHIA-II: a knowledge/database system for managing performance data and recommending scientific software. ACM Transactions on Mathematical Software, 26(2):227–253, June 2000. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL http://www.acm.org/pubs/citations/journals/
toms/2000-26-2/p227-houstis/. Also published in [BH02].

Houstis:2002:GMP

[HCT^{+}02a] E. N. Houstis, A. C. Catlin, P. Tsompanopoulou, D. Gottfried, G. Balakrishnan, K. Su, and J. R. Rice. GasTurbnLab: a multidisciplinary problem solving environment for gas turbine engine design on a network of nonhomogeneous machines. Journal of Computational and Applied Mathematics, 149(1):83–100, December 1, 2002. CODEN JCAMDI. ISSN 0377-0427 (print),

1879-1778 (electronic). URL http://www.sciencedirect.com/
science/article/pii/S0377042702005228.

Houstis:2002:GLM

[HCT^{+}02b] E. N. Houstis, Ann C. Catlin, P. Tsompanopoulou, D. Gottfried, B. Balakrishnan, K. Su, and J. R. Rice. GasTurbine Lab: a multidisciplinary problem solving environment for gas turbine engine design on a network of non-homogeneous machines. J. Comp. Engr. Math., 149(??):83–100, ???? 2002. CODEN ???? ISSN ????

Hennell:1980:PAN

[HD80] M. A. Hennell and L. M. Delves, editors. Production and Assessment of Numerical Software. Academic Press, New York, NY, USA, 1980. ISBN 0-12-340940-3. LCCN QA297 .C645 1979.

Houstis:1997:GEI

[HGBR97a] Elias Houstis, Efstratios Gallopoulos, Randall Bramley, and John Rice. Guest Editors’ introduction: Problem-solving environments for computational science. IEEE Computational Science & Engineering, 4(3):18–21, July/September 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-

190X (electronic).  URL http://dlib.computer.org/cs/
books/cs1997/pdf/c3018.pdf;   http://www.computer.org/
cse/cs1998/c3018abs.htm.

Houstis:1997:PSE

[HGBR97b] Elias N. Houstis, Efstratios Gallopoulos, Randall Bramley, and John R. Rice. Problem-solving environments for computational science. IEEE Computational Science & Engineering, 4(3):18–21, ???? 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X

(electronic). URL http://ieeexplore.ieee.org/stamp/stamp.
jsp?tp=&arnumber=615427.

Houstis:2000:ETC

[HGRB00] E. N. Houstis, E. Gallopoulos, J. R. Rice, and R. Bramley, editors. Enabling technologies for computational science: frameworks, middleware, and environments, volume SECS 548 of The Kluwer international series in engineering and computer science. Kluwer Academic Publishers Group, Norwell, MA, USA, and Dordrecht, The Netherlands, 2000. ISBN 0-7923-7809-1. xxi + 402 pp. LCCN QA76 .E548 2000.

Houstis:1990:AKB

[HHK^{+}90] C. Houstis, E. Houstis, M. Katzouraki, T. S. Papatheodorou, John R. Rice, and P. Varodoglou. ATHENA: a knowledge base system for //ELLPACK. Technical report TR-950, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, February 1990. 10 pp. URL http://www.cs.purdue.

edu/research/technical_reports/1990/TR%2090-950.pdf.

Houstis:1984:PAP

[HHR84] C. E. Houstis, E. N. Houstis, and J. R. Rice. Partitioning and allocation of PDE computations in distributed systems. In Engquist et al. [ESY84], pages 67–85. ISBN 0-444-87620-0. LCCN QA377 .I44 1983; QA377.I44. With a discussion.

Houstis:1986:PPC

[HHR86a] C. Houstis, E. Houstis, and John R. Rice. Partitioning PDE computations: Methods and performance evaluation. Technical report TR-614, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, July 1986. 37

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1986/TR%2086-614.pdf.

Houstis:1986:PEM

[HHR86b] C. Houstis, E. Houstis, and John R. Rice. Performance evaluation modeling for distributed computing. Technical report TR-576, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, January 1986. 20

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1986/TR%2086-576.pdf.

Houstis:1987:PPC

[HHR87] Catherine E. Houstis, Elias N. Houstis, and John R. Rice. Partitioning PDE computations: methods and performance evaluation. Parallel Computing, 5(1–2):141–163, July 1987. CODEN PACOEJ. ISSN 0167-8191 (print), 1872-7336 (electronic). Proceedings of the international conference on vector and parallel computing—issues in applied research and development (Loen, 1986).

Houstis:1988:UGA

[HHR^{+}88] C. Houstis, E. Houstis, John R. Rice, Alexandrakis, and Samartzis. A User Guide to the algorithm mapper: a system for modeling and evaluating parallel applications/architecture pairs. Technical report TR-793, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, July 1988.

81 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1988/TR%2088-793.pdf.

Hadjidimos:1989:SIMa

[HHR^{+}89a] A. Hadjidimos, E. Houstis, John R. Rice, M. Samartzis, and E. Vavalis. Semi iterative methods on distributed memory multiprocessor architectures. Technical report TR-856, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, February 1989. 12 pp. URL http://www.cs.purdue.

edu/research/technical_reports/1989/TR%2089-856.pdf.

Hadjidimos:1989:SIMb

[HHR^{+}89b] A. Hadjidimos, E. N. Houstis, J. R. Rice, M. K. Samartzis, and E. A. Vavalis. Semi iterative methods on distributed memory multiprocessor architectures. In Houstis [Hou89], pages 82–90. ISBN ???? LCCN ????

Houstis:1989:AMS

[HHR^{+}89c] C. Houstis, E. Houstis, John R. Rice, Alexandrakis, and Samartzis. The algorithm mapper: a system for modeling and evaluating parallel applications/architecture pairs. Technical report TR-854, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, February 1989.

43 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1989/TR%2089-854.pdf.

Houstis:1991:AKB

[HHR^{+}91] C. E. Houstis, E. N. Houstis, J. R. Rice, P. Varodoglou, and T. S. Papatheodorou. Athena: a knowledge base system for //ELL-PACK. In Diday and Lechevallier [DL91], pages 459–467. ISBN 1-56072-042-5. LCCN QA278 .S96 1991.

Hoffmann:1993:SVL

[HHR^{+}93] C. Hoffmann, Elias N. Houstis, John R. Rice, Ann C. Catlin, M. Gaitatzes, S. Weerawarana, N-H Wang, C. Takoudis, and D. Taylor. Softlab — a virtual laboratory for computational science. Technical report TR-93-06l, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA,

September 1993. 13 pp. URL http://www.cs.purdue.edu/
research/technical_reports/1993/TR%2093-06l.pdf.

Hoffmann:1994:SVL

[HHR^{+}94] Christoph M. Hoffman, Elias N. Houstis, John R. Rice, Ann Christine Catlin, Margaret Gaitatzes, Sanjiva Weerawarana, N-H. Linda Wang, Christos G. Takoudis, and David G. Taylor. SoftLab — a virtual laboratory for computational science. Mathematics and Computers in Simulation, 36(4–6):479–491, October 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166

(electronic). URL https://www.sciencedirect.com/science/
article/pii/0378475494900809.

Houstis:1990:AMS

[HHRS90] C. E. Houstis, E. N. Houstis, J. R. Rice, and S. M. Samartzis. The algorithm mapper: a system for modeling and evaluating parallel applications/architecture pairs. In Houstis et al. [HRV90], pages 87–102. ISBN 0-444-88682-6. LCCN QA76.95 .I46 1988.

Hadjidimos:1991:LCS

[HHRV91a] A. Hadjidimos, Elias N. Houstis, John R. Rice, and E. A. Vavalis. Line cubic spline collocation methods for elliptic partial differential equations in multidimensions. Technical report TR-91-040, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, April 1991. 33

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1991/TR%2091-040.pdf.

Hadjidimos:1991:ISL

[HHRV91b] A. Hadjidimos, Elias N. Houstis, John R. Rice, and E. A. Vavalis. On the iterative solution of line spline collocation schemes for elliptic PDEs. Technical report TR-91-020, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107,

USA, March 1991. 27 pp. URL http://www.cs.purdue.edu/
research/technical_reports/1991/TR%2091-020.pdf.

Hadjidimos:1993:ILC

[HHRV93] A. Hadjidimos, E. N. Houstis, J. R. Rice, and E. A. Vavalis. Iterative line cubic spline collocation methods for elliptic partial differential equations in several dimensions. SIAM Journal on Scientific Computing, 14(3):715–734, May 1993. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Hadjidimos:1999:AIL

[HHRV99] A. Hadjidimos, E. N. Houstis, J. R. Rice, and E. Vavalis. Analysis of iterative line spline collocation methods for elliptic partial differential equations. SIAM Journal on Matrix Analysis and Applications, 21(2):508–521, ???? 1999. CODEN SJMAEL. ISSN 0895-4798 (print), 1095-7162 (electronic). URL http://epubs.

siam.org/sam-bin/dbq/article/30982.

Houstis:1995:MMPb

[HJR95] E. N. Houstis, A. Joshi, and J. R. Rice. MPSE: Multidisciplinary problem solving environments. White paper for America in the Age of Information: a forum. Communications on Information and Communications, National Science and Technology Council, ??(??):10, ???? 1995. CODEN ???? ISSN ???? URL http://

www.hpcc.gov/cic/forum/White_Papers.

Houstis:1997:TMP

[HJR^{+}97a] E. N. Houstis, A. Joshi, J. R. Rice, T. T. Drashansky, and S. Weerawarana. Towards multidisciplinary problem solving environments. High Performance Computing Users News, 1(1):6, ???? 1997. CODEN ???? ISSN ???? URL http://www.hpcu.org/

hpcunews/vol1/issue1/mpse.htm.

Houstis:1997:INS

[HJR^{+}97b] E. N. Houstis, A. Joshi, J. R. Rice, S. Weerawarana, and N. Ramakrishnan. Intelligent networked scientific computing. In Sydow [Syd97], pages 785–790. ISBN 3-89685-550-6 (set), 3-89685-551-4 (vol. 1), 3-89685-552-2 (vol. 2), 3-89685-553-0 (vol. 3),3-89685-554-9 (vol. 4), 3-89685-555-7 (vol. 5), 3-89685-556-5 (vol. 6). LCCN Q183.9 .I46 1997. In cooperation with R.-P. Schafer, W. Rufeger, and H. Lehmann.

Houstis:1974:AEP

[HLPR74] Elias N. Houstis, Robert E. Lynch, T. S. Papatheodorou, and John R. Rice. Algorithms for elliptic partial differential equations: Metalgorithms and selection. Technical report TR-121, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, June 1974. 18

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1974/TR%2074-121.pdf.

Houstis:1975:DESb

[HLPR75a] E. N. Houstis, R. E. Lynch, T. S. Papatheodorou, and J. R. Rice. Development, evaluation, and selection of methods for elliptic partial differential equations. Annales de l’Association Internationale pour le Calcul Analogique, 11(??):98–105, ???? 1975. CODEN AAICAE. ISSN 0020-594X.

Houstis:1975:DESa

[HLPR75b] Elias N. Houstis, Robert E. Lynch, T. S. Papatheodorou, and John R. Rice. Development, evaluation & selection of methods for elliptic partial differential equations. Technical report TR-134, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, January 1975. 13

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1975/TR%2075-134.pdf.

Houstis:1975:ENM

[HLPR75c] Elias N. Houstis, Robert E. Lynch, T. S. Papatheodorou, and John R. Rice. Evaluation of numerical methods for elliptic partial differential equations. Technical report TR-204, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, January 1975. 75 pp. URL http://www.cs.purdue.

edu/research/technical_reports/1975/TR%2075-204.pdf.

Houstis:1978:ENM

[HLR78] E. N. Houstis, R. E. Lynch, and J. R. Rice. Evaluation of numerical methods for elliptic partial differential equations. Journal of Computational Physics, 27(3):323–350, June 1978. CODEN JCTPAH. ISSN 0021-9991 (print), 1090-2716 (electronic).

Hopkins:1990:RRK

[HM90] Tim Hopkins and David Morse. Remark on “Algorithm 620: References and keywords for Collected Algorithms of the ACM”. ACM Transactions on Mathematical Software, 16(4):401–403, December 1990. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See [RH84, Ham85].

Houstis:1972:AAP

[HMR72] E. N. Houstis, W. F. Mitchell, and J. R. Rice. ACM Algorithm 438: Product type two-point Gauss–Legendre–Simpson’s integration. Communications of the ACM, 15(6):1071, June 1972. CO-DEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Houstis:1985:AGC

[HMR85a] Elias N. Houstis, W. F. Mitchell, and John R. Rice. Algorithm 637: GENCOL: Collocation of general domains with bicubic Hermite polynomials. ACM Transactions on Mathematical Software, 11(4):413–415, December 1985. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://www.acm.org/

pubs/citations/journals/toms/1985-11-4/p413-houstis/.

Houstis:1985:AIH

[HMR85b] Elias N. Houstis, W. F. Mitchell, and John R. Rice. Algorithm 638: INTCOL and HERMCOL: Collocation on rectangular domains with bicubic Hermite polynomials. ACM Transactions on Mathematical Software, 11(4):416–418, December 1985. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL http://www.acm.org/pubs/citations/journals/
toms/1985-11-4/p416-houstis/.

Houstis:1985:CSS

[HMR85c] Elias N. Houstis, W. F. Mitchell, and John R. Rice. Collocation software for second-order elliptic partial differential equations. ACM Transactions on Mathematical Software, 11(4): 379–412, December 1985. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://www.acm.org/

pubs/citations/journals/toms/1985-11-4/p379-houstis/;
 http://www.acm.org/pubs/toc/Abstracts/toms/6191.html.

Houstis:1989:S

[Hou89] E. N. Houstis, editor. Supercomputing 1989. ACM Press, New York, NY 10036, USA, 1989. ISBN ???? LCCN ????

Houstis:1988:SIC

[HPP88] E. N. Houstis, T. S. Papatheodorou, and C. D. Polychronopoulos, editors. Supercomputing: 1st International Conference, Athens, Greece, June 8–12, 1987: proceedings, volume 297 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988. CODEN LNCSD9. ISBN 0-387-18991-2 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43

no.297. URL http://link.springer-ny.com/link/service/
series/0558/tocs/t0297.htm;   http://www.springerlink.
com/openurl.asp?genre=issue&issn=0302-9743&volume=297.

The conference was organized and sponsored by the Computer Technology Institute (C.T.I.) of Greece.

Hobby:1965:MPA

[HR65] Charles R. Hobby and John R. Rice. A moment problem in L_{1} approximation. Proceedings of the American Mathematical Society, 16(??):665–670, 1965. CODEN PAMYAR. ISSN 0002-9939 (print), 1088-6826 (electronic).

Hobby:1967:ACF

[HR67] C. R. Hobby and John R. Rice. Approximation from a curve of functions. Archive for rational mechanics and analysis, 24(??): 91–106, 1967. CODEN AVRMAW. ISSN 0003-9527 (print), 1432-0673 (electronic).

Hill:1973:EPS

[HR73] R. Hill and J. R. Rice. Elastic potentials and the structure of inelastic constitutive laws. SIAM Journal on Applied Mathematics, 25(3):448–461, November 1973. CODEN SMJMAP. ISSN 0036-1399 (print), 1095-712X (electronic).

Houstis:1975:SLE

[HR75] Elias N. Houstis and John R. Rice. Software for linear elliptic problems on general two dimensional domains. Technical report TR-220, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, January 1975.

23 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1975/TR%2075-220.pdf.

Houstis:1978:EDC

[HR78a] Elias N. Houstis and John R. Rice. An experimental design for the computational evaluation of elliptic partial differential equations. Technical report TR-264, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, May 1978.

17 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1978/TR%2078-264.pdf.

Houstis:1978:PPD

[HR78b] Elias N. Houstis and John R. Rice. A population of partial differential equations for evaluating methods. Technical report TR-263, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, May 1978. 74

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1978/TR%2078-263.pdf.

Houstis:1980:EDC

[HR80a] E. N. Houstis and J. R. Rice. An experiment design for the computational evaluation of elliptic partial differential equation solvers. In Hennell and Delves [HD80], pages 57–66. ISBN 0-12-340940-3. LCCN QA297 .C645 1979.

Houstis:1980:HOM

[HR80b] Elias N. Houstis and John R. Rice. High order methods for elliptic partial differential equations with singularities. Technical report TR-341, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, January 1980.

26 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1980/TR%2080-341.pdf.

Houstis:1982:HOM

[HR82] Elias N. Houstis and John R. Rice. High order methods for elliptic partial differential equations with singularities. International Journal for Numerical Methods in Engineering, 18(5):737–754, ???? 1982. CODEN IJNMBH. ISSN 0029-5981 (print), 1097-0207 (electronic).

Houstis:1989:PEP

[HR89] Elias N. Houstis and John R. Rice. Parallel (//) ELLPACK PDE solving system. Technical report TR-912, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107,

USA, October 1989. 64 pp. URL http://www.cs.purdue.edu/
research/technical_reports/1989/TR%2089-912.pdf.

Houstis:1992:PED

[HR92a] E. N. Houstis and J. R. Rice. Parallel ELLPACK: a development and problem solving environment for high performance computing machines. In Gaffney and Houstis [GH92], pages 229–241. ISBN 0-444-89176-5. LCCN QA76.6 .I1782 1992.

Houstis:1992:APS

[HR92b] Elias N. Houstis and John R. Rice. The architecture of PDE solving systems. Technical report TR-92-022, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107,

USA, April 1992. 14 pp. URL http://www.cs.purdue.edu/
research/technical_reports/1992/TR%2092-022.pdf.

Houstis:1992:AIE

[HR92c] Elias N. Houstis and John Rischard Rice, editors. Artificial Intelligence, Expert Systems and Symbolic Computing. Selected and Revised Papers from the IMACS 13th World Congress, Dublin, Ireland, July 22–26, 1991. North-Holland, Amsterdam, The Netherlands, 1992. ISBN 0-444-89703-8. LCCN Q334 .I45 1991.

Houstis:2000:WSP

[HR00a] E. N. Houstis and J. R. Rice. The WebPDELab server: a problem solving environment for partial differential equations. In Deville and Owens [DO00], page ?? ISBN 3-9522075-1-9. LCCN QA76.4 .P75 2000; QA76.4.P75. 1 computer optical disc.

Houstis:2000:FPSb

[HR00b] Elias N. Houstis and John R. Rice. Future problem solving environments for computational science. Mathematics and Computers in Simulation, 54(4–5):243–257, ???? 2000. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic).

Houstis:2000:FPSa

[HR00c] Elias N. Houstis and John R. Rice. On the future of problem solving environments. Technical report TR-00-009, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, March 2000. 79 pp. URL http://www.cs.purdue.

edu/research/technical_reports/2000/TR%2000-009.pdf.

Houstis:1990:ENSa

[HRC^{+}90a] E. Houstis, J. Rice, N. Chrisochoides, H. Karathanasis, P. Papachiou, M. Samartzis, E. Vavalis, and K. Wang. //ELLPACK: a numerical simulation programming environment for parallel MIMD machines. Technical report TR-949, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107,

USA, January 1990. 26 pp. URL http://www.cs.purdue.edu/
research/technical_reports/1990/TR%2090-949.pdf.

Houstis:1990:ENSb

[HRC^{+}90b] E. N. Houstis, J. R. Rice, N. Chrisochoides, H. Karathanasis, P. Papachiou, M. K. Samartzis, E. A. Vavalis, K. Wang, and S. Weerawarana. //ELLPACK: a numerical simulation programming environment for parallel MIMD machines. In Sopka and de Groot [SdG90], pages 96–107. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Houstis:1988:PSS

[HRCV88] E. N. Houstis, J. R. Rice, C. C. Christara, and E. A. Vavalis. Performance of scientific software. In Mathematical aspects of scientific software (Minneapolis, Minn., 1986/87), volume 14 of IMA Vol. Math. Appl., pages 123–155. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988. ISBN 0-387-96706-0.

Houstis:1995:MMPa

[HRJ^{+}95] Elias N. Houstis, John R. Rice, A. Joshi, S. Weerawarana, E. Sacks, V. Rego, N. H. L. Wang, C. Takoudis, A. H. Sameh, and E. Gallopoulos. MPSE: Multidisciplinary problem solving environments. Technical report TR-95-047, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA,

July 1995. 10 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1995/TR%2095-047.pdf.

Houstis:1983:AGC

[HRM83a] Elias N. Houstis, John R. Rice, and W. F. Mitchell. Algorithm GENCOL: Collocation on general domains with bi-cubic Hermite polynomials. Technical report TR-444, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107,

USA, January 1983. 4 pp. URL http://www.cs.purdue.edu/
research/technical_reports/1983/TR%2083-444.pdf.

Houstis:1983:CSS

[HRM83b] Elias N. Houstis, John R. Rice, and W. F. Mitchell. Collocation software for second order elliptic partial differential equations. Technical report TR-446, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, Jan-

uary 1983. 40 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1983/TR%2083-446.pdf.

Houstis:1997:NBS

[HRMW97] E. N. Houstis, J. R. Rice, S. Markus, and S. Weerawarana. Network based scientific problem solving environments. High Performance Computing Users News, 1(1):5, ???? 1997. CODEN

???? ISSN ???? URL http://www.hpcu.org/hpcunews/vol1/
issue1/netswp.htm.

Houstis:1988:PEE

[HRP88] Elias N. Houstis, John R. Rice, and T. S. Papatheodorou. Parallel (//) ELLPACK: An expert system for parallel processing of partial differential equations. Technical report TR-831, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, November 1988. 16

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1988/TR%2088-831.pdf.

Houstis:1989:PEE

[HRP89] E. N. Houstis, J. R. Rice, and T. S. Papatheodorou. Parallel (∥) ELLPACK: an expert system for parallel processing of partial differential equations. Mathematics and Computers in Simulation, 31(4–5):497–507, October 1989. CO-DEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (elec-

tronic).  URL https://www.sciencedirect.com/science/
article/pii/0378475489901432. Reprinted in [HRP90].

Houstis:1990:PEE

[HRP90] E. N. Houstis, J. R. Rice, and T. S. Papatheodorou. Parallel ELL-PACK: An expert system for parallel processing of partial differential equations. In Houstis et al. [HRV90], pages 63–73. ISBN 0-444-88682-6. LCCN QA76.95 .I46 1988. Reprint of [HRP89].

Hull:1980:PPI

[HRR80] Thomas E. Hull, Christian H. Reinsch, and John R. Rice. Principles, preferences and ideals for computer arithmetic. Technical report TR-339, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, June 1, 1980.

13 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1980/TR%2080-339.pdf.

Houstis:1998:MPS

[HRR^{+}98] E. N. Houstis, J. R. Rice, N. Ramakrishnan, T. T. Drashansky, S. Weerawarana, A. Joshi, and C. E. Houstis. Multidisciplinary problem solving environments for computational science.

In Zelkowitz [Zel98], pages 401–438. ISBN 0-12-012146-8. LCCN QA76 .A346 1998eb.

Houstis:1986:CHC

[HRV86] Elias N. Houstis, John R. Rice, and Vavalis. Convergence of O(h^{4}) cubic spline collocation methods for elliptic partial differential equations. Technical report TR-596, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA,

April 1986. 29 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1986/TR%2086-596.pdf.

Houstis:1988:PES

[HRV88a] E. N. Houstis, J. R. Rice, and E. A. Vavalis. Performance evaluation of scientific software. In Rice [Ric88d], pages 123–155. ISBN 3-540-96706-0. LCCN QA76.76.D47 M366 1988.

Houstis:1988:SSV

[HRV88b] Elias N. Houstis, John R. Rice, and Vavalis. A Schwarz splitting variant of cubic spline collocation methods for elliptic PDEs. Technical report TR-745, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, March 1988.

21 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1988/TR%2088-745.pdf.

Houstis:1989:BMH

[HRV89a] E. Houstis, John R. Rice, and E. Vavalis. Benchmarking of MIMD hardware on subdomain splitting elliptic PDE solvers. Technical report TR-874, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, March 1989.

16 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1989/TR%2089-874.pdf.

Houstis:1989:ESN

[HRV89b] E. N. Houstis, J. R. Rice, and R. Vichnevetsky. Expert systems for numerical computing. Mathematics and Computers in Simulation, 31(??):315–517, ???? 1989. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). Special double issue.

Houstis:1990:IMS

[HRV90] E. N. Houstis, John Rischard Rice, and Robert Vichnevetsky, editors. Intelligent mathematical software systems: proceedings of the First IMACS/IFAC International Conference on Expert Systems for Numerical Computing, Purdue University, USA., 5–7 December, 1988. North-Holland, Amsterdam, The Netherlands, 1990. ISBN 0-444-88682-6. LCCN QA76.95 .I46 1988.

Houstis:1992:ESS

[HRV92] E. N. Houstis, J. R. Rice, and R. Vichnevetsky, editors. Expert systems for scientific computing: proceedings of the Second IMACS International Conference on Expert Systems for Numerical Computing, Purdue University, USA, 24–26 April, 1990. Elsevier Science Publishers, Amsterdam, The Netherlands, 1992. ISBN 0-444-89226-5. LCCN QA76.95 .I48 1990.

Houstis:1994:TIC

[HRV94] E. N. Houstis, J. R. Rice, and R. Vichnevetsky. Third international conference on expert systems for scientific computing. Mathematics and Computers in Simulation, 36(4–6):267–503, October 1994. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). Special triple issue.

Houstis:1994:OSP

[HRW94] Elias N. Houstis, John R. Rice, and S. Weerawarana. An open structure for PDE solving systems. Technical report TR-94-035, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, May 1994. 6

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1994/TR%2094-035.pdf.

Houstis:1996:PEP

[HRW^{+}96] Elias N. Houstis, John R. Rice, S. Weerawarana, Ann C. Catlin, P. Papachiou, K-Y Wang, and M. Gaitatzes. Parallel (//) ELLPACK: a problem solving environment for PDE based applications on multicomputer platforms. Technical report TR-96-070, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, November 1996. 34

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1996/TR%2096-070.pdf.

Houstis:1998:PPS

[HRW^{+}98] Elias N. Houstis, John R. Rice, Sanjiva Weerawarana, Ann C. Catlin, P. Papachiou, K.-Y. Wang, and M. Gaitatzes. PELL-PACK: a problem-solving environment for PDE-based applications on multicomputer platforms. ACM Transactions on Mathematical Software, 24(1):30–73, ???? 1998. CODEN ACMSCU.

ISSN 0098-3500 (print), 1557-7295 (electronic). URL http:

//www.acm.org/pubs/citations/journals/toms/1998-24-1/

p30-houstis/. Abridged version in [HRW^{+}00].

Houstis:2000:PPS

[HRW^{+}00] E. N. Houstis, J. R. Rice, S. Weerawarana, A. C. Catlin, P. Papachiou, K. Y. Yang, and M. Gaitatzes. PELLPACK: a problem solving environment for PDE based applications on multicomputer platforms. In Houstis et al. [HGRB00], chapter 14, pages 171–185. ISBN 0-7923-7809-1. LCCN QA76 .E548 2000. Full version in [HRW^{+}98].

Houstis:1998:PIK

[HVC^{+}98] Elias N. Houstis, V. S. Verykios, Ann C. Catlin, N. Ramakrishnan, and John R. Rice. Pythia-II: a knowledge/data base for testing and recommending scientific software. Technical report TR-98-031, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, September 1998.

33 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1998/TR%2098-031.pdf.

Houstis:1999:PIKb

[HVC^{+}99] Elias N. Houstis, V. S. Verykios, Ann C. Catlin, N. Ramakrishnan, and John R. Rice. PYTHIA-II: a knowledge/data base system for testing and recommending scientific. Technical report TR-99-012, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, April 1999. 33

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1999/TR%2099-012.pdf.

Houstis:2000:DME

[HVC^{+}00] E. N. Houstis, V. S. Verykios, Ann C. Catlin, N. Ramakrishnan, and J. R. Rice. A data mining environment for modeling the performance of scientific software. In Houstis et al. [HGRB00], chapter 21, pages 261–271. ISBN 0-7923-7809-1. LCCN QA76 .E548 2000.

Houstis:1988:CCS

[HVR88] E. N. Houstis, E. A. Vavalis, and J. R. Rice. Convergence of O(h^{4}) cubic spline collocation methods for elliptic partial differential equations. SIAM Journal on Numerical Analysis, 25(1):54– 74, February 1988. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Hoaglin:1976:PNI

[HW76] David C. Hoaglin and Roy E. Welsch, editors. Proceedings of the Ninth Interface Symposium on Computer Science and Statistics, Harvard University, Massachusetts Institute of Technology, April 1–2, 1976. Prindle, Weber and Schmidt, Boston, MA, USA, 1976. ISBN 0-87150-237-2. LCCN QA276.A1 I53 1976.

Hwang:1984:TSP

[Hwa84] Kai Hwang, editor. Tutorial—Supercomputers: Design and Applications. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1984. ISBN 0-8186-0581-2. viii + 640 pp. LCCN TK 7888.3 H82 1984.

Houstis:1995:PP

[HWJR95] E. N. Houstis, S. Weerawarana, A. Joshi, and J. R. Rice. The PYTHIA project. In Aityan et al. [AGH^{+}95], pages 215– 218. ISBN 0-9640398-9-3 (hardback) 0-9640398-8-5 (paperback). LCCN QA76.87 .I58 1995.

IEEE:1993:PSI

[IEE93] IEEE, editor. Proceedings / Seventh International Parallel Processing Symposium, April 13–16, 1993, Newport Beach, California. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-3442-1. LCCN QA 76.58 I56 1993. IEEE catalog no. 93TH0513-2.

IEEE:1995:IIP

[IEE95a] IEEE, editor. IPPS ’95: 9th International parallel processing symposium — April 25–28, 1995, Santa Barbara, CA, International Parallel Processing Symposium. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7074-6. ISSN 1063-7133. LCCN QA76.58 .I58 1995.

IEEE:1995:PIC

[IEE95b] IEEE, editor. Proceedings of the 7th International Conference: Tools with Artificial Intelligence, Herndon, VA, November 1995. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. CODEN ???? ISBN 0-8186-7312-5, 0-8186-7313-3, 0-8186-7314-1 (microfiche). ISSN 1063-6730. LCCN Q334; Q334 .I574, 1995; Q334 .I574 1995; Q334 .I574, 1995eb.

IEEE:1996:PIIa

[IEE96] IEEE, editor. Proceedings of the 1996 IEEE International Conference on Neural Networks, June 3–6, 1996, Sheraton Washington Hotel, Washington, DC, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-7803-3210-5 (softbound), 0-7803-3211-3 (casebound), 0-7803-3212-1 (microfiche). LCCN QA76.87.I3437 1996. Four volumes. IEEE catalog number 96CH35907.

IEEE:1997:IIC

[IEE97] IEEE, editor. The 1997 IEEE International Conference on Neural Networks, June 9–12, 1997, Westin Galleria Hotel, Houston, Texas, USA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-7803-4122-8, 0-7803-4123-6, 0-7803-4124-4. LCCN QA76.87 .I3437 1997; QA76.87; QA76.87 .I3437 1997eb; Internet; QA76.87 .I347 1997. Four volumes. IEEE catalog number: 97CH36109.

IEEE:2003:PIP

[IEE03] IEEE, editor. Proceedings: 17th International Parallel & Distributed Processing Symposium: IPDPS 2003: 22–26 April, 2003, Nice, France. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2003. ISBN 0-7695-1926-1. LCCN QA76.58 .M47 2003.

IFIP:1995:KWC

[IFI95] IFIP Working Group 2.5, editor. Kyoto Workshop 1995: Current Directions in Numerical Software and High Performance Computing, 19–20 October 1995, Kyoto, Japan. ????, ????, 1995. ISBN

???? LCCN ???? URL http://www.nsc.liu.se/~boein/ifip/
kyoto/kyoto.html#reid;   http://www.nsc.liu.se/~boein/
ifip/kyoto/workshop-info/proceedings/.

Jaja:1990:SFM

[Jáj90] Joseph Jájá, editor. The 3rd Symposium on the Frontiers of Massively Parallel Computation: Proceedings, Held at the University of Maryland, College Park, Maryland, October 8–10, 1990. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1990. ISBN 0-8186-9053-4. LCCN QA76.58 .S95 1990.

Joshi:1995:LAM

[JDR^{+}95] A. Joshi, T. Drashansky, John R. Rice, S. Weerawarana, and E. Houstis. On learning and adaptation in multiagent systems: a scientific computing perspective. Technical report TR-95-040, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, June 1995. 32

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1995/TR%2095-040.pdf.

Joshi:1996:MAS

[JDR^{+}96] A. Joshi, T. Drashansky, John R. Rice, S. Weerawarana, and Elias N. Houstis. Multi-agent simulation of complex heterogeneous models in scientific computing. Technical report TR-96-025, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, May 1996. 33

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1996/TR%2096-025.pdf.

Joshi:1997:MAS

[JDR^{+}97] A. Joshi, T. T. Drashansky, J. R. Rice, S. Weerawarana, and E. N. Houstis. Multi-agent simulation of complex heterogeneous models in scientific computing. Math. Comp. Soft., 44(??):43–59, ???? 1997. CODEN ???? ISSN ????

Jamieson:1987:CPA

[JGD87] Leah H. Jamieson, Dennis B. Gannon, and Robert J. Douglass, editors. The Characteristics of Parallel Algorithms, volume 30 of MIT Press series in scientific computation. MIT Press, Cambridge, MA, USA, 1987. ISBN 0-262-10036-3. LCCN QA76.6 .C42981 1987.

Jing:1990:PTP

[JR90] J. Jing and John R. Rice. Problems to test parallel and vector languages – II. Technical report TR-1016, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, September 1990. 49 pp. URL http://www.cs.purdue.edu/

research/technical_reports/1990/TR%2090-1016.pdf.

Joshi:1997:ABS

[JRD^{+}97] A. Joshi, N. Ramakrishnan, T. T. Drashansky, Elias N. Houstis, John R. Rice, S. Weerawarana, and L. H. Tsoukalas. Agent based systems to support multidisciplinary problem solving environments. Technical report TR-97-031, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA,

June 1997. 27 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1997/TR%2097-031.pdf.

Joshi:2001:ABN

[JRD^{+}01] A. Joshi, N. Ramakrishnan, T. T. Drashansky, E. N. Houstis, J. R. Rice, S. Weerawarana, and C. H. Tsoukalas. An agent-based netcentric framework for MPSEs. International Journal of Computational Engineering Science (IJCES), ??(??):??, ???? 2001. CODEN ???? ISSN 1465-8763 (print), 2047-6086 (electronic). To appear.

Joshi:1996:NNF

[JRHR96] A. Joshi, N. Ramakrishnan, Elias N. Houstis, and J. R. Rice. On neurobiological, neuro-fuzzy and statistical pattern recognition techniques. Technical report TR-96-004, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107,

USA, January 1996. 23 pp. URL http://www.cs.purdue.edu/
research/technical_reports/1996/TR%2096-004.pdf.

Joshi:1997:NNF

[JRHR97] A. Joshi, N. Ramakrishnan, E. N. Houstis, and J. R. Rice. On neurobiological, neuro-fuzzy and statistical pattern recognition techniques. IEEE Transactions on Neural Networks, 8(1):18–31, ???? 1997. CODEN ITNNEP. ISSN 1045-9227 (print), 1941-0093 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.jsp?
tp=&arnumber=554188.

Joshi:1995:NFA

[JRRH95] A. Joshi, N. Ramakrishnan, John R. Rice, and E. N. Houstis. A neuro-fuzzy approach to agglomerative clustering. Technical report TR-95-066, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, October 1995.

7 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1995/TR%2095-066.pdf.

Joshi:1996:NFA

[JRRH96] Aunupam Joshi, Narendran Ramakrishnan, John R. Rice, and Elias N. Houstis. A neuro-fuzzy approach to agglomerative clustering. In IEEE [IEE96], pages 1028–1033. ISBN 0-7803-3210-5 (softbound), 0-7803-3211-3 (casebound), 0-7803-3212-1 (microfiche). LCCN QA76.87.I3437 1996. URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=549039. Four vol-

umes. IEEE catalog number 96CH35907.

Joshi:1997:NSM

[JWH^{+}97] A. Joshi, S. Weerawarana, Elias N. Houstis, J. R. Rice, T. T. Drashansky, and S. Markus. Network servers for multidisciplinary problem solving. Technical report TR-97-023, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, April 1997. 7 pp. URL http://www.cs.purdue.edu/

research/technical_reports/1997/TR%2097-023.pdf.

Joshi:1995:NNF

[JWR^{+}95] A. Joshi, S. Weerawarana, N. Ramakrishnan, E. N. Houstis, and John R. Rice. Neural and neuro-fuzzy approaches to support “Intelligent” scientific problem solving. Technical report TR-95-039, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, June 1995. 21

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1995/TR%2095-039.pdf.

Joshi:1996:NFS

[JWR^{+}96] Anupam Joshi, Sanjiva Weerawarana, Narendran Ramakrishnan, Elias N. Houstis, and John R. Rice. Neuro-fuzzy support for problem-solving environments: A step toward automated solution of PDEs. IEEE Computational Science & Engineering, 3(1): 44–56, Spring 1996. CODEN ISCEE4. ISSN 1070-9924 (print),

1558-190X (electronic). URL http://www.computer.org/cse/
cs1998/c1044abs.htm.

Kalenich:1965:IPP

[Kal65] Wayne A. Kalenich, editor. Information processing 1965: proceedings of IFIP congress 65; New York City May 24–29, 1965. Spartan Books, New York, NY, USA, 1965. LCCN ???? Two volumes.

Kershbaum:2004:PFV

[KAMR04] F. Kershbaum, M. J. Atallah, D. M’Raihi, and J. R. Rice. Private fingerprint verification without local storage. In Zhang and Jain [ZJ04], page 7. CODEN LNCSD9. ISBN 3-540-22146-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76 A1 L43 3072 (LC). URL http:

//link.springer-ny.com/link/service/series/0558/tocs/
t3072.htm;   http://www.springerlink.com/openurl.asp?

genre=issue&issn=0302-9743&volume=3072;   http://www.
springerlink.com/openurl.asp?genre=volume&id=doi:10.1007/
b98225.

Kant:1992:KPG

[KDM^{+}92] E. Kant, F. Daube, W. Macgregor, J. Wald, E. N. Houstis, J. R. Rice, and R. Vichnevetsky. Knowledge-based program generation for mathematical modeling. In Houstis et al. [HRV92], pages 371– 392. ISBN 0-444-89226-5. LCCN QA76.95 .I48 1990.

Kim:1992:MPS

[KHHR92] S. B. Kim, A. Hadjidimos, Elias N. Houstis, and J. R. Rice. Multi-parameterized Schwarz splittings. Technical report TR-92-073, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, October 1992. 34

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1992/TR%2092-073.pdf.

Kim:1995:MPS

[KHHR95] S-B Kim, A. Hadjidimos, Elias N. Houstis, and J. R. Rice. Multiparameterized Schwarz alternating methods for elliptic boundary value problems. Technical report TR-95-005, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107,

USA, February 1995. 30 pp. URL http://www.cs.purdue.edu/
research/technical_reports/1995/TR%2095-005.pdf.

Kim:1996:MPS

[KHHR96] S.-B. Kim, A. Hadjidimos, E. N. Houstis, and J. R. Rice. Multiparameterized Schwarz alternating methods for elliptic boundary value problems. Mathematics and Computers in Simulation, 42(1):47–76, September 1996. CODEN MCSIDR. ISSN 0378-4754 (print), 1872-7166 (electronic). URL https://www.

sciencedirect.com/science/article/pii/0378475495001115.

Klerer:1968:ISE

[KR68] Melvin Klerer and Juris Reinfelds, editors. Interactive Systems for Experimental Applied Mathematics: proceedings, Washington, DC. Academic Press, New York, NY, USA, 1968.

Kumar:1991:PFI

[Kum91] V. K. Prasanna Kumar, editor. Proceedings, the Fifth International Parallel Processing Symposium, Anaheim, CA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1991. ISBN 0-8186-9167-0. LCCN QA76.5 .I559 1991; QA76.5 .I57 1991eb; QA76.58 .I563 1991; QA76.5. IEEE catalog number 91TH0363-2.

Kling:1993:TCB

[KWRW93] Rob Kling, Peter Wegner, John R. Rice, and Eric A. Weiss. Technical correspondence: Broadening computer science. Communications of the ACM, 36(2):15–20, February 1993. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Keys:1994:DDM

[KX94] David E. Keys and Jinchao Xu, editors. Domain decomposition methods in scientific and engineering computing: proceedings of the Seventh International Conference on Domain Decomposition, October 27–30, 1993, the Pennsylvania State University, volume 180 of Contemporary mathematics. American Mathematical Society, Providence, RI, USA, 1994. ISBN 0-8218-5171-3. ISSN 0271-4132 (print), 1098-3627 (electronic). LCCN QA402.2 .I55 1993.

Lorentz:1973:ATC

[LBCS73] G. G. Lorentz, H. Berens, E. W. Cheney, and L. L. Schumaker, editors. Approximation theory: proceedings of an International Symposium conducted by the University of Texas and the National Science Foundation at Austin, Texas, January 22–24, 1973. Academic Press, New York, NY, USA, 1973. ISBN 0-12-455750-3. LCCN QA221 .A651; QA221 .A651 1973; QA221 .A65.

Leininger:1983:CAD

[Lei83] G. G. Leininger, editor. Computer aided design of multivariable technological systems: proceedings of the second IFAC symposium, West Lafayette, Indiana, USA, 15–17 September 1982, IFAC proceedings series. Pergamon, New York, NY, USA, 1983. ISBN 0-08-029357-3. LCCN TA174 .C5826 1982; TA174 .C5826 1983; TA174 .C5826 1983 06/19/86 CT. US$150.00.

Levin:1998:RAS

[Lev98] Stewart A. Levin. Remark on Algorithm 622: a simple macroprocessor. ACM Transactions on Mathematical Software, 24(3): 336–340, September 1998. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). URL http://www.acm.org:80/

pubs/citations/journals/toms/1998-24-3/p336-levin/. See

[RRW84].

Lai:1992:ISH

[LHHR92] Y-L Lai, A. Hadjidimos, Elias N. Houstis, and J. R. Rice. On the iterative solution of Hermite collocation equations. Technical report TR-92-094, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, December 1992.

44 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1992/TR%2092-094.pdf.

Lai:1994:AIE

[LHHR94a] Y-L Lai, A. Hadjidimos, Elias N. Houstis, and J. R. Rice. The analysis of iterative elliptic PDE solvers based on the cubic Hermite collocation discretization. Technical report TR-94-036, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, May 1994. 83 pp. URL http:/

/www.cs.purdue.edu/research/technical_reports/1994/TR%
2094-036.pdf.

Lai:1994:GIHa

[LHHR94b] Y-L Lai, A. Hadjidimos, Elias N. Houstis, and J. R. Rice. General interior Hermite collocation methods for second order elliptic partial differential equations. Technical report TR-94-004, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, January 1994. 17

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1994/TR%2094-004.pdf.

Lai:1994:GIHb

[LHHR94c] Yu-Ling Lai, Apostolos Hadjidimos, Elias N. Houstis, and John R. Rice. General interior Hermite collocation methods for second-order elliptic partial differential equations. Applied Numerical Mathematics: Transactions of IMACS, 16(1– 2):183–200, December 1, 1994. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). URL http://

www.elsevier.com/cgi-bin/cas/tree/store/apnum/cas_sub/
browse/browse.cgi?year=1994&volume=16&issue=1-2&aid=531.

A Festschrift to honor Professor Robert Vichnevetsky on his 65th birthday.

Lai:1995:ISH

[LHHR95] Yu-Ling Lai, Apostolos Hadjidimos, Elias N. Houstis, and John R. Rice. On the iterative solution of Hermite collocation equations. SIAM Journal on Matrix Analysis and Applications, 16(1):254– 277, January 1995. CODEN SJMAEL. ISSN 0895-4798 (print),

1095-7162 (electronic). URL http://epubs.siam.org/sam-bin/
dbq/article/24230.

Lin:1993:CFW

[LHRL93] Herbert S. Lin, Juris Hartmanis, John R. Rice, and Morton Lowengrub. Computing the future: Whither computer science and engineering? In Buck and Kwasny [BK93], page 530. ISBN 0-89791-558-5, 0-89791-559-3. LCCN QA75.5; QA75.5 .A26

1993; Internet. URL http://www.acm.org/pubs/citations/
proceedings/csc/170791/p530-lin/.

Lynch:1968:CRM

[LR68a] Robert E. Lynch and John R. Rice. Convergence rates of ADI methods with smooth initial error. Mathematics of Computation, 22(??):311–335, 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lynch:1968:CRA

[LR68b] Robert E. Lynch and John R. Rice. Convergence rates of ADI methods with smooth initial error. Mathematics of Computation, 22(102):311–335, April 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lemme:1975:SPA

[LR75a] James M. Lemme and John R. Rice. Speedup in parallel algorithms for adaptive quadratures. Technical report TR-192, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, January 1975. 17

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1975/TR%2075-192.pdf.

Lynch:1975:HAF

[LR75b] Robert E. Lynch and John R. Rice. High accuracy finite difference approximation to solutions of elliptic partial differential equations. Technical report TR-223, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, Jan-

uary 1975. 17 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1975/TR%2075-223.pdf.

Lynch:1975:HOD

[LR75c] Robert E. Lynch and John R. Rice. A high order difference method for differential equation. Technical report TR-244, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, January 1975. 57

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1975/TR%2075-244.pdf.

Lynch:1975:HM

[LR75d] Robert E. Lynch and John R. Rice. The Hodie method. Technical report TR-170, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, January 1975.

13 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1975/TR%2075-170.pdf.

Lynch:1975:HMO

[LR75e] Robert E. Lynch and John R. Rice. The Hodie method for ordinary differential equations. Technical report TR-188, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, January 1975. 45 pp. URL http://www.

cs.purdue.edu/research/technical_reports/1975/TR%2075-
188.pdf.

Lynch:1975:CTI

[LR75f] Robert Emmett Lynch and John Rischard Rice. Computers: Their Impact and Use — with BASIC Programming. Holt, Rinehart, and Winston, New York, NY, USA, 1975. ISBN 0-03-088526-4. xi + 398 pp. LCCN QA76 .L891; QA76 .L89; QA76 .L989c.

Lynch:1977:CTI

[LR77] Robert Emmett Lynch and John Rischard Rice. Computers, Their Impact and Use: Structured Programming in Fortran. Holt, Rinehart, and Winston, New York, NY, USA, January 1977. ISBN 0-03-088525-6. ix + 453 pp. LCCN QA76 .L893 Bar.

US$12.95. URL http://www.cbooks.com/sqlnut/SP/search/
gtsumt?source=&isbn=0030885256. Includes index.

Lynch:1978:HAF

[LR78a] Robert E. Lynch and John R. Rice. High accuracy finite difference approximation to solutions of elliptic partial differential equations. Proceedings of the National Academy of Sciences of the United States of America, 75(6):2541–2544, 1978. CODEN PNASA6. ISSN 0027-8424 (print), 1091-6490 (electronic).

Lynch:1978:HMP

[LR78b] Robert E. Lynch and John R. Rice. The Hodie method and its performance for solving elliptic partial differential equations. In de Boor and Golub [dBG78], pages 143–175. ISBN 0-12-208360-1. LCCN QA3 .U45 no. 41; QA297 S994 1978. URL http://

catalog.hathitrust.org/api/volumes/oclc/65991765.html;
 http://www.gbv.de/dms/hbz/toc/ht001227727.pdf.

Lynch:1978:CTI

[LR78c] Robert Emmett Lynch and John Rischard Rice. Computers, their impact and use: structured programming in PL/1. Holt, Rinehart, and Winston, New York, NY, USA, 1978. ISBN 0-03-088527-2. ix + 442 pp. LCCN QA76.73.P25 L96.

Lemme:1979:SPA

[LR79] James M. Lemme and John R. Rice. Speedup in parallel algorithms for adaptive quadrature. Journal of the ACM, 26(1):65– 71, January 1979. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Lemme:1980:AQA

[LR80a] J. M. Lemme and J. R. Rice. Adaptive quadrature algorithms for the ILLIAC IV. International Journal of Computer and Information Sciences, 9(??):63–72, ???? 1980. CODEN IJCIAH. ISSN 0091-7036.

Lynch:1980:HOD

[LR80b] Robert E. Lynch and John R. Rice. A high-order difference method for differential equations. Mathematics of Computation, 34(150):333–372, April 1980. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Lambers:1991:NQG

[LR91] J. V. Lambers and John R. Rice. Numerical quadrature for general two-dimensional domains. Technical report TR-91-067, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, September 1991. 35

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1991/TR%2091-067.pdf.

Lynch:1964:TPA

[LRT64a] R. E. Lynch, J. R. Rice, and D. H. Thomas. Tensor product analysis of partial difference equations. Bulletin of the American Mathematical Society, 70(??):378–384, 1964. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881x (electronic).

Lynch:1964:DSP

[LRT64b] Robert E. Lynch, John R. Rice, and Donald H. Thomas. Direct solution of partial difference equations by tensor product methods. Numerische Mathematik, 6(??):185–199, 1964. CODEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Lynch:1965:TPA

[LRT65] Robert E. Lynch, John R. Rice, and Donald H. Thomas. Tensor product analysis of alternating direction implicit methods. Journal of the Society for Industrial and Applied Mathematics, 13(4): 995–1006, December 1965. CODEN JSIMAV. ISSN 0368-4245 (print), 1095-712X (electronic).

Law:1976:TAA

[LS76] Alan G. (Alan Greenwell) Law and Badri N. Sahney, editors. Theory of approximation, with applications: proceedings of a conference conducted by the University of Calgary and the University of Regina, at the University of Calgary, Alberta, Canada, August 11–13, 1975. Academic Press, New York, NY, USA, 1976. ISBN 0-12-438950-3. LCCN QA297.5 .C68 1975. Dedicated to the memory of Eckard Schmidt.

Makedon:1994:PDS

[Mak94] F. Makedon, editor. Proceedings of the DAGS’94 Symposium. Dartmouth Institute for Advanced Graduate Studies, Dartmouth, NH, USA, 1994. ISBN ???? LCCN ????

Mason:1987:AAB

[MC87] J. C. Mason and M. G. Cox, editors. Algorithms for approximation: based on the proceedings of the IMA Conference on Algorithms for the Approximation of Functions and Data, held at the Royal Military College of Science, Shrivenham, July 1985, volume 10 of The Institute of Mathematics and Its Applications conference series, new series. Clarendon Press, Oxford, UK, 1987. ISBN 0-19-853612-7. LCCN QA221 .A5361 1987; QA221 .I47 1985. US$90.

Meinardus:1964:AFI

[Mei64] Gunter Meinardus. Approximation von Funktionen und ihre numerische Behandlung. (German) [Approximation of functions and their numerical treatment], volume 4 of Springer tracts in natural philosophy. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1964. ISBN 3-540-03219-3, 3-642-85646-2, 3-642-85647-0 (print). viii + 180 pp. LCCN QA320 .M4.

Meinardus:1967:AFT

[Mei67] Günter Meinardus. Approximation of functions: Theory and numerical methods, volume 13 of Springer Tracts in Natural Philosophy. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1967. ISBN 0-387-03985-6, 3-540-03985-6. ISSN 0081-3877. viii + 198 pp. LCCN QA221 .M3813. URL

http://catalogue.bnf.fr/ark:/12148/cb37378349d.

Macaraeg:1989:IAM

[MH89] M. G. Macaraeg and M. Y. Hussaini, editors. 12th International Association for Mathematics and Computation in Simulation (IMACS) World Congress on Scientific Computation, volume 6(1–2) of Applied Numerical Mathematics: Transactions of IMACS. North-Holland, Amsterdam, The Netherlands, December 1989. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Markus:1999:ABN

[MHC^{+}99] Shahani Markus, Elias N. Houstis, Ann C. Catlin, John R. Rice, P. Tsompanopoulou, Emmanuel Vavalis, Dana Gottfried, and Ke Su. An agent-based netcentric framework for multidisciplinary problem solving environments (MPSE). Technical report TR-99-025, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, August 1999. 36

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1999/TR%2099-025.pdf.

Markus:2000:ABNb

[MHC^{+}00a] S. Markus, E. N. Houstis, Ann C. Catlin, J. R. Rice, P. Tsompanopoulou, E. A. Vavalis, D. Gottfried, K. Su, and G. Balakrishnan. An agent-based netcentric framework for multi disciplinary problem solving environments. International Journal of Computational Engineering Science (IJCES), ??(??):33–60, ???? 2000. CODEN ???? ISSN 1465-8763 (print), 2047-6086 (electronic).

Markus:2000:ABNa

[MHC^{+}00b] S. Markus, Elias N. Houstis, Ann C. Catlin, J. R. Rice, P. Tsompanopoulou, E. Vavalis, D. Gottfried, Ke Su, and G. Balakrishnan. An agent-based netcentric framework for multidisciplinary problem solving environments (MPSE). Technical report TR-00-003, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, February 2000.

27 pp. URL http://www.cs.purdue.edu/research/technical_
reports/2000/TR%2000-003.pdf.

Martin:1995:ALB

[MMR95] I. M. Martin, D. C. Marinescu, and J. R. Rice. Adaptive load balancing strategies for solving irregular problems on distributed memory MIMD systems. In IEEE [IEE95a], pages 57–64. ISBN 0-8186-7074-6. ISSN 1063-7133. LCCN QA76.58 .I58 1995.

Muller:1995:PSI

[MN95] Hausi A. Müller and Ronald J. Norman, editors. Seventh International Workshop on Computer-Aided Software Engineering: July 10-14, 1995, Toronto, Ontario, Canada: proceedings. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7078-9 (softbound), 0-7803-2946-5 (casebound), 0-7803-2962-7 (microfiche). LCCN QA76.758 .I575 1995. IEEE catalog number 95CH35827.

Marinescu:1986:DOA

[MR86] D. C. Marinescu and John R. Rice. Domain oriented analysis of PDE splitting algorithms. Technical report TR-627, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, September 1986. 24 pp. URL http://www.

cs.purdue.edu/research/technical_reports/1986/TR%2086-
627.pdf.

Marinescu:1987:NPC

[MR87a] D. Marinescu and J. Rice. Nonhomogeneous parallel computation I. Synchronization analysis of parallel algorithms. Technical Report TR-683, Department of Computer Science, Purdue University, 1987.

Marinescu:1987:DOA

[MR87b] D. C. Marinescu and J. R. Rice. Domain oriented analysis of PDE splitting algorithms. Journal of Information Science, Principles and Practice, 43(??):3–24, ???? 1987. CODEN JISCDI. ISSN 0165-5515 (print), 1741-6485 (electronic).

Marinescu:1987:NHP

[MR87c] D. C. Marinescu and John R. Rice. Non homogeneous parallel computations I. synchronization analysis of parallel numerical algorithms. Technical report TR-683, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA,

May 1987. 26 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1987/TR%2087-683.pdf.

McFaddin:1987:PVP

[MR87d] H. Scott McFaddin and John R. Rice. Parallel and vector problems on the FLEX/32. Technical report TR-661, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107,

USA, February 1987. 87 pp. URL http://www.cs.purdue.edu/
research/technical_reports/1987/TR%2087-661.pdf.

McFaddin:1987:PIP

[MR87e] H. Scott McFaddin and John R. Rice. PROTRAN II: Preliminary report. Technical report TR-698, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA,

July 1987. 76 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1987/TR%2087-698.pdf.

Marinescu:1988:MHS

[MR88a] D. C. Marinescu and J. R. Rice. Modeling hardware-software interactions in parallel and distributed systems using stochastic high level Petri nets. IEEE Distr. Proc. News, 10(??):28–34, ???? 1988. CODEN ???? ISSN ????

Marinescu:1988:MAI

[MR88b] D. C. Marinescu and John R. Rice. Multilevel asynchronous iteration for PDEs. Technical report TR-820, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107,

USA, November 1988. 20 pp. URL http://www.cs.purdue.edu/
research/technical_reports/1988/TR%2088-820.pdf.

Marinescu:1988:NAL

[MR88c] D. C. Marinescu and John R. Rice. Non-algorithmic load imbalance effects for domain decomposition methods on a hypercube. Technical report TR-832, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, Decem-

ber 1988. 25 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1988/TR%2088-832.pdf.

Marinescu:1988:ESP

[MR88d] D. C. Marinescu and John R. Rice. On the effects of synchronization in parallel computing. Technical report TR-750, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, March 1988. 21 pp. URL http://www.

cs.purdue.edu/research/technical_reports/1988/TR%2088-
750.pdf.

Mu:1988:EPA

[MR88e] Mo Mu and John R. Rice. An experimental performance analysis for the rate of convergence on collocation on general domains. Technical report TR-738, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, January

1988. 51 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1988/TR%2088-738.pdf.

Mu:1988:ROG

[MR88f] Mo Mu and John R. Rice. Row oriented Gauss elimination on distributed memory multiprocessors. Technical report TR-845, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, December 1988. 30

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1988/TR%2088-845.pdf.

Marinescu:1989:ATL

[MR89a] D. C. Marinescu and J. R. Rice. Analysis of a two level asynchronous algorithm for PDEs. In Wright [Wri89], pages 23–33. ISBN 0-444-87310-4. LCCN QA76.5 .I2775 1988.

Marinescu:1989:SNP

[MR89b] Dan C. Marinescu and John R. Rice. Synchronization of nonhomogeneous parallel computations. In Rodrigue [Rod89], pages 363–367. ISBN 0-89871-228-9. LCCN QA76.5 .S515 1987. Also Computer Science Tech. Report CSD-TR-683, Purdue University, May, 1987.

Mu:1989:EPA

[MR89c] Mo Mu and John R. Rice. An experimental performance analysis for the rate of convergence of collocation on general domains.

Numerical Methods for Partial Differential Equations, 5(1):45–52, ???? 1989. CODEN NMPDEB. ISSN 0749-159x (print), 1098-2426 (electronic).

Mu:1989:GBS

[MR89d] Mo Mu and John R. Rice. A grid based subtree-subcube assignment strategy for solving PDEs on hypercubes. Technical report TR-869, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, February 1989.

17 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1989/TR%2089-869.pdf.

Mu:1989:SLS

[MR89e] Mo Mu and John R. Rice. Solving linear systems with sparse matrices on hypercubes. Technical report TR-870, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, February 1989. 16 pp. URL http://www.cs.purdue.

edu/research/technical_reports/1989/TR%2089-870.pdf.

Marinescu:1990:MLA

[MR90a] D. C. Marinescu and J. R. Rice. Multi-level asynchronous iteration for PDEs. In Iterative methods for large linear systems (Austin, TX, 1988), pages 193–213. Academic Press, New York, NY, USA, 1990. ISBN 0-12-407475-8.

Marinescu:1990:ECL

[MR90b] D. C. Marinescu and John R. Rice. The effects of communication latency upon synchronization and dynamic load balance on a hypercube. Technical report TR-1032, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107,

USA, October 1990. 17 pp. URL http://www.cs.purdue.edu/
research/technical_reports/1990/TR%2090-1032.pdf.

Marinescu:1990:SLI

[MR90c] D. C. Marinescu and John R. Rice. Synchronization and load ImBalance effects in distributed memory multiprocessor systems. Technical report TR-1000, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, July 1990.

36 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1990/TR%2090-1000.pdf.

Marinescu:1990:APM

[MR90d] Dan Marinescu and John R. Rice. Application of the E/T performance modeling methodology to a computation on a 128 processor NCUBE. Technical report TR-998, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA,

July 1990. 27 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1990/TR%2090-998.pdf.

Marinescu:1990:HLC

[MR90e] Dan Marinescu and John R. Rice. On high level characterization of parallelism. Technical report TR-1011, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107,

USA, August 1990. 27 pp. URL http://www.cs.purdue.edu/
research/technical_reports/1990/TR%2090-1011.pdf.

Marinescu:1990:SPC

[MR90f] Dan Cristian Marinescu and John R. Rice. On single parameter characterization of parallelism. In Jájá [Jáj90], pages 235–237. ISBN 0-8186-9053-4. LCCN QA76.58 .S95

1990. URL http://ieeexplore.ieee.org/stamp/stamp.jsp?
tp=&arnumber=89464.

Mu:1990:ISG

[MR90g] M. Mu and John R. Rice. Incomplete sparse Gauss factorization for solving PDEs. Technical report TR-991, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, July 1990. 30 pp. URL http://www.cs.purdue.edu/

research/technical_reports/1990/TR%2090-991.pdf.

Mu:1990:PSD

[MR90h] M. Mu and John R. Rice. PARALLEL SPARSE: Data structure and organization. Technical report TR-974, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, April 1990. 24 pp. URL http://www.cs.purdue.

edu/research/technical_reports/1990/TR%2090-974.pdf.

Mu:1990:SPS

[MR90i] M. Mu and John R. Rice. The structure of parallel sparse matrix algorithms for solving partial differential equations on hypercubes. Technical report TR-976, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, April 1990.

42 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1990/TR%2090-976.pdf.

Marinescu:1991:CCS

[MR91a] D. C. Marinescu and J. R. Rice. Communication and control in SPMD parallel numerical computations. In Anonymous [Ano90], pages 37–69. ISBN ???? LCCN ???? ARO Report 91-1.

Marinescu:1991:ECL

[MR91b] D. C. Marinescu and J. R. Rice. The effects of communication and latency on synchronization and dynamic load balance on a hypercube. In Kumar [Kum91], pages 18–25. ISBN 0-8186-9167-0. LCCN QA76.5 .I559 1991; QA76.5 .I57 1991eb; QA76.58 .I563 1991; QA76.5. URL http://ieeexplore.ieee.

org/stamp/stamp.jsp?tp=&arnumber=153751. IEEE catalog

number 91TH0363-2.

Marinescu:1991:SLI

[MR91c] D. C. Marinescu and J. R. Rice. Synchronization and load imbalance effects in distributed memory multi-processor systems. Concurrency: Practice and Experience, 3(6):593–625, December 1991. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

McFaddin:1991:CPS

[MR91d] H. Scott McFaddin and John R. Rice. Collaborating PDE solvers. Technical report TR-91-068, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, Septem-

ber 1991. 24 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1991/TR%2091-068.pdf.

McFaddin:1991:RSP

[MR91e] S. McFaddin and John R. Rice. RELAX: a software platform for PDE interface relaxation methods. Technical report TR-91-018, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, March 1991. 13

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1991/TR%2091-018.pdf.

Mu:1991:OSG

[MR91f] M. Mu and John R. Rice. An organization of sparse Gauss elimination for solving PDEs on distributed memory machines. Technical report TR-91-080, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, November 1991.

20 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1991/TR%2091-080.pdf.

Mu:1991:PPS

[MR91g] M. Mu and John R. Rice. Performance of PDE sparse solvers on hypercubes. Technical report TR-91-002, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107,

USA, January 1991. 23 pp. URL http://www.cs.purdue.edu/
research/technical_reports/1991/TR%2091-002.pdf.

Mu:1991:USP

[MR91h] M. Mu and John R. Rice. Unstructured scheduling in parallel PDE sparse solvers on distributed memory machines. Technical report TR-91-077, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, November 1991.

49 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1991/TR%2091-077.pdf.

Marinescu:1992:SCC

[MR92a] D. C. Marinescu and John R. Rice. Speedup, communication complexity and blocking — à la recherche du temps perdu. Technical report TR-92-057, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, August 1992.

26 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1992/TR%2092-057.pdf.

McFaddin:1992:ARP

[MR92b] H. Scott McFaddin and John R. Rice. Architecture of the RE-LAX problem solving environment. Technical report TR-92-081, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, August 1992. 22 pp. URL http:/

/www.cs.purdue.edu/research/technical_reports/1992/TR%
2092-081.pdf.

McFaddin:1992:CPS

[MR92c] H. Scott McFaddin and John R. Rice. Collaborating PDE solvers. Applied Numerical Mathematics: Transactions of IMACS, 10(??): 279–295, ???? 1992. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

McFaddin:1992:RPS

[MR92d] H. Scott McFaddin and John R. Rice. RELAX: a platform for software relaxation. expert systems for scientific computing. In Houstis et al. [HRV92], pages 175–194. ISBN 0-444-89226-5. LCCN QA76.95 .I48 1990.

Mu:1992:PPS

[MR92e] M. Mu and J. R. Rice. Performance of PDE sparse solvers on hypercubes. In Mehrotra et al. [MSV92], pages 345–370. ISBN 0-262-13272-9. LCCN Q183.9 .U57 1992. Papers presented at a workshop held by ICASE in Nags Head, NC in October 1990.

Mu:1992:ROG

[MR92f] M. Mu and J. R. Rice. Row oriented Gauss elimination on distributed memory multiprocessors. International Journal of High Speed Computing (IJHSC), 4(2):143–168, June 1992. CODEN IH-SCEZ. ISSN 0129-0533.

Mu:1992:PDD

[MR92g] M. Mu and John R. Rice. Preconditioning for domain decomposition through function approximation. Technical report TR-92-091, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, November 1992. 18

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1992/TR%2092-091.pdf.

Mu:1992:GBS

[MR92h] Mo Mu and J. R. Rice. A grid-based subtree-subcube assignment strategy for solving partial differential equations on hypercubes. SIAM Journal on Scientific and Statistical Computing, 13(3):826– 839, May 1992. CODEN SIJCD4. ISSN 0196-5204.

Mu:1992:PMP

[MR92i] Mo Mu and John Rice. Performance of massively parallel distributed memory multiprocessors for a PDE sparse solver benchmark. Technical report TR-92-018, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107,

USA, March 1992. 47 pp. URL http://www.cs.purdue.edu/
research/technical_reports/1992/TR%2092-018.pdf.

Marinescu:1993:SCC

[MR93a] Dan Cristian Marinescu and John R. Rice. Speedup, communication complexity and blocking — à la recherche du temps perdue. In IEEE [IEE93], pages 712–721. ISBN 0-8186-3442-1. ISSN 1070-9924 (print), 1558-190X (electronic). LCCN QA 76.58 I56

1993. URL http://ieeexplore.ieee.org/stamp/stamp.jsp?

tp=&arnumber=262793. IEEE catalog no. 93TH0513-2.

Mu:1993:CPS

[MR93b] Mo Mu and John R. Rice. Collaborating PDE solvers with interface smoothing. Technical report TR-93-024, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, April 1993. 20 pp. URL http://www.cs.purdue.

edu/research/technical_reports/1993/TR%2093-024.pdf.

Mu:1993:OSG

[MR93c] Mo Mu and John R. Rice. An organization of sparse Gauss elimination for solving partial differential equations on distributed memory machines. Numerical Methods for Partial Differential Equations, 9(2):175–189, ???? 1993. CODEN NMPDEB. ISSN 0749-159x (print), 1098-2426 (electronic).

Mu:1993:PCRa

[MR93d] Mo Mu and John R. Rice. Preconditioner construction with rational approximation. Technical report TR-93-001, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, January 1993. 5 pp. URL http://www.cs.purdue.

edu/research/technical_reports/1993/TR%2093-001.pdf.

Mu:1993:PCRb

[MR93e] Mo Mu and John R. Rice. Preconditioner construction with rational approximation. In Sincovec et al. [SKR^{+}93], pages 678–682. ISBN 0-89871-315-3. LCCN QA76.58 .S55 1993 v.1-2. Two volumes.

Marinescu:1994:HLC

[MR94a] Dan C. Marinescu and John R. Rice. On high level characterization of parallelism. Journal of Parallel and Distributed Computing, 20(1):107–113, January 1994. CODEN JPDCER. ISSN 0743-7315 (print), 1096-0848 (electronic). URL http:/

/www.idealibrary.com/links/doi/10.1006/jpdc.1994.1011/
production;   http://www.idealibrary.com/links/doi/10.
1006/jpdc.1994.1011/production/pdf. Also as Tech report

1011, 1990, Purdue University, Department of Computer Science.

Marinescu:1994:SAP

[MR94b] Dan C. Marinescu and John R. Rice. On the scalability of asynchronous parallel computations. Journal of Parallel and Distributed Computing, 22(3):538–546, ???? 1994. CODEN JPD-CER. ISSN 0743-7315 (print), 1096-0848 (electronic).

Mu:1994:MCPa

[MR94c] Mo Mu and John R. Rice. Modeling with collaborating PDE solvers — theory and practice. Technical report TR-94-056, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, August 1994. 16 pp. URL http:/

/www.cs.purdue.edu/research/technical_reports/1994/TR%
2094-056.pdf.

Mu:1994:MCPb

[MR94d] Mo Mu and John R. Rice. Modeling with collaborating PDE solvers: theory and practice. In Keys and Xu [KX94], pages 427–438. ISBN 0-8218-5171-3. ISSN 0271-4132 (print), 1098-3627 (electronic). LCCN QA402.2 .I55 1993.

Mu:1994:PDD

[MR94e] Mo Mu and John R. Rice. Preconditioning for domain decomposition through function approximation. SIAM Journal on Scientific Computing, 15(6):1452–1466, November 1994. CODEN SJOCE3. ISSN 1064-8275 (print), 1095-7197 (electronic).

Mu:1995:MCP

[MR95] M. Mu and J. R. Rice. Modeling with collaborating PDE solvers — theory and practice. Computing Systems in Engineering: an International Journal, 6(??):87–95, ???? 1995. CODEN COSEEO. ISSN 0956-0521 (print), 1873-6211 (electronic).

Mu:1997:SCP

[MR97] M. Mu and John R. Rice. Solving composite problems with interface relaxation. Technical report TR-97-029, Department of Computer Science, Purdue University, West Lafayette, IN 47907-

2107, USA, May 1997. 30 pp. URL http://www.cs.purdue.edu/
research/technical_reports/1997/TR%2097-029.pdf.

Marinescu:1992:MED

[MRCH^{+}92] D. C. Marinescu, John R. Rice, M. A. Cornea-Hasegan, Robert E. Lynch, and M. G. Rossmann. Macromolecular electron density averaging on distributed memory MIMD systems. Technical report TR-92-019, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, March 1992.

28 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1992/TR%2092-019.pdf.

Marinescu:1993:MED

[MRCH^{+}93] D. C. Marinescu, J. R. Rice, M. A. Cornea-Hasegan, R. E. Lynch, and M. G. Rossmann. Macromolecular electron density averaging on distributed memory MIMD systems. Concurrency: Practice and Experience, 5(8):635–657, December 1993. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-9128 (electronic).

Marinescu:1990:PIM

[MRV90a] D. C. Marinescu, John R. Rice, and E. Vavalis. Performance of iteration methods for distributed memory processors. Technical report TR-979, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, May 1990.

27 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1990/TR%2090-979.pdf.

Marinescu:1990:CCS

[MRV90b] D. C. Marinescu, John R. Rice, and E. A. Vavalis. Communication and control in SPMD parallel numerical computations. Technical report TR-981, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, May 1990.

35 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1990/TR%2090-981.pdf.

Marinescu:1991:PIM

[MRV91] D. C. Marinescu, J. R. Rice, and E. A. Vavalis. Performance of iterative methods for distributed memory machines (extended abstract). In Vichnevetsky and Miller [VM91], pages 684–685. ISBN ???? LCCN ???? See full paper in [MRV93].

Marinescu:1993:PIM

[MRV93] D. C. Marinescu, J. R. Rice, and E. A. Vavalis. Performance of iterative methods for distributed memory machines. Applied Numerical Mathematics: Transactions of IMACS, 12(5):421–430, July 1993. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic). See extended abstract in [MRV91].

Marinescu:1990:DS

[MRW^{+}90] Dan Cristian Marinescu, John R. Rice, B. Waltsburger, Catherine E. Houstis, Thomas Kunz, and Helmut Waldschmidt. Distributed supercomputing. In Second IEEE Workshop on Future Trends of Distributed Computing Systems, 1990. Proceedings, pages 381–387. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA,

1990. URL http://ieeexplore.ieee.org/stamp/stamp.jsp?
tp=&arnumber=138350.

Mu:1991:PEO

[MRW91] Mo Mu, John R. Rice, and J. W. Wang. Performance experiments and optimizations of PDE sparse solvers on hypercubes. Technical report TR-91-045, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, June 1991.

22 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1991/TR%2091-045.pdf.

Mehrotra:1992:USC

[MSV92] Piyush Mehrotra, Joel Saltz, and Robert Voigt, editors. Unstructured Scientific Computation on Scalable Microprocessors. MIT Press, Cambridge, MA, USA, 1992. ISBN 0-262-13272-9. LCCN Q183.9 .U57 1992. Papers presented at a workshop held by ICASE in Nags Head, NC in October 1990.

Michopoulos:2003:DAD

[MTH^{+}03] J. Michopoulos, P. Tsompanopoulou, E. N. Houstis, J. R. Rice, C. Farhat, M. Lesoinne, and F. Lechenault. Design architecture of a data driven environment for multiphysics applications. Report DETC 2003/CIE-48268, Amer. Soc. Mech. Engr., 2003. 8 pp.

Markus:1997:SCWa

[MWHR97a] S. Markus, S. Weerawarana, Elias N. Houstis, and J. R. Rice. Scientific computing via the World Wide Web: The Net//ELLPACK PSE server. Technical report TR-97-022, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107,

USA, March 1997. 22 pp. URL http://www.cs.purdue.edu/
research/technical_reports/1997/TR%2097-022.pdf.

Markus:1997:SCWb

[MWHR97b] Shahani Markus, Sanjiva Weerawarana, Elias N. Houstis, and John R. Rice. Scientific computing via the Web: The Net Pellpack PSE server. IEEE Computational Science & Engineering, 4(3):43–51, July/September 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://dlib.

computer.org/cs/books/cs1997/pdf/c3043.pdf;   http://
ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=615430;
 http://www.computer.org/cse/cs1998/c3043abs.htm.

Markus:2000:NBP

[MWHR00] S. Markus, S. Weerawarana, E. N. Houstis, and J. R. Rice. Network based PSEs for PDE computing. In Houstis et al. [HGRB00], chapter 20, pages 249–260. ISBN 0-7923-7809-1. LCCN QA76 .E548 2000.

Nash:1990:HSC

[Nas90] Stephen G. Nash, editor. A History of Scientific Computing. ACM Press history series. Addison-Wesley and ACM Press, Addison-Wesley and New York, NY 10036, USA, 1990. ISBN 0-201-50814-1. xix + 359 pp. LCCN QA76.17 .H59 1990.

Newman:2002:MSC

[NR02] Ken B. Newman and John Rice. Modeling the survival of chinook salmon smolts outmigrating through the lower Sacramento river system. Journal of the American Statistical Association, 97(460):983–??, December 2002. CODEN JSTNAL. ISSN 0162-1459 (print), 1537-274X (electronic).

URL http://fidelio.ingentaselect.com/cgi-bin/linker?
ini=asa&reqidx=/cw/asa/01621459/v97n460/s6/p983.

Ortega:1970:NSN

[OR70] James M. Ortega and Werner C. Rheinboldt, editors. Numerical Solutions of Nonlinear Problems: a collection of papers presented at Symposia in Numerical Solution of Nonlinear problems sponsored by the Office of Naval Research at the Fall meeting of the Society for Industrial and Applied Mathematics held at Philadelphia, Pennsylvania, October 21–23, 1968, volume 2 of Studies in Numerical Analysis. SIAM Press, Philadelphia, PA, USA, 1970. LCCN ????

Rais:1984:CMV

[Raĭ84] Dzh. Raĭs. Matrichnye vychisleniya i matematicheskoe obespechenie. In V. V. Voevodin, editor, ????, page 264. “Mir”, Moscow, USSR, 1984. Translated from the English by O. B. Arushanyan.

Rice:1981:SEP

[RB81] John R. Rice and Ronald F. Boisvert. Solving elliptic problems using ELLPACK. Technical report TR-414, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107,

USA, January 1981. 128 pp. URL http://www.cs.purdue.edu/
research/technical_reports/1981/TR%2081-414.pdf.

Rice:1985:SEP

[RB85] John Rischard Rice and Ronald F. Boisvert, editors. Solving Elliptic Problems Using ELLPACK, volume 2 of Springer Series in Computational Mathematics. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1985. ISBN 0-387-90910-9, 3-540-90910-9. x + 497 pp. LCCN QA377 .R53 1985. With appendices by W. R. Dyksen, E. N. Houstis, Rice, J. F. Brophy, C. J. Ribbens and W. A. Ward.

Rice:1996:SSL

[RB96] John R. Rice and Ronald F. Boisvert. From scientific software libraries to problem-solving environments. IEEE Computational Science & Engineering, 3(3):44–53, Fall 1996. CO-DEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.jsp?
tp=&arnumber=537091;   http://www.computer.org/cse/cs1998/
c3044abs.htm.

Rice:2000:SSL

[RB00] J. R. Rice and R. F. Boisvert. Scalable software libraries and problem solving environments. In Houstis et al. [HGRB00], chapter 3, pages 33–43. ISBN 0-7923-7809-1. LCCN QA76 .E548 2000.

Rice:1968:LEG

[RD68] John R. Rice and Edsger W. Dijkstra. Letter to the Editor: The go to statement reconsidered. Communications of the ACM, 11(8): 538, August 1968. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rice:1981:SVN

[RD81] John R. Rice and Wayne R. Dyksen. Symmetric versus nonsymmetric differencing for self-adjoint elliptic problems. Technical report TR-415, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, January 1981.

5 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1981/TR%2081-415.pdf.

Rice:1994:SCS

[RD94] John Rischard Rice and Richard A. DeMillo, editors. Studies in computer science: in honor of Samuel D. Conte, Software science and engineering. Plenum Press, New York, NY, USA; London, UK, 1994. ISBN 0-306-44697-9. LCCN QA76 .S848 1994.

Rice:1986:EPS

[RDHR86] John R. Rice, Wayne R. Dyksen, Elias N. Houstis, and Calvin J. Ribbens. ELLPACK Project status report. Technical report TR-579, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, February 1986.

32 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1986/TR%2086-579.pdf.

Reid:1982:RBN

[Rei82] J. K. Reid, editor. The Relationship Between Numerical Computation and Programming Languages: Proceedings of the IFIP TC2 Working Conference on the Relationship between Numerical Computation and Programming Languages, Boulder, Colorado, USA, 3–7 August, 1981. Elsevier North-Holland, Inc., New York, NY, USA, 1982. ISBN 0-444-86377-X. LCCN QA297 .I34 1981.

Rice:1979:NCN

[RGO^{+}79] J. R. Rice, C. W. Gear, J. Ortega, B. Parlett, M. Schultz, L. F. Shampine, P. Wolfe, and J. F. Traub. Numerical computation: its nature and research directions. ACM SIGNUM Newsletter, 14(3S (Special issue)):1–48, February 1979. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Rice:1984:ARK

[RH84] John R. Rice and Richard J. Hanson. Algorithm 620: References and keywords for Collected Algorithms of the ACM. ACM Transactions on Mathematical Software, 10(4):359–360, December 1984. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See also [Ham85, HM90].

Rice:1981:PLS

[RHD81] John R. Rice, Elias N. Houstis, and Wayne R. Dyksen. A population of linear, second order, elliptic partial differential equations on rectangular domains. I, II. Mathematics of Computation, 36(154):475–484 (loose microfiche suppl.), April 1981. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic). Includes loose microfiche supplement.

Rice:1959:CAB

[Ric59a] J. R. Rice. On the convergence of an algorithm for best Tchebycheff approximations. Journal of the Society for Industrial and Applied Mathematics, 7(??):133–142, 1959. CODEN JSIMAV. ISSN 0368-4245 (print), 1095-712X (electronic).

Rice:1959:CBN

[Ric59b] John Rischard Rice. The Characterization of Best Nonlinear Tchebycheff Approximations. Ph.D. thesis, California Institute of Technology, Pasadena, CA, USA, January 1959. ????

pp. URL http://resolver.caltech.edu/CaltechETD:etd-
02102006-083607;   http://thesis.library.caltech.edu/
594/.

Rice:1960:CBN

[Ric60a] John R. Rice. The characterization of best nonlinear Tchebycheff approximations. Transactions of the American Mathematical Society, 96(??):322–340, 1960. CODEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (electronic).

Rice:1960:CA

[Ric60b] John R. Rice. Chebyshev approximation by ab^{x} + c. Journal of the Society for Industrial and Applied Mathematics, 8(4):691– 702, December 1960. CODEN JSIMAV. ISSN 0368-4245 (print), 1095-712X (electronic).

Rice:1960:CEE

[Ric60c] John R. Rice. Criteria for the existence and equioscillation of best Tchebycheff approximations. Journal of Research of the National Bureau of Standards. Section B, Mathematics and Mathematical Physics, 64B(??):91–93, 1960. CODEN JNBBAU. ISSN 0022-4340.

Rice:1960:STB

[Ric60d] John R. Rice. Sequence transformations based on Tchebycheff approximations. Journal of Research of the National Bureau of Standards. Section B, Mathematics and Mathematical Physics, 64B (??):227–235, 1960. CODEN JNBBAU. ISSN 0022-4340.

Rice:1960:SRK

[Ric60e] John R. Rice. Split Runge–Kutta method for simultaneous equations. Journal of Research of the National Bureau of Standards. Section B, Mathematics and Mathematical Physics, 64B(??):151– 170, 1960. CODEN JNBBAU. ISSN 0022-4340.

Rice:1961:INT

[Ric61a] J. R. Rice. Introduction to nonlinear Tchebycheff approximation. Industrial Mathematics, 11(??):55–68, 1961. CODEN IMTHAI. ISSN 0019-8528.

Rice:1961:ACA

[Ric61b] John R. Rice. Algorithms for Chebyshev approximation by ab^{x}+c. Journal of the Society for Industrial and Applied Mathematics, 9 (4):571–583, December 1961. CODEN JSIMAV. ISSN 0368-4245 (print), 1095-712X (electronic).

Rice:1961:BAI

[Ric61c] John R. Rice. Best approximations and interpolating functions. Transactions of the American Mathematical Society, 101(??):477– 498, 1961. CODEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (electronic).

Rice:1961:TAF

[Ric61d] John R. Rice. Tchebycheff approximations by functions unisolvent of variable degree. Transactions of the American Mathematical Society, 99(??):298–302, 1961. CODEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (electronic).

Rice:1962:CAE

[Ric62a] John R. Rice. Chebyshev approximation by exponentials. Journal of the Society for Industrial and Applied Mathematics, 10(1):149– 161, March 1962. CODEN JSIMAV. ISSN 0368-4245 (print), 1095-712X (electronic).

Rice:1962:TAC

[Ric62b] John R. Rice. Tchebycheff approximation in a compact metric space. Bulletin of the American Mathematical Society, 68 (4):405–410, ???? 1962. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881x (electronic). URL http://projecteuclid.

org/euclid.bams/1183524686.

Rice:1963:AFD

[Ric63a] J. R. Rice. The applicability of finite differences to lubrication problems. In ????, editor, Proc. 1963 Lubrication Symposium, page ?? American Society of Mechanical Engineering, ????, 1963. LCCN ????

Rice:1963:ACC

[Ric63b] J. R. Rice. Approximation with convex constraints. Journal of the Society for Industrial and Applied Mathematics, 11(??):15– 32, 1963. CODEN JSIMAV. ISSN 0368-4245 (print), 1095-712X (electronic).

Rice:1963:NNS

[Ric63c] J. R. Rice. Note on numerical solutions for gas-lubricated bearings. Transactions of the ASME. Series C, 85(??):187–189, ???? 1963. CODEN ???? ISSN ????

Rice:1963:TAS

[Ric63d] John R. Rice. Tchebycheff approximation in several variables. Transactions of the American Mathematical Society, 109(??):444– 466, 1963. CODEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (electronic).

Rice:1964:TAF

[Ric64a] John R. Rice. The Approximation of Functions, volume Volume 1, Linear Theory. Addison-Wesley, Reading, MA, USA, 1964. 206 pp.

Rice:1964:AFV

[Ric64b] John R. Rice. The approximation of functions. Vol. I: Linear theory, volume 1. Addison-Wesley, Reading, MA, USA, 1964. xi + 203 pp. LCCN QA221 .R5 V.1-2.

Rice:1964:NA

[Ric64c] John R. Rice. On nonlinear L_{1} approximation. Archive for rational mechanics and analysis, 17(??):61–66, 1964. CODEN AVR-MAW. ISSN 0003-9527 (print), 1432-0673 (electronic).

Rice:1964:WA

[Ric64d] John R. Rice. On the L_{∞} Walsh arrays for Γ(x) and Erfc(x). Mathematics of Computation, 18(88):617–626, October 1964. CO-DEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rice:1964:CAE

[Ric64e] John R. Rice. On the computation of L_{1} approximations by exponentials, rationals, and other functions. Mathematics of Computation, 18(87):390–396, July 1964. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rice:1964:ECB

[Ric64f] John R. Rice. On the existence and characterization of best nonlinear Tchebycheff approximations. Transactions of the American Mathematical Society, 110(??):88–97, 1964. CODEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (electronic).

Rice:1965:APA

[Ric65a] J. R. Rice. Automatic procedures for approximating functions and data. In Kalenich [Kal65], pages 572–573. LCCN ???? Two volumes.

Rice:1965:MTH

[Ric65b] J. R. Rice. Misalignment torques of hydrodynamic gas-lubricated journal bearings. Transactions of the ASME. Series D, Journal of basic engineering, 85(??):193–198, ???? 1965. CODEN ???? ISSN 0021-9223.

Rice:1965:BRG

[Ric65c] John R. Rice. Book review: Günter Meinardus, Approximation von Funktionen und ihre numerische Behandlung, Springer-Verlag, Berlin, 1964, viii + 180 pp., 23 cm. Price DM 49. Mathematics of Computation, 19(91):517, July 1965. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic). URL

https://www.jstor.org/stable/2003711.

Rice:1965:NA

[Ric65d] John R. Rice. Nonlinear approximation. In Garabedian [Gar65], pages 111–133. LCCN ????

Rice:1965:CPR

[Ric65e] John R. Rice. On the conditioning of polynomial and rational forms. Numerische Mathematik, 7(5):426–435, October 1965. CO-DEN NUMMA7. ISSN 0029-599X (print), 0945-3245 (electronic).

Rice:1966:EGS

[Ric66a] John R. Rice. Experiments on Gram–Schmidt orthogonalization (in Technical Notes and Short Papers). Mathematics of Computation, 20(94):325–328, April 1966. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rice:1966:TC

[Ric66b] John R. Rice. A theory of condition. SIAM Journal on Numerical Analysis, 3(2):287–310, June 1966. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rice:1967:NAA

[Ric67a] J. R. Rice. Numerical analysis: Approximation or estimation? ACM SIGNUM Newsletter, 2(2):??, ???? 1967. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Rice:1967:BRB

[Ric67b] John R. Rice. Book review: Tchebycheff Systems: With Applications in Analysis and Statistics (Samuel Karlin and William J. Studden). SIAM Review, 9(2):257–258, ???? 1967. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rice:1967:CCA

[Ric67c] John R. Rice. Characterization of Chebyshev approximations by splines. SIAM Journal on Numerical Analysis, 4(4):557–565, December 1967. CODEN SJNAAM. ISSN 0036-1429 (print), 1095-7170 (electronic).

Rice:1967:NAI

[Ric67d] John R. Rice. Nonlinear approximation. II. Curvature in Minkowski geometry and local uniqueness. Transactions of the American Mathematical Society, 128(??):437–459, 1967. CODEN TAMTAM. ISSN 0002-9947 (print), 1088-6850 (electronic).

Rice:1968:AFP

[Ric68a] J. R. Rice. Approximation formulas for physical data. Pyrodynamics, 6(??):231–256, ???? 1968. CODEN ???? ISSN 0555-8344.

Rice:1968:CPA

[Ric68b] J. R. Rice. On the construction of polyalgorithms for automatic numerical analysis. In Klerer and Reinfelds [KR68], pages 301– 313.

Rice:1968:PII

[Ric68c] J. R. Rice. A path independent integral and the approximate analysis of strain concentration by notches and cracks. Journal of Applied Mechanics, 35(2):379–386, June 1968. CODEN JAMCAV. ISSN 0021-8936 (print), 1528-9036 (electronic).

Rice:1969:AFT

[Ric69a] J. R. Rice. Approximation des fonctions. Théorie linéaire. Traduit de l’anglais par M. Saloin. Dunod, Paris, France, 1969. viii + 215 pp. Traduit de l’anglais par M. Saloin.

Rice:1969:PASb

[Ric69b] J. R. Rice. A polyalgorithm for the automatic solution of nonlinear equations. In ACM [ACM69], pages 179–183. LCCN ????

Rice:1969:AFV

[Ric69c] John R. Rice. The approximation of functions. Vol. 2: Nonlinear and multivariate theory, volume 2. Addison-Wesley, Reading, MA, USA, 1969. xiii + 334 pp. LCCN QA221 .R5 V.1-2.

Rice:1969:DCE

[Ric69d] John R. Rice. The degree of convergence for entire functions. Technical report TR-36, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, June 1969.

19 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1969/TR%2069-36.pdf.

Rice:1969:DCN

[Ric69e] John R. Rice. On the degree of convergence of nonlinear spline approximation. In Schoenberg [Sch69], pages 349–365. ISBN 0-12-628850-X. LCCN QA3 .U45 no. 23.

Rice:1969:PASa

[Ric69f] John R. Rice. A polyalgorithm for the automatic solution of nonlinear equations. Technical report TR-32, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107,

USA, February 1969. 11 pp. URL http://www.cs.purdue.edu/
research/technical_reports/1969/TR%2069-32.pdf.

Rice:1969:STF

[Ric69g] John R. Rice. A set of 74 test functions for nonlinear equation solvers. Technical report TR-34, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA,

April 1969. 9 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1969/TR%2069-34.pdf.

Rice:1970:GPC

[Ric70a] J. R. Rice. General purpose curve fitting. In Talbot [Tal70], pages 191–204. LCCN QA221 .A66????

Rice:1970:CEF

[Ric70b] John R. Rice. A characterization of entire functions in terms of degree of convergence. Bulletin of the American Mathematical Society, 76(??):129, 1970. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881x (electronic). URL http://projecteuclid.

org/euclid.bams/1183531406.

Rice:1970:MTN

[Ric70c] John R. Rice. Minimization and techniques in nonlinear approximation. In Ortega and Rheinboldt [OR70], pages 80–98. LCCN ????

Rice:1970:NAF

[Ric70d] John R. Rice. Numerical analysis — final report. Technical report TR-49, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, October 1970. 12

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1970/TR%2070-49.pdf.

Rice:1971:FNA

[Ric71a] J. R. Rice. On the future of numerical analysis. In ????, editor, Proceedings of Share XXXVI, Los Angeles, CA, March 1971, page 8. ????, ????, March 1971.

Rice:1971:SAL

[Ric71b] J. R. Rice. SQUARES: An algorithm for least squares approximations. In Rice [Ric71d], pages 451–476. ISBN 0-12-587250-X. LCCN QA1 .M26. Based on the proceedings of the Mathematical Software Symposium held at Purdue University, Lafayette, Indiana, USA, April 1–3, 1970.

Rice:1971:DCE

[Ric71c] John R. Rice. The degree of convergence for entire functions. Duke Mathematical Journal, 38(3):429–440, ???? 1971. CODEN DUMJAO. ISSN 0012-7094 (print), 1547-7398 (electronic). URL

http://projecteuclid.org/euclid.dmj/1077379786.

Rice:1971:MS

[Ric71d] John R. Rice, editor. Mathematical Software. Academic Press, New York, NY, USA, 1971. ISBN 0-12-587250-X. LCCN QA1 .M26. Based on the proceedings of the Mathematical Software Symposium held at Purdue University, Lafayette, Indiana, USA, April 1–3, 1970.

Rice:1971:MRN

[Ric71e] John R. Rice. Matrix representations of nonlinear equation iterations—Application to parallel computation. Mathematics of Computation, 25(116):639–647, October 1971. CODEN MCM-PAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rice:1971:RO

[Ric71f] John R. Rice. Running orthogonalization. Journal of Approximation Theory, 4(??):332–338, 1971. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Rice:1972:PFS

[Ric72] John R. Rice. On the present and future of scientific computation. Communications of the ACM, 15(7):637–639, July 1972. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rice:1973:NLS

[Ric73a] J. R. Rice. NAPSS-like systems: Problems and prospects. In AFIPS [AFI73], pages 43–47. LCCN ????

Rice:1973:EAQa

[Ric73b] John R. Rice. An educational adaptive quadrature algorithm. Technical report TR-90, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, Febru-

ary 1973. 15 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1973/TR%2073-90.pdf.

Rice:1973:EAQb

[Ric73c] John R. Rice. An educational adaptive quadrature algorithm. Technical report TR-89, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, Febru-

ary 1973. 15 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1973/TR%2073-89.pdf.

Rice:1973:EAQc

[Ric73d] John R. Rice. An educational adaptive quadrature algorithm. ACM SIGNUM Newsletter, 8(2):27–41, April 1973. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic). See errata [Rob73].

Rice:1973:CCA

[Ric73e] John R. Rice. On the computational complexity of approximation operators. In Lorentz et al. [LBCS73], pages 449–455. ISBN 0-12-455750-3. LCCN QA221 .A651; QA221 .A651 1973; QA221 .A65.

Rice:1973:PAAb

[Ric73f] John R. Rice. Parallel algorithms for adaptive quadrature — convergence. Technical report TR-104, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA,

September 1973. 19 pp. URL http://www.cs.purdue.edu/
research/technical_reports/1973/TR%2073-104.pdf.

Rice:1973:PAAa

[Ric73g] John R. Rice. Parallel algorithms for adaptive quadrature II — metalgorithm correctness. Technical report TR-107, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, November 1973. 29 pp. URL http://www.

cs.purdue.edu/research/technical_reports/1973/TR%2073-
107.pdf.

Rice:1974:BCC

[Ric74a] J. R. Rice. B74-1 complexity of computer computations. IEEE Transactions on Computers, C-23(1):109–110, January 1974. CODEN ITCOB4. ISSN 0018-9340 (print), 1557-9956 (elec-

tronic). URL http://ieeexplore.ieee.org/stamp/stamp.jsp?
tp=&arnumber=1672384.

Rice:1974:AA

[Ric74b] John R. Rice. Adaptive approximation. Technical report TR-120, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, June 1974. 10

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1974/TR%2074-120.pdf.

Rice:1974:AQC

[Ric74c] John R. Rice. Adaptive quadrature: convergence of parallel and sequential algorithms. Bulletin of the American Mathematical Society, 80(6):1250–1254, ???? 1974. CODEN BAMOAD. ISSN 0002-9904 (print), 1936-881x (electronic). URL http://

projecteuclid.org/euclid.bams/1183536040.

Rice:1974:ASPc

[Ric74d] John R. Rice. The algorithm selection problem abstract models. Technical report TR-116, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, May 1974.

20 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1974/TR%2074-116.pdf.

Rice:1974:ASPb

[Ric74e] John R. Rice. The algorithm selection problem II two concrete problems: Numerical analysis quadrature algorithms operating systems — scheduling design. Technical report TR-117, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, May 1974. 20 pp. URL http://www.

cs.purdue.edu/research/technical_reports/1974/TR%2074-
117.pdf.

Rice:1974:ASPa

[Ric74f] John R. Rice. The algorithm selection problem III — approximation theory machinery. Technical report TR-130, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, December 1974. 41 pp. URL http://www.cs.purdue.

edu/research/technical_reports/1974/TR%2074-130.pdf.

Rice:1974:BRB

[Ric74g] John R. Rice. Book review: Complexity of Computer Computations (Raymond E. Miller and James W. Thatcher, eds.). SIAM Review, 16(3):407–409, ???? 1974. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rice:1974:PAAb

[Ric74h] John R. Rice. Parallel algorithms for adaptive quadrature — convergence. In Rosenfeld [Ros74], pages 600–604. ISBN 0-444-10689-8. LCCN QA 76 I615.

Rice:1974:PAAa

[Ric74i] John R. Rice. Parallel algorithms for adaptive quadrature III — program correctness. Technical report TR-112, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, March 1974. 51 pp. URL http://www.cs.purdue.

edu/research/technical_reports/1974/TR%2074-112.pdf.

Rice:1975:AAA

[Ric75a] John R. Rice. Algorithm — adapt, adaptive smooth curve fitting. Technical report TR-166, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, Jan-

uary 1975. 15 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1975/TR%2075-166.pdf.

Rice:1975:ASP

[Ric75b] John R. Rice. The algorithm selection problem. Technical report TR-152, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, January 1975. 78

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1975/TR%2075-152.pdf.

Rice:1975:APS

[Ric75c] John R. Rice. Algorithmic progress in solving partial differential equations. Technical report TR-173, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107,

USA, January 1975. 21 pp. URL http://www.cs.purdue.edu/
research/technical_reports/1975/TR%2075-173.pdf.

Rice:1975:ECE

[Ric75d] John R. Rice. ELLPACK–A cooperative effort for the study of numerical methods for elliptic partial differential equations. Technical report TR-203, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, January 1975.

6 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1975/TR%2075-203.pdf.

Rice:1975:MAQ

[Ric75e] John R. Rice. A metalgorithm for adaptive quadrature. Journal of the ACM, 22(1):61–82, January 1975. CODEN JACOAH. ISSN 0004-5411 (print), 1557-735X (electronic).

Rice:1975:NAF

[Ric75f] John R. Rice. Nonlinear approximation — final report for NSF Grant GP-3294OX. Technical report TR-235, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, January 1975. 8 pp. URL http://www.cs.purdue.

edu/research/technical_reports/1975/TR%2075-235.pdf.

Rice:1975:NSD

[Ric75g] John R. Rice. Note on the software design for an elliptic PDE solver. Technical report TR-193, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107,

USA, January 1975. 12 pp. URL http://www.cs.purdue.edu/
research/technical_reports/1975/TR%2075-193.pdf.

Rice:1975:NCB

[Ric75h] John R. Rice. Numerical computation breakthroughs and future challenges. Technical report TR-183, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107,

USA, January 1975. 7 pp. URL http://www.cs.purdue.edu/
research/technical_reports/1975/TR%2075-183.pdf.

Rice:1975:CCA

[Ric75i] John R. Rice. On the computational complexity of approximation operators II. Technical report TR-136, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107,

USA, January 1975. 12 pp. URL http://www.cs.purdue.edu/
research/technical_reports/1975/TR%2075-136.pdf.

Rice:1975:PAA

[Ric75j] John R. Rice. Parallel algorithms for adaptive quadrature. II. Metalgorithm correctness. Acta Informatica, 5(4):273–285, December 19, 1975. CODEN AINFA2. ISSN 0001-5903 (print), 1432-0525 (electronic).

Rice:1975:PS

[Ric75k] John R. Rice. Purpose and scope. ACM Transactions on Mathematical Software, 1(1):1–3, March 1975. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Rice:1975:RPP

[Ric75l] John R. Rice. Remarks on piecewise polynomial approximation. Technical report TR-163, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, Jan-

uary 1975. 19 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1975/TR%2075-163.pdf.

Rice:1975:SNC

[Ric75m] John R. Rice. Software for numerical computation. Technical report TR-214, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, January 1975.

30 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1975/TR%2075-214.pdf.

Rice:1975:SPP

[Ric75n] John R. Rice. Software package policy. ACM Transactions on Mathematical Software, 1(3):193–195, September 1975. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Rice:1975:SDO

[Ric75o] John R. Rice. Some data and observation on research publication (A) numerical computation, (B) programming languages and systems. Technical report TR-197, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, Jan-

uary 1975. 8 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1975/TR%2075-197.pdf.

Rice:1975:SCV

[Ric75p] John R. Rice. Statistical computing — the vanguard of the future of education. Technical report TR-186, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107,

USA, January 1975. 3 pp. URL http://www.cs.purdue.edu/
research/technical_reports/1975/TR%2075-186.pdf.

Rice:1975:SES

[Ric75q] John R. Rice. A study of the effect of structured programming on the execution time of a particular program. Technical report TR-147, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, January 1975.

20 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1975/TR%2075-147.pdf.

Rice:1976:PEA

[Ric76a] J. R. Rice. Programming effort analysis of the ELLPACK language. ACM SIGNUM Newsletter, 14(??):109–111, ???? 1976. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Rice:1976:SDO

[Ric76b] J. R. Rice. Some data and observations on research publication in the areas of numerical computation and programming languages and systems. ACM SIGNUM Newsletter, 11(3):28–32, October 1976. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Rice:1976:SCV

[Ric76c] J. R. Rice. Statistical computing: The vanguard of the revolution in education. In Hoaglin and Welsch [HW76], pages 1–4. ISBN 0-87150-237-2. LCCN QA276.A1 I53 1976.

Rice:1976:AA

[Ric76d] John R. Rice. Adaptive approximation. Journal of Approximation Theory, 16(4):329–337, 1976. CODEN JAXTAZ. ISSN 0021-9045 (print), 1096-0430 (electronic).

Rice:1976:ASP

[Ric76e] John R. Rice. The algorithm selection problem. In Rubinoff and Yovits [RY76], pages 65–118. ISBN 0-12-012115-8. ISSN 0065-2458. LCCN QA76.

Rice:1976:APS

[Ric76f] John R. Rice. Algorithmic progress in solving partial differential equations. ACM SIGNUM Newsletter, 11(4):6–10, December 1976. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Rice:1976:APP

[Ric76g] John R. Rice. On adaptive piecewise polynomial approximation. In Law and Sahney [LS76], pages 359–386. ISBN 0-12-438950-3. LCCN QA297.5 .C68 1975. Dedicated to the memory of Eckard Schmidt.

Rice:1976:CCA

[Ric76h] John R. Rice. On the computational complexity of approximation operators. II. In Traub [Tra76], pages 191–204. ISBN 0-12-697560-4. LCCN QA297 .S915 1975.

Rice:1976:PAA

[Ric76i] John R. Rice. Parallel algorithms for adaptive quadrature. III. Program correctness. ACM Transactions on Mathematical Software, 2(1):1–30, March 1976. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Rice:1976:TPS

[Ric76j] John R. Rice. TOMS policy statement: The rights of program authors in the evaluation of programs. ACM Transactions on Mathematical Software, 2(2):113–114, June 1976. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Rice:1977:ERT

[Ric77a] J. R. Rice. ELLPACK: a research tool for elliptic partial differential equations software. In Rice [Ric77c], pages 319–342. ISBN 0-12-587260-7. LCCN QA3 .U45 no. 39; QA297 .M36 1977.

URL https://www.sciencedirect.com/book/9780125872607/
mathematical-software.

Rice:1977:RPP

[Ric77b] J. R. Rice. Remarks on piecewise polynomial approximation. In Stechkin [Ste77], pages 305–310. ISBN 0-8218-3038-4. LCCN QA1 .A413 no. 138; QA221; QA3 .A462; QA221 .P7413 1977.

Rice:1977:MSI

[Ric77c] John R. Rice, editor. Mathematical software III: Proceedings of a symposium conducted by the Mathematics Research Center, the University of Wisconsin–Madison, March 28–30, 1977, number 39 in Publication of the Mathematics Research Center, the University of Wisconsin, Madison. Academic Press, New York, NY, USA, 1977. ISBN 0-12-587260-7. LCCN QA3 .U45 no.

39; QA297 .M36 1977. URL https://www.sciencedirect.com/
book/9780125872607/mathematical-software.

Rice:1978:EWP

[Ric78a] John R. Rice. 1979 ELLPACK workshop progress report and a proposal for a 2-year program. Technical report TR-315, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, November 1978. 17

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1978/TR%2078-315.pdf.

Rice:1978:AMA

[Ric78b] John R. Rice. Adaptive multivariate approximation theory and applications. Technical report TR-293, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107,

USA, November 1978. 20 pp. URL http://www.cs.purdue.edu/
research/technical_reports/1978/TR%2078-293.pdf.

Rice:1978:AAA

[Ric78c] John R. Rice. Algorithm 525: ADAPT, adaptive smooth curve fitting [E2]. ACM Transactions on Mathematical Software, 4(1): 82–94, March 1978. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Rice:1978:ECG

[Ric78d] John R. Rice. ELLPACK 77 contributor’s guide. Technical report TR-267, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, May 1978. 35

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1978/TR%2078-267.pdf.

Rice:1978:EUGb

[Ric78e] John R. Rice. ELLPACK 77 user’s guide. Technical report TR-289, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, August 1978. 76

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1978/TR%2078-289.pdf.

Rice:1978:EUGa

[Ric78f] John R. Rice. ELLPACK ’78 user’s guide-preliminary version. Technical report TR-306, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, Novem-

ber 1978. 20 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1978/TR%2078-306.pdf.

Rice:1978:EW

[Ric78g] John R. Rice. ELLPACK workshop. Technical report TR-285, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, August 1978. 15 pp. URL http:/

/www.cs.purdue.edu/research/technical_reports/1978/TR%
2078-285.pdf.

Rice:1978:MPP

[Ric78h] John R. Rice. Multivariate piecewise polynomial approximation. In Handscomb [Han78], pages 261–277. ISBN 0-12-323350-X. LCCN QA297.5 .M84.

Rice:1978:PEA

[Ric78i] John R. Rice. Programming effort, analysis of the ELLPACK language. Technical report TR-288, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107,

USA, August 1978. 4 pp. URL http://www.cs.purdue.edu/
research/technical_reports/1978/TR%2078-288.pdf.

Rice:1979:MAS

[Ric79a] J. R. Rice. Methodology for the algorithm selection problem. In Fosdick [Fos79], pages 301–307. ISBN 0-444-85330-8. LCCN QA297 .I18 1978.

Rice:1979:PEAa

[Ric79b] J. R. Rice. Programming effort analysis of the ELLPACK language. In Vichnevetsky and Stepleman [VS79], page 28. ISBN ???? LCCN ????

Rice:1979:SNC

[Ric79c] J. R. Rice. Software for numerical computation. In Wegner [Weg79], pages 688–708. ISBN 0-262-23096-8. LCCN QA76.6 .R45.

Rice:1979:PEAb

[Ric79d] John R. Rice. “Programming effort” analysis of the ELLPACK language. ACM SIGNUM Newsletter, ??(??):109–111, ???? 1979. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Rice:1980:EUG

[Ric80a] John R. Rice. ELLPACK user’s guide supplement. Technical report TR-363, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, January 1980.

13 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1980/TR%2080-363.pdf.

Rice:1980:MSL

[Ric80b] John R. Rice. Is mathematical software a legitimate research area? Technical report TR-362, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, Jan-

uary 1980. 11 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1980/TR%2080-362.pdf.

Rice:1980:MCE

[Ric80c] John R. Rice. Machine and compiler effects on the performance of elliptic PDE software. Technical report TR-359, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, January 1980. 9 pp. URL http://www.cs.purdue.

edu/research/technical_reports/1980/TR%2080-359.pdf.

Rice:1980:EIGa

[Ric80d] John R. Rice. On the effectiveness of iteration for the Galerkin method equations. Technical report TR-362A, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, January 1980. 11 pp. URL http://www.cs.purdue.

edu/research/technical_reports/1980/TR%2080-362A.pdf.

Rice:1980:EIGb

[Ric80e] John R. Rice. On the effectiveness of iteration for the Galerkin method equations. Technical report TR-353, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107,

USA, January 1980. 13 pp. URL http://www.cs.purdue.edu/
research/technical_reports/1980/TR%2080-353.pdf.

Rice:1981:ATD

[Ric81a] John R. Rice. ALGORITHM: a two dimensional domain processor. Technical report TR-417, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, Jan-

uary 1981. 15 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1981/TR%2081-417.pdf.

Rice:1981:AFP

[Ric81b] John R. Rice. Array facilities in programming languages. Technical report TR-380, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, January 1981.

30 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1981/TR%2081-380.pdf.

Rice:1981:MSL

[Ric81c] John R. Rice. Is mathematical software a legitimate research area? ACM SIGNUM Newsletter, 16(1):23–25, March 1981. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Rice:1981:NCG

[Ric81d] John R. Rice. Numerical computation with general two dimensional domains. Technical report TR-416, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107,

USA, January 1981. 11 pp. URL http://www.cs.purdue.edu/
research/technical_reports/1981/TR%2081-416.pdf.

Rice:1981:PAM

[Ric81e] John R. Rice. Performance analysis of 13 methods to solve the Galerkin method equations. Technical report TR-369, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, January 1981. 19 pp. URL http://www.

cs.purdue.edu/research/technical_reports/1981/TR%2081-
369.pdf.

Rice:1981:SMPa

[Ric81f] John R. Rice. A simple macro processor — user’s guide. Technical report TR-403, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, January 1981.

21 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1981/TR%2081-403.pdf.

Rice:1981:SPL

[Ric81g] John R. Rice. Survey on programming language facilities for numerical computation. ACM SIGNUM Newsletter, 16(3):14–20, September 1981. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Rice:1981:MCM

[Ric81h] John Rischard Rice. Matrix Computations and Mathematical Software. McGraw-Hill Computer Science Series. McGraw-Hill, New York, NY, USA, 1981. ISBN 0-07-052145-X. xii + 248 pp. LCCN QA188 .R52.

Rice:1982:PLP

[Ric82a] J. R. Rice. Programming languages: Power, trends, and facilities for numerical computation. In Reid [Rei82], pages 3–16. ISBN 0-444-86377-X. LCCN QA297 .I34 1981.

Rice:1982:SSC

[Ric82b] J. R. Rice. Software for scientific computation. In Leininger [Lei83], pages 339–344. ISBN 0-08-029357-3. LCCN TA174 .C5826 1982; TA174 .C5826 1983; TA174 .C5826 1983 06/19/86 CT. US$150.00.

Rice:1983:MS

[Ric83a] J. R. Rice. Mathematical software. In Ralston and Reilly, Jr. [RR83], pages 925–927. ISBN 0-442-24496-7. LCCN QA76.15 .E48 1983.

Rice:1983:EPE

[Ric83b] John R. Rice. Expansion of the performance evaluation capabilities of ELLPACK. Technical report TR-451, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107,

USA, January 1983. 8 pp. URL http://www.cs.purdue.edu/
research/technical_reports/1983/TR%2083-451.pdf.

Rice:1983:MCE

[Ric83c] John R. Rice. Machine and compiler effects on the performance of elliptic PDE software. In Stepleman et al. [S^{+}83], pages 97–101. ISBN 0-444-86607-8. LCCN Q172 .I46 1982.

Rice:1983:NMSa

[Ric83d] John R. Rice. Numerical Methods, Software, and Analysis. Mc-Graw-Hill, New York, NY, USA, 1983. ISBN 0-07-052208-1. xii + 483 pp. LCCN QA297 .R49 1983. US$19.95.

Rice:1983:NMSb

[Ric83e] John R. Rice. Numerical Methods, Software, and Analysis. McGraw-Hill International, Tokyo, Japan, 1983. ISBN 0-07-066507-9, 0-07-052208-1. 483 pp.

Rice:1983:PWV

[Ric83f] John R. Rice. PARVEC workshop on very large least squares problems and supercomputers. Technical report TR-464, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, January 1983. 22 pp. URL http://www.

cs.purdue.edu/research/technical_reports/1983/TR%2083-
464.pdf.

Rice:1983:PAM

[Ric83g] John R. Rice. Performance analysis of 13 methods to solve the Galerkin method equations. Linear Algebra and its Applications, 52/53(??):533–546, ???? 1983. CODEN LAAPAW. ISSN 0024-3795 (print), 1873-1856 (electronic).

Rice:1983:SPE

[Ric83h] John R. Rice. Software parts for elliptic PPE software. Technical report TR-448, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, January 1983.

13 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1983/TR%2083-448.pdf.

Rice:1983:NMSc

[Ric83i] John Rischard Rice. Numerical methods, software, and analysis. McGraw-Hill, New York, NY, USA, IMSL reference edition edition, 1983. ISBN 0-07-052209-X. x + 661 pp. LCCN QA297 .R49 1983b.

Rice:1984:VLL

[Ric84a] J. R. Rice. Very large least squares problems and supercomputers. In Hwang [Hwa84], pages 404–419. ISBN 0-8186-0581-2. LCCN TK 7888.3 H82 1984.

Rice:1984:ATD

[Ric84b] John R. Rice. Algorithm 625: a two-dimensional domain processor. ACM Transactions on Mathematical Software, 10(4):453– 462, December 1984. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Rice:1984:BLA

[Ric84c] John R. Rice. BLAS, linear, algebra modules and supercomputers parallel workshop #4. Technical report TR-501, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, November 1984. 16 pp. URL http://www.

cs.purdue.edu/research/technical_reports/1984/TR%2084-
501.pdf.

Rice:1984:BEP

[Ric84d] John R. Rice. Building elliptic problem solvers with ELLPACK. In Birkhoff and Schoenstadt [BS84], pages 3–22. ISBN 0-12-100560-7. LCCN QA377 .E533 1983. Proceedings of the Elliptic Problem Solvers Conference, sponsored by the Mathematics and Mechanics Branch, Office of Naval Research, held in Monterey, California, January 10–12, 1983.

Rice:1984:EEP

[Ric84e] John R. Rice. ELLPACK: An evolving problem solving environment. Technical report TR-497, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, Octo-

ber 1984. 9 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1984/TR%2084-497.pdf.

Rice:1984:FEP

[Ric84f] John R. Rice. Fortran extensions for parallel and vector computation. Technical report TR-470, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, Jan-

uary 1984. 8 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1984/TR%2084-470.pdf.

Rice:1984:NCG

[Ric84g] John R. Rice. Numerical computation with general twodimensional domains. ACM Transactions on Mathematical Software, 10(4):443–452, December 1984. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (electronic).

Rice:1984:SPE

[Ric84h] John R. Rice. Software parts for elliptic PDE software. In Engquist et al. [ESY84], pages 123–138. ISBN 0-444-87620-0. LCCN QA377 .I44 1983; QA377.I44. With a discussion.

Rice:1985:ATP

[Ric85a] John R. Rice. Adaptive tensor product grids for singular problems. Technical report TR-536, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, Septem-

ber 1985. 12 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1985/TR%2085-536.pdf.

Rice:1985:ARS

[Ric85b] John R. Rice. The aspect ratio significant for finite element problem. Technical report TR-535, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, Septem-

ber 1985. 14 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1985/TR%2085-535.pdf.

Rice:1985:PTP

[Ric85c] John R. Rice. Problems to test parallel and vector languages. Technical report TR-516, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, May 1985.

95 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1985/TR%2085-516.pdf.

Rice:1985:UST

[Ric85d] John R. Rice. Using supercomputers today and tomorrow. Technical report TR-529, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, August 1985.

12 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1985/TR%2085-529.pdf.

Rice:1986:PSPb

[Ric86a] J. R. Rice. Parallelism in solving PDEs. In Stone and Winkler [SW86], pages 540–546. ISBN 0-8186-4743-4 (microfiche), 0-8186-8743-6 (hardcover), 0-8186-0743-2 (paperback). LCCN QA75.5 .F35 1986. IEEE catalog number 86CH2345-7. Computer Society order number 743. ACM order number 401860.

Rice:1986:DTP

[Ric86b] John R. Rice. Design of a tensor product population of PDE problems. Technical report TR-628, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, Septem-

ber 1986. 12 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1986/TR%2086-628.pdf.

Rice:1986:LIP

[Ric86c] John R. Rice. Language independent PROTRAN. Technical report TR-633, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, October 1986. 11

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1986/TR%2086-633.pdf.

Rice:1986:MFM

[Ric86d] John R. Rice. Multi-flex machines preliminary report. Technical report TR-612, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, July 1986.

20 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1986/TR%2086-612.pdf.

Rice:1986:PMP

[Ric86e] John R. Rice. Parallel methods for PDES. Technical report TR-587, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, March 1986. 28

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1986/TR%2086-587.pdf.

Rice:1986:PSPa

[Ric86f] John R. Rice. Parallelism in solving PDEs. Technical report TR-604, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, June 1986. 21

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1986/TR%2086-604.pdf.

Rice:1987:ATP

[Ric87a] J. R. Rice. Adaptive tensor product grids for singular problems. In Mason and Cox [MC87], pages 493–506. ISBN 0-19-853612-7. LCCN QA221 .A5361 1987; QA221 .I47 1985. US$90.

Rice:1987:EEP

[Ric87b] J. R. Rice. ELLPACK: An evolving problem solving environment. In Ford and Chaitin-Chatelin [FCC87], pages 233–245. ISBN 0-444-70254-7. LCCN Q183.9 .I35 1985.

Rice:1987:MSA

[Ric87c] J. R. Rice. Mathematical software and ACM publications. In Crane [Cra87], pages https://doi.org/10.1145/41579.4158459–62. ISBN 0-89791-229-2. LCCN QA76 .A25 1987.

Rice:1987:MAS

[Ric87d] John R. Rice. Mathematical aspects of scientific software. Technical report TR-713, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, September 1987.

41 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1987/TR%2087-713.pdf.

Rice:1987:PMP

[Ric87e] John R. Rice. Parallel methods for partial differential equations. In Jamieson et al. [JGD87], pages 209–231. ISBN 0-262-10036-3. LCCN QA76.6 .C42981 1987.

Rice:1987:SAPa

[Ric87f] John R. Rice. Supercomputing about physical objects. Technical report TR-708, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, September 1987.

16 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1987/TR%2087-708.pdf.

Rice:1988:SAP

[Ric88a] J. R. Rice. Supercomputing about physical objects. In Houstis et al. [HPP88], pages 443–455. CODEN LNCSD9. ISBN 0-387-18991-2 (USA). ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA267.A1 L43 no.297. URL http:

//link.springer-ny.com/link/service/series/0558/tocs/
t0297.htm;   http://www.springerlink.com/openurl.asp?
genre=issue&issn=0302-9743&volume=297. The conference was

organized and sponsored by the Computer Technology Institute (C.T.I.) of Greece.

Rice:1988:MSD

[Ric88b] John A. Rice. Mathematical Statistics and Data Analysis. Cole Statistics / Probability Series. Wadsworth and Brooks, ????, 1988. Good intermediary statistics book.

Rice:1988:MASb

[Ric88c] John R. Rice. Mathematical aspects of scientific software. In Mathematical Aspects of Scientific Software [Ric88d], pages 1–39. ISBN 3-540-96706-0. LCCN QA76.76.D47 M366 1988.

Rice:1988:MASa

[Ric88d] John R. Rice, editor. Mathematical Aspects of Scientific Software, volume 14 of The IMA volumes in mathematics and its applications. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 1988. ISBN 3-540-96706-0. xi + 208 pp. LCCN QA76.76.D47 M366 1988.

Rice:1988:PED

[Ric88e] John R. Rice. Potential evaluation directions for ELLPACK. Technical report TR-752, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, March 1988.

20 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1988/TR%2088-752.pdf.

Rice:1989:CLS

[Ric89] John R. Rice. Composition of libraries, software parts and problem solving environments. Technical report TR-852, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, January 1989. 15 pp. URL http://www.

cs.purdue.edu/research/technical_reports/1989/TR%2089-
852.pdf.

Rice:1990:CRI

[Ric90a] J. R. Rice. Is computing research an isolated science? Computing Research News: the News Journal of the Computing Research Association, 2(2):1, ???? 1990. CODEN ???? ISSN 1069-384X.

Rice:1990:MSA

[Ric90b] J. R. Rice. Mathematical software and ACM publications. In Nash [Nas90], pages 217–227. ISBN 0-201-50814-1. LCCN QA76.17 .H59 1990.

Rice:1990:SPE

[Ric90c] John R. Rice. Software performance evaluation papers in TOMS, volumes 1–15. Technical report TR-1026, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107,

USA, October 1990. 62 pp. URL http://www.cs.purdue.edu/
research/technical_reports/1990/TR%2090-1026.pdf.

Rice:1991:APC

[Ric91a] J. R. Rice. Academic programs in computational engineering and science available. Computing Research News: the News Journal of the Computing Research Association, 3(1):11–12, ???? 1991. CO-DEN ???? ISSN 1069-384X. See also expanded version [Ric93a].

Rice:1991:PCP

[Ric91b] J. R. Rice. Pros and cons of postdocs in CS. Computing Research News: the News Journal of the Computing Research Association, 3(??):10, ???? 1991. CODEN ???? ISSN 1069-384X.

Rice:1991:FCS

[Ric91c] John R. Rice. And the first computer science department was? Communications of the ACM, 34(2):19, February 1991. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rice:1991:LEF

[Ric91d] John R. Rice. Learning: An experiment with five selection forms. Technical report TR-91-043, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA,

May 1991. 16 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1991/TR%2091-043.pdf.

Rice:1991:LTO

[Ric91e] John R. Rice. Learning, teaching, optimization and approximation. Technical report TR-91-032, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA,

April 1991. 34 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1991/TR%2091-032.pdf.

Rice:1991:PEN

[Ric91f] John R. Rice. Performance evaluation of numerical software. ACM SIGNUM Newsletter, 26(1):7–??, January 1991. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Rice:1992:FRD

[Ric92a] J. R. Rice. Future research directions in problem solving environments for computational science (extended abstract). In Gaffney and Houstis [GH92], pages 363–366. ISBN 0-444-89176-5. LCCN QA76.6 .I1782 1992.

Rice:1992:GBRb

[Ric92b] J. R. Rice. Geometry based representations for mechanical design: Creation of optimization shells. In Houstis et al. [HRV92], pages 203–207. ISBN 0-444-89226-5. LCCN QA76.95 .I48 1990. See also extended abstract [Ric92c].

Rice:1992:GBRa

[Ric92c] J. R. Rice. Geometry based representations for mechanical design: Creation of optimization shells (extended abstract). In Brezinski and Kulisch [BK92], pages 1018–1019. ISBN 0-444-89701-1. LCCN QA76.95 .C62 1991.

Rice:1992:LTO

[Ric92d] J. R. Rice. Learning, teaching, optimization and approximation. In Houstis et al. [HRV92], pages 89–123. ISBN 0-444-89226-5. LCCN QA76.95 .I48 1990.

Rice:1992:NMS

[Ric92e] John R. Rice. Numerical Methods, Software, and Analysis. Academic Press, San Diego, CA, USA, second edition, 1992. ISBN 0-12-587755-2. ???? pp. LCCN QA297 .R49 1993. URL http://

www.loc.gov/catdir/description/els032/93114990.html;
http://www.loc.gov/catdir/toc/els032/93114990.html.

Rice:1993:APCb

[Ric93a] J. R. Rice. Academic programs in computational engineering and science available. Computer, 26(10):11–12, October 1993. CODEN CPTRB4. ISSN 0018-9162 (print), 1558-0814 (electronic).

Rice:1993:MS

[Ric93b] J. R. Rice. Mathematical software. In Ralston and Reilly, Jr. [RR93], pages 826–828. ISBN 0-442-27679-6. LCCN QA76.15 .E48 1993.

Rice:1993:LTP

[Ric93c] John R. Rice. The 1981, 1986, and 1990 long term plans of the Department of Computer Sciences. Technical report TR-93-036, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, June 1993. 68

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1993/TR%2093-036.pdf.

Rice:1993:APCa

[Ric93d] John R. Rice. Academic programs in computational science and engineering. Technical report TR-93-042, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107,

USA, July 1993. 19 pp. URL http://www.cs.purdue.edu/
research/technical_reports/1993/TR%2093-042.pdf.

Rice:1993:PMD

[Ric93e] John R. Rice. Policy manual: Department of Computer Sciences. Technical report TR-93-044, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA,

July 1993. 38 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1993/TR%2093-044.pdf.

Rice:1993:NMS

[Ric93f] John Rischard Rice. Numerical methods, software, and analysis. Academic Press, New York, NY, USA, second edition, 1993. ISBN 0-12-587755-2. xiv + 720 pp. LCCN QA297 .R49 1993.

Rice:1994:CSE

[Ric94a] J. R. Rice. Computational science and engineering at Purdue. In Ames [Ame94], pages 424–427. ISBN ???? LCCN ???? Three volumes.

Rice:1994:APC

[Ric94b] John R. Rice. Academic programs in computational science and engineering. IEEE Computational Science & Engineering, 1(1): 13–21, Spring 1994. CODEN ISCEE4. ISSN 1070-9924 (print),

1558-190X (electronic). URL http://ieeexplore.ieee.org/
stamp/stamp.jsp?tp=&arnumber=295373.

Rice:1994:EPW

[Ric94c] John R. Rice. Editor’s perspectives: Whither computational science and engineering. IEEE Computational Science & Engineering, 1(2):96, Summer 1994. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Rice:1994:FBW

[Ric94d] John R. Rice. Foreword [on the 65th birthday of Walter Gautschi]. In Zahar [Zah94], pages ix–x. ISBN 0-8176-3753-2. LCCN QA221 .A634 1994.

Rice:1994:MEA

[Ric94e] John R. Rice. Metrics to evaluate academic departments. Technical report TR-94-048, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, July 1994.

24 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1994/TR%2094-048.pdf.

Rice:1994:PPI

[Ric94f] John R. Rice. Processing PDE interface conditions. Technical report TR-94-041, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, June 1994.

29 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1994/TR%2094-041.pdf.

Rice:1994:RSE

[Ric94g] John R. Rice. Recognition in science and engineering. Technical report TR-94-015, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, March 1994.

7 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1994/TR%2094-015.pdf.

Rice:1995:PSE

[Ric95a] John Rice. Problem solving environments for scientific computing. In IFIP Working Group 2.5 [IFI95], page ?? ISBN

???? LCCN ???? URL http://www.nsc.liu.se/~boein/ifip/
kyoto/workshop-info/proceedings/rice/rice1.html.

Rice:1995:CSFa

[Ric95b] John R. Rice. Computational science and the future of computing research. Technical report TR-95-045, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA,

July 1995. 11 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1995/TR%2095-045.pdf.

Rice:1995:CSFb

[Ric95c] John R. Rice. Computational science and the future of computing research (linked to full text). IEEE Computational Science & Engineering, 2(4):35–41, Winter 1995. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic). URL http://ieeexplore.

ieee.org/stamp/stamp.jsp?tp=&arnumber=476366;   http:/
/www.computer.org/cse/cs1998/c4035abs.htm.

Rice:1996:PSE

[Ric96a] John Rice. Problem solving environments — Web site. ACM SIGNUM Newsletter, 31(4):35–??, October 1996. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic).

Rice:1996:CSOa

[Ric96b] John R. Rice. Computational science as one driving force for all aspects of computing research. Technical report TR-96-026, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, May 1996. 10 pp. URL http://www.

cs.purdue.edu/research/technical_reports/1996/TR%2096-
026.pdf.

Rice:1996:CSOb

[Ric96c] John R. Rice. Computational science as one driving force for all aspects of computing research. ACM Computing Surveys, 28 (4es):24, December 1996. CODEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (electronic). URL http://www.acm.org/pubs/

citations/journals/surveys/1996-28-4es/a24-rice/.

Rice:1996:CWA

[Ric96d] John R. Rice. Conferences & workshops: Announcing the first IEEE-CS Workshop on Computational Science and Engineering, Purdue University, Indiana, October 21–26, 1996. IEEE Computational Science & Engineering, 3(1):88, Spring 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Rice:1996:CWF

[Ric96e] John R. Rice. Conferences & workshops: First IEEE Computer Society Workshop on CSE. IEEE Computational Science & Engineering, 3(4):86, Winter 1996. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-190X (electronic).

Rice:1996:MPP

[Ric96f] John R. Rice. Measuring the performance of parallel computation (session overview). In Siegel [Sie96], pages 8–9. ISBN 0-8186-7623-X. LCCN QA76.58 .I29 1996; QA76.58; QA76.58 .I327 1996; QA76.58 .I34 1996; QA76.58 .I34 1996eb. URL http://

ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=539063.

Rice:1996:SSS

[Ric96g] John R. Rice. Scalable scientific software libraries and problem solving environments. Technical report TR-96-001, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, January 1996. 25 pp. URL http://www.cs.purdue.

edu/research/technical_reports/1996/TR%2096-001.pdf.

Rice:1996:WAE

[Ric96h] John R. Rice. What is an answer? An essay on the theory of functions. Technical report TR-96-035, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA,

June 1996. 20 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1996/TR%2096-035.pdf.

Rice:1997:FCS

[Ric97a] J. R. Rice. Future challenges for scientific simulation. In Sydow [Syd97], pages 1–8. ISBN 3-89685-550-6 (set), 3-89685-551-4 (vol. 1), 3-89685-552-2 (vol. 2), 3-89685-553-0 (vol. 3),3-89685-554-9 (vol. 4), 3-89685-555-7 (vol. 5), 3-89685-556-5 (vol. 6). LCCN Q183.9 .I46 1997. Also published in [HGRB00].

Rice:1997:WAE

[Ric97b] J. R. Rice. What is an answer? An essay on the theory of functions. In Hermis’96, pages 89–107. LEA Publications, Athens, Greece, 1997. ISBN ???? LCCN ???? Also in [Ric00d].

Rice:1997:CCC

[Ric97c] John R. Rice. The CSE community converges: First IEEE computer society workshop on computational science and engineering. IEEE Computational Science & Engineering, 4(2):10–11, April/June 1997. CODEN ISCEE4. ISSN 1070-9924 (print),

1558-190X (electronic). URL http://dlib.computer.org/cs/
books/cs1997/pdf/c2010.pdf;   http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=609826;   http://www.
computer.org/cse/cs1998/c2010abs.htm.

Rice:1997:FSSa

[Ric97d] John R. Rice. Future scientific software systems. Technical report TR-97-007, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, January 1997. 7

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1997/TR%2097-007.pdf.

Rice:1997:FSSb

[Ric97e] John R. Rice. Future scientific software systems. IEEE Computational Science & Engineering, 4(2):44–48, April/June 1997. CODEN ISCEE4. ISSN 1070-9924 (print), 1558-

190X (electronic).  URL http://dlib.computer.org/cs/
books/cs1997/pdf/c2044.pdf;   http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=609831;   http://www.
computer.org/cse/cs1998/c2044abs.htm.

Rice:1997:SPU

[Ric97f] John R. Rice. Solving PDEs using an agent based architecture. Technical report TR-97-043, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, Septem-

ber 1997. 12 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1997/TR%2097-043.pdf.

Rice:1998:AAS

[Ric98] John Rice. An agent architecture for solving partial differential equation. SIAM News, 31(6):1, 6–7, July/August 1998. CODEN ???? ISSN 0036-1437.

Rice:1999:CKA

[Ric99a] John R. Rice. Can KIVA accurately simulate a gas turbine combustor? Technical report TR-99-008, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107,

USA, March 1999. 10 pp. URL http://www.cs.purdue.edu/
research/technical_reports/1999/TR%2099-008.pdf.

Rice:1999:PCS

[Ric99b] John R. Rice. A perspective on computational science in the 21st Century. Computing in Science and Engineering, 1(2):14–16, March/April 1999. CODEN CSENFA. ISSN 1521-9615 (print),

1558-366X (electronic). URL http://dlib.computer.org/cs/
books/cs1999/pdf/c2014.pdf;   http://ieeexplore.ieee.
org/stamp/stamp.jsp?tp=&arnumber=753042.

Rice:2000:FSS

[Ric00a] J. R. Rice. Future scientific software systems (abstract). Bulletin of the American Physical Society, 45(??):339, ???? 2000. CODEN BAPSA6. ISSN 0003-0503.

Rice:2000:MS

[Ric00b] J. R. Rice. Mathematical software. In Encyclopedia of computer science [RRH00], pages 1093–1096. ISBN 0-333-77879-0, 1-56159-248-X (Grove’s Dictionaries). LCCN QA76.15 .E48 2000. Previous ed.: Anthony Ralston, Edwin D. Reilly, Caryl Ann Dahlin. London: Chapman and Hall, 1993.

Rice:2000:PSE

[Ric00c] J. R. Rice. Problem solving environments. In Encyclopedia of computer science [RRH00], pages 1440–1441. ISBN 0-333-77879-0, 1-56159-248-X (Grove’s Dictionaries). LCCN QA76.15 .E48 2000. Previous ed.: Anthony Ralston, Edwin D. Reilly, Caryl Ann Dahlin. London: Chapman and Hall, 1993.

Rice:2000:WAE

[Ric00d] John R. Rice. What is an answer? An essay on the theory of functions. HERMIS-μπ, 1(??):73–91, ???? 2000. ISSN 1108-7609.

Rice:2002:CC

[Ric02] J. R. Rice. CS&E in the 21st Century. In Shigeo Kawata, Y. Tago, and Y. Umetani, editors, Computational Science, pages 16–19. Nikkei Science, Tokyo, Japan, March 2002. ISBN ???? LCCN ???? In Japanese.

Rice:2003:CGU

[Ric03] J. R. Rice. ET: Come and gone unobserved? Contact in context, ??(??):4, March 2003. CODEN ???? ISSN 1547-8890. URL http:

//cic.setileague.org.

Rice:20xx:BPN

[Ricxx] J. R. Rice. Biographical and professional notes. Mathematics of Computation, ??(??):??, ???? 20xx. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Ramakrishnan:1997:CES

[RJHR97a] N. Ramakrishnan, A. Joshi, Elias N. Houstis, and J. R. Rice. Collaborative environments for scientific computing the task of algorithm/software selection. Technical report TR-97-057, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, December 1997. 7

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1997/TR%2097-057.pdf.

Ramakrishnan:1997:NFA

[RJHR97b] Narendran Ramakrishnan, Aunupam Joshi, Elias N. Houstis, and John R. Rice. Neuro-fuzzy approaches to collaborative scientific computing. In IEEE [IEE97], pages 473–478. ISBN 0-7803-4122-8, 0-7803-4123-6, 0-7803-4124-4. LCCN QA76.87 .I3437 1997; QA76.87; QA76.87 .I3437 1997eb; Internet; QA76.87 .I347

1997. URL http://ieeexplore.ieee.org/stamp/stamp.jsp?

tp=&arnumber=611714. Four volumes. IEEE catalog number: 97CH36109.

Ramakrishnan:1995:UNF

[RJW^{+}95a] N. Ramakrishnan, A. Joshi, S. Weerawarana, E. N. Houstis, and J. R. Rice. Using neuro-fuzzy methods in scientific computing. In Anonymous [Ano95], pages 279–284. ISBN ???? LCCN ????

Ramakrishnan:1995:NFS

[RJW^{+}95b] N. Ramakrishnan, A. Joshi, S. Weerawarana, E. N. Houstis, and John R. Rice. Neuro-fuzzy systems for intelligent scientific computation. Technical report TR-95-026, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA,

April 1995. 17 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1995/TR%2095-026.pdf.

Rice:1974:MSP

[RLC^{+}74] John R. Rice, Charles L. Lawson, William J. Cody, C. William Gear, Hans J. Oser, and Barry W. Boehm. A panel session: Mathematical software: patterns for the future. In American Federation of Information Processing Societies: 1974 National Computer Conference, 6–10 May 1974, Chicago, Illinois, USA, volume 43 of AFIPS Conference Proceedings, page 971. AFIPS Press, Montvale, NJ, USA, 1974.

Rice:2000:MML

[RLR00] John R. Rice, M. Lujan, and P. Rao. MOL: Method of lines application. Technical report TR-00-014, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107,

USA, October 2000. 27 pp. URL http://www.cs.purdue.edu/
research/technical_reports/2000/TR%2000-014.pdf.

Rice:1988:ATL

[RM88a] John R. Rice and D. C. Marinescu. Analysis of a two level asynchronous algorithm for PDEs. Technical report TR-800, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, August 1988. 16

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1988/TR%2088-800.pdf.

Rice:1988:EPA

[RM88b] John R. Rice and Mu. An experimental performance analysis for the rate of convergence of 5-point star on general domains. Technical report TR-747, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, March 1988.

34 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1988/TR%2088-747.pdf.

Roberts:1973:E

[Rob73] O. P. Roberts. Errata: SIGNUM Newsletter 8(2) 32 April (1973). ACM SIGNUM Newsletter, 8(3):11, July 1973. CODEN SNEWD6. ISSN 0163-5778 (print), 1558-0237 (electronic). See [Ric73d].

Rodrigue:1989:PPS

[Rod89] G. Rodrigue, editor. Parallel processing for scientific computing. SIAM Press, Philadelphia, PA, USA, 1989. ISBN 0-89871-228-9. LCCN QA76.5 .S515 1987.

Rosenfeld:1974:IPP

[Ros74] Jack L. Rosenfeld, editor. Information processing, 1974; Proceedings of IFIP Congress 74, Stockholm, Sweden, August 5–10, 1974. North-Holland, Amsterdam, The Netherlands, 1974. ISBN 0-444-10689-8. LCCN QA 76 I615.

Rice:1966:NNAa

[RR66a] J. R. Rice and S. Rosen. NAPSS — a numerical analysis problem solving system. In ACM [ACM66], pages 51–56. LCCN ????

Rice:1966:NNAb

[RR66b] John R. Rice and Saul Rosen. NAPSS—a Numerical Analysis Problem-Solving System. Communications of the ACM, 9(7):476, July 1966. CODEN CACMA2. ISSN 0001-0782 (print), 1557-7317 (electronic).

Rice:1969:ICS

[RR69] John K. Rice and John R. Rice. Introduction to computer science; problems, algorithms, languages, information and computers. Holt, Rinehart, and Winston, New York, NY, USA, 1969. xv + 463 pp. LCCN QA76.5 .R472. Algol and Fortran appendices by Edouard Desautels.

Rice:1973:ICFa

[RR73a] John K. Rice and John Rischard Rice. Introduction to computing with BASIC. Holt, Rinehart, and Winston, New York, NY, USA, 1973. ISBN 0-03-086306-7. ix + 226 pp. LCCN QA76.73.F25 R51.

Rice:1973:ICFb

[RR73b] John K. Rice and John Rischard Rice. Introduction to computing with Fortran. Holt, Rinehart, and Winston, New York, NY, USA, 1973. ISBN 0-03-086217-5. ix + 384 pp. LCCN QA76.73.F25 R51.

Rice:1980:NSF

[RR80] J. R. Rice and J. W. Rudnicki. A note on some features of the theory of localization of deformation. International Journal of Solids and Structures, 16(7):597–605, 1980. CODEN IJSOAD. ISSN 0020-7683 (print), 1879-2146 (electronic).

Rice:1982:BEB

[RR82] J. Rice and M. Rosenblatt. Boundary effects on the behavior of smoothing splines. In Statistics and probability: essays in honor of C. R. Rao, pages 635–643. North-Holland, Amsterdam, The Netherlands, 1982. ISBN 0-444-86130-0.

Ralston:1983:ECS

[RR83] Anthony Ralston and Edwin D. Reilly, Jr., editors. Encyclopedia of Computer Science and Engineering. Van Nostrand Reinhold, New York, NY, USA, second edition, 1983. ISBN 0-442-24496-7. xxix + 1664 pp. LCCN QA76.15 .E48 1983.

Ribbens:1986:RPS

[RR86] Calvin J. Ribbens and John R. Rice. Realistic PDE solutions for non-rectangular domains. Technical report TR-639, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, November 1986. 36

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1986/TR%2086-639.pdf.

Rice:1990:HCS

[RR90a] John R. Rice and Saul Rosen. History of the Computer Science Department at Purdue University. Technical report TR-1003, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, August 1990. 26

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1990/TR%2090-1003.pdf.

Rosen:1990:OCC

[RR90b] Saul Rosen and John R. Rice. The origins of computing and computer science at Purdue University. Technical report TR-1004, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, August 1990. 13

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1990/TR%2090-1004.pdf.

Ralston:1993:ECS

[RR93] Anthony Ralston and Edwin D. Reilly, Jr., editors. Encyclopedia of Computer Science and Engineering. Van Nostrand Reinhold, New York, NY, USA, third edition, 1993. ISBN 0-442-27679-6. xxv + 1558 pp. LCCN QA76.15 .E48 1993.

Rice:1994:HCS

[RR94a] J. R. Rice and S. Rosen. History of the Computer Science Department of Purdue University. In Rice and DeMillo [RD94], pages 45–72. ISBN 0-306-44697-9. LCCN QA76 .S848 1994.

Rosen:1994:OCP

[RR94b] S. Rosen and J. R. Rice. The origins of computing at Purdue University. In Rice and DeMillo [RD94], pages 31–44. ISBN 0-306-44697-9. LCCN QA76 .S848 1994.

Ramakrishnan:1996:GAA

[RR96a] N. Ramakrishnan and John R. Rice. GAUSS: An automatic algorithm selection system for quadrature. Technical report TR-96-048, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, August 1996. 14

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1996/TR%2096-048.pdf.

Ramakrishnan:1996:KDC

[RR96b] N. Ramakrishnan and John R. Rice. Knowledge discovery in computational science: a case study in algorithm selection. Technical report TR-96-081, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, December 1996.

21 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1996/TR%2096-081.pdf.

Rice:1999:HGF

[RR99] John R. Rice and N. Ramakrishnan. How to get a free lunch (at no cost). Technical report TR-99-014, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA,

April 1999. 8 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1999/TR%2099-014.pdf.

Ramakrishnan:2002:GOA

[RR02a] N. Ramakrishnan and J. R. Rice. GAUSS: An online algorithm recommender system for one-dimensional numerical quadrature. Advances in Engineering Software, 23(1):27–36, ???? 2002. CO-DEN AESODT. ISSN 0141-1195, 0965-9978.

Rice:2002:CSP

[RR02b] John R. Rice and Saul Rosen. Computer sciences at Purdue University 1962 to 2000. Technical report TR-02-027, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, December 2002. 18 pp. URL http://www.cs.purdue.

edu/research/technical_reports/2002/TR%2002-027.pdf.

Rice:2004:CSP

[RR04] John R. Rice and Saul Rosen. Computer sciences at Purdue University—1962 to 2000. IEEE Annals of the History of Computing, 26(2):48–61, April/June 2004. CODEN IAHCEX. ISSN 1058-6180 (print), 1934-1547 (electronic). URL http://

csdl.computer.org/comp/mags/an/2004/02/a2048abs.htm;
http://csdl.computer.org/dl/mags/an/2004/02/a2048.htm;
 http://csdl.computer.org/dl/mags/an/2004/02/a2048.pdf;
 http://ieeexplore.ieee.org/stamp/stamp.jsp?tp=&arnumber=
1299659.

Reilly:2000:ECS

[RRH00] Edwin D. Reilly, Anthony Ralston, and David Hemmendinger. Encyclopedia of computer science. Nature Publishing Group, London, UK, fourth edition, 2000. ISBN 0-333-77879-0, 1-56159-248-X (Grove’s Dictionaries). xxix + 2034 + [16] pp. LCCN QA76.15 .E48 2000. Previous ed.: Anthony Ralston, Edwin D. Reilly, Caryl Ann Dahlin. London: Chapman and Hall, 1993.

Rice:1984:ASM

[RRW84] John R. Rice, Calvin Ribbens, and William A. Ward. Algorithm 622: a simple macroprocessor. ACM Transactions on Mathematical Software, 10(4):410–416, December 1984. CODEN ACM-SCU. ISSN 0098-3500 (print), 1557-7295 (electronic). See remark [Lev98].

Rice:1983:IISb

[RS83a] J. R. Rice and H. D. Schwetman. Interface issues in a software parts technology. In Biggerstaff [Big83], pages 129–137. ISBN ???? LCCN ???? Reprinted in [RS87, RS89].

Rice:1983:IISa

[RS83b] John R. Rice and H. D. Schwetman. Interface issues in a software parts technology. Technical report TR-447, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107,

USA, January 1983. 9 pp. URL http://www.cs.purdue.edu/
research/technical_reports/1983/TR%2083-447.pdf.

Rice:1987:IIS

[RS87] J. R. Rice and H. D. Schwetman. Interface issues in a software parts technology. In Freeman [Fre87], pages 96–104. ISBN 0-8186-0750-5 (paperback), 0-8186-4750-7 (microfiche). LCCN QA76.76.R47 T87 1987. Reprint of [RS83a]. Also reprinted in [RS89].

Rice:1989:IIS

[RS89] J. R. Rice and H. D. Schwetman. Interface issues in a software parts technology. In Software reusability [BP89], pages 125– 139. ISBN 0-201-08017-6 (vol. 1), 0-201-50018-3 (vol. 2). LCCN QA76.76.R47 B543 1989; QA76.76.R47 S64 1989. Reprint of [RS83a]. Also reprinted in [RS87].

Rice:1997:RPI

[RTV97] John R. Rice, P. Tsompanopoulou, and E. Vavalis. Review and performance of interface relaxation methods for elliptic differential equations. Technical report TR-97-004, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107,

USA, January 1997. 31 pp. URL http://www.cs.purdue.edu/
research/technical_reports/1997/TR%2097-004.pdf.

Rice:1998:AER

[RTV98a] John R. Rice, P. Tsompanopoulou, and E. Vavalis. Automated estimation of relaxation for interface relaxation. Technical report TR-98-018, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, June 1998. 8

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1998/TR%2098-018.pdf.

Rice:1998:FTI

[RTV98b] John R. Rice, P. Tsompanopoulou, and E. Vavalis. Fine tuning interface relaxation methods for elliptic differential equations. Technical report TR-98-017, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, June 1998.

23 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1998/TR%2098-017.pdf.

Rice:1998:STU

[RTV98c] John R. Rice, P. Tsompanopoulou, and E. Vavalis. SciAgents tool: User’s guide. Technical report TR-98-043, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107,

USA, November 1998. 18 pp. URL http://www.cs.purdue.edu/
research/technical_reports/1998/TR%2098-043.pdf.

Rice:1999:AER

[RTV99] John R. Rice, P. Tsompanopoulou, and E. Vavalis. Automated estimation of relaxation parameters for interface relaxation. Technical report TR-99-015, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, April 1999.

8 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1999/TR%2099-015.pdf.

Rice:2000:IRM

[RTV00a] J. R. Rice, P. Tsompanopoulou, and E. Vavalis. Interface relaxation methods for elliptic differential equations. Applied Numerical Mathematics: Transactions of IMACS, 32(2):219–245, February 2000. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460

(electronic). URL http://www.elsevier.nl/gej-ng/10/10/28/
59/27/32/abstract.html;   http://www.elsevier.nl/gej-
ng/10/10/28/59/27/32/article.pdf.

Rice:2000:FTI

[RTV00b] John R. Rice, T. Tsompanopoulou, and E. Vavalis. Fine tuning interface relaxation methods for elliptic differential. Technical report TR-00-007, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, March 2000.

23 pp. URL http://www.cs.purdue.edu/research/technical_
reports/2000/TR%2000-007.pdf.

Rice:2002:FTI

[RTV02] J. R. Rice, P. Tsompanopoulou, and E. Vavalis. Fine tuning interface relaxation methods for elliptic differential equations. Applied Numerical Mathematics: Transactions of IMACS, 43(4):459–481, December 2002. CODEN ANMAEL. ISSN 0168-9274 (print), 1873-5460 (electronic).

Rice:1968:LAE

[RU68] John R. Rice and Karl H. Usow. The Lawson algorithm and extensions. Mathematics of Computation, 22(101):118–127, January 1968. CODEN MCMPAF. ISSN 0025-5718 (print), 1088-6842 (electronic).

Rice:1998:CAM

[RV98] J. R. Rice and E. A. Vavalis. Collaborating agents for modeling air pollution. Systems analysis, modelling, simulation, 32(??):93–101, ???? 1998. CODEN SAMSEC. ISSN 0232-9298.

Rice:1990:SIC

[RVH90] John R. Rice, R. Vichnevetsky, and Elias N. Houstis. Second international conference on expert systems for numerical computing. Technical report TR-963, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, March 1990.

85 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1990/TR%2090-963.pdf.

Rice:1993:TIC

[RVH93] John R. Rice, R. Vichnevetsky, and Elias N. Houstis. The third international conference on expert systems for numerical computing. Technical report TR-93-028, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA,

May 1993. 118 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1993/TR%2093-028.pdf.

Rice:1993:CAN

[RVY93] John R. Rice, E. A. Vavalis, and D. Yang. Convergence analysis of a non-overlapping domain decomposition method for elliptic PDEs. Technical report TR-93-048, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA,

July 1993. 8 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1993/TR%2093-048.pdf.

Rice:1997:AND

[RVY97] J. R. Rice, E. A. Vavalis, and Daoqi Yang. Analysis of a nonoverlapping domain decomposition method for elliptic partial differential equations. Journal of Computational and Applied Mathematics, 87(1):11–19, December 18, 1997. CO-DEN JCAMDI. ISSN 0377-0427 (print), 1879-1778 (electronic).

URL http://www.sciencedirect.com/science/article/pii/
S0377042797001726.

Rice:1964:NSE

[RW64] John R. Rice and John S. White. Norms for smoothing and estimation. SIAM Review, 6(3):243–256, ???? 1964. CODEN SIREAD. ISSN 0036-1445 (print), 1095-7200 (electronic).

Rice:1981:SMPb

[RW81] John R. Rice and William A. Ward. A simple macro processor. Technical report TR-400, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, Jan-

uary 1981. 8 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1981/TR%2081-400.pdf.

Rubinoff:1976:AC

[RY76] Morris Rubinoff and Marshall C. Yovits, editors. Advances in Computers, volume 15. Academic Press, New York, NY, USA, 1976. ISBN 0-12-012115-8. ISSN 0065-2458. xiii + 301 pp. LCCN QA76.

Stepleman:1983:SCA

[S^{+}83] R. S. (Robert S.) Stepleman et al., editors. Scientific computing: applications of mathematics and computing to the physical sciences, volume 1 of IMACS transactions on scientific computation. North-Holland, Amsterdam, The Netherlands, 1983. ISBN 0-444-86607-8. LCCN Q172 .I46 1982.

Schoenberg:1969:ASE

[Sch69] I. J. Schoenberg, editor. Approximations, with special emphasis on spline functions; proceedings of a symposium conducted by the Mathematics Research Center, United States Army, at the University of Wisconsin, Madison, May 5–7, 1969, Publication no. 23 of the Mathematics Research Center, United States Army, the University of Wisconsin. Academic Press, New York, NY, USA, 1969. ISBN 0-12-628850-X. LCCN QA3 .U45 no. 23.

Sameh:1996:CSE

[SCK^{+}96] A. Sameh, G. Cybenko, M. Kalos, K. Neves, J. Rice, D. Sorensen, and F. Sullivan. Computational science and engineering. ACM Computing Surveys, 28(4):810–817, December 1996. CO-DEN CMSVAN. ISSN 0360-0300 (print), 1557-7341 (elec-

tronic). URL http://www.acm.org/pubs/citations/journals/
surveys/1996-28-4/p810-sameh/.

Snyder:1980:PQ

[SCRS80] Lawrence Snyder, Douglas E. Comer, John R. Rice, and H. D. Schwetman. Project QUANTA. Technical report TR-366, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, January 1980. 72

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1980/TR%2080-366.pdf.

Sopka:1990:ICS

[SdG90] John R. Sopka and D. de Groot, editors. 1990 International Conference on Supercomputing: June 11–15, 1990, Amsterdam, The Netherlands: conference proceedings, volume 18(3) of ACM SIGARCH Computer Architecture News. ACM Press, New York, NY 10036, USA, 1990. CODEN CANED2. ISSN 0163-5964 (ACM), 0884-7495 (IEEE).

Siegel:1996:PIW

[Sie96] Howard Jay Siegel, editor. Proceedings of the 1996 ICPP Workshop on Challenges for Parallel Processing, August 12, 1996. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1996. ISBN 0-8186-7623-X. LCCN QA76.58 .I29 1996; QA76.58; QA76.58 .I327 1996; QA76.58 .I34 1996; QA76.58 .I34 1996eb.

Sincovec:1993:PSS

[SKR^{+}93] Richard F. Sincovec, David E. Keyes, L. M. R., L. R. Petzold, and D. A. Reed, editors. Proceedings of the Sixth SIAM Conference on Parallel Processing for Scientific Computing, held March 22–24, 1993, in Norfolk, VA, USA. SIAM Press, Philadelphia, PA, USA, 1993. ISBN 0-89871-315-3. LCCN QA76.58 .S55 1993 v.1-2. Two volumes.

Stechkin:1977:AFO

[Ste77] S. B. Stechkin, editor. Approximation of functions and operators, volume 138 (1975) of Proceedings of the Steklov Institute of Mathematics. American Mathematical Society, Providence, RI, USA, 1977. ISBN 0-8218-3038-4. iii + 211 pp. LCCN QA1 .A413 no. 138; QA221; QA3 .A462; QA221 .P7413 1977.

Stone:1986:FJC

[SW86] H. S. Stone and S. Winkler, editors. Fall Joint Computer Conference Proceedings: November 2–6, 1986, INFOMART, Dallas, Texas. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1986. ISBN 0-8186-4743-4 (microfiche), 0-8186-8743-6 (hardcover), 0-8186-0743-2 (paperback). LCCN QA75.5 .F35 1986. IEEE catalog number 86CH2345-7. Computer Society order number 743. ACM order number 401860.

Schnebly:1989:Ca

[SWM^{+}89] Dexter A. Schnebly, Bruce W. Walker, Richard R. Moore, John R. Rice, J. H. Matheny, and George Shaw. Correspondence. ACM Fortran Forum, 8(1):21–32, January 1989. CODEN ???? ISSN 1061-7264 (print), 1931-1311 (electronic).

Sydow:1997:IWC

[Syd97] Achim Sydow, editor. 15th IMACS World Congress on Scientific Computation, Modelling and Applied Mathematics: Berlin, August 1997: proceedings, IMACS World Congress. Wissenschaft and Technik, Berlin, Germany, 1997. ISBN 3-89685-550-6 (set), 3-89685-551-4 (vol. 1), 3-89685-552-2 (vol. 2), 3-89685-553-0 (vol. 3),3-89685-554-9 (vol. 4), 3-89685-555-7 (vol. 5), 3-89685-556-5 (vol. 6). LCCN Q183.9 .I46 1997. In cooperation with R.-P. Schafer, W. Rufeger, and H. Lehmann.

Talbot:1970:ATP

[Tal70] Alan Talbot, editor. Approximation theory: proceedings of a symposium held at Lancaster, July 1969. Academic Press, New York, NY, USA, 1970. LCCN QA221 .A66????

Tsompanopoulou:1999:DSAa

[TBMR99a] P. Tsompanopoulou, L. Boloni, D. C. Marinescu, and John R. Rice. The design of software agents for a network of PDE solvers. Technical report TR-99-013, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA,

April 1999. 11 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1999/TR%2099-013.pdf.

Tsompanopoulou:1999:DSAb

[TBMR99b] P. Tsompanopoulou, X. Boloni, D. C. Marinescu, and J. R. Rice. The design of software agents for a network of PDE solvers. In Anonymous [Ano99], page ?? ISBN ???? LCCN ????

Traub:1976:ACC

[Tra76] J. F. (Joseph Frederick) Traub, editor. Analytic computational complexity: Proceedings of the Symposium on Analytic Computational Complexity, held by the Computer Science Department, Carnegie–Mellon University, Pittsburgh, Pennsylvania, on April 7–8, 1975. Academic Press, New York, NY, USA, 1976. ISBN 0-12-697560-4. LCCN QA297 .S915 1975.

Vermeulen:1994:SOO

[Ver94] A. Vermeulen, editor. Second Object-Oriented Numerics Conference. Rogue Wave Software, Corvallis, OR, USA, 1994. ISBN ???? LCCN ????

Verykios:1998:KDM

[VHR98] V. S. Verykios, Elias N. Houstis, and John R. Rice. A knowledge discovery methodology for the performance valuation of scientific software. Technical report TR-98-042, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107,

USA, November 1998. 19 pp. URL http://www.cs.purdue.edu/
research/technical_reports/1998/TR%2098-042.pdf.

Verykos:1999:MPC

[VHR99] V. S. Verykos, Elias N. Houstis, and John R. Rice. Mining the performance of complex systems. In 1999 International Conference on Information Intelligence and Systems, 1999. Proceedings, pages 606–612. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA,

1999. URL http://ieeexplore.ieee.org/stamp/stamp.jsp?
tp=&arnumber=810355.

Verykios:2000:KDM

[VHR00] V. S. Verykios, E. N. Houstis, and J. R. Rice. A knowledge discovery methodology for the performance evaluation of scientific software. Neural, Parallel and Scientific Computations, 8(??):115– 132, ???? 2000. CODEN NPACEM. ISSN 1061-5369. Abridged version in [MWHR00].

Vichnevetsky:1991:IWC

[VM91] Robert Vichnevetsky and John J. H. Miller, editors. IMACS ’91: 13th World Congress on Computation and Applied Mathematics, July 22–26, 1991, Trinity College Dublin, Ireland: proceedings. IMACS, Department of Computer Science, Rutgers University, New Brunswick, NJ, USA, 1991. ISBN ???? LCCN ???? Four volumes.

Vichnevetsky:1979:ACM

[VS79] Robert Vichnevetsky and R. S. Stepleman, editors. Advances in Computer Methods for Partial Differential Equations III: Proceedings of the third IMACS International Symposium on Computer Methods for Partial Differential Equations held at Lehigh University, Bethlehem, Pennsylvania, USA., June 20–22, 1979, volume III. IMACS, Department of Computer Science, Rutgers University, New Brunswick, NJ, USA, 1979. ISBN ???? LCCN ????

Weerawarana:1992:ISN

[WCHR92] S. Weerawarana, Ann C. Catlin, Elias N. Houstis, and John R. Rice. Integrated symbolic-numeric computing in //ELLPACK: Experiences and plans. Technical report TR-92-092, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, March 1992. 18 pp. URL http://www.

cs.purdue.edu/research/technical_reports/1992/TR%2092-
092.pdf.

Wegner:1979:RDS

[Weg79] Peter Wegner, editor. Research directions in software technology, volume 2 of The MIT Press series in computer science. MIT Press, Cambridge, MA, USA, 1979. ISBN 0-262-23096-8. xiii + 869 pp. LCCN QA76.6 .R45.

Weerawarana:1995:ESS

[WHCR95] S. Weerawarana, Elias N. Houstis, Ann C. Catlin, and John R. Rice. //ELLPACK: a system for simulating partial differential equations. Technical report TR-95-043, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA,

June 1995. 6 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1995/TR%2095-043.pdf.

Werrawarana:1990:ISN

[WHR90] S. Werrawarana, Elias N. Houstis, and John R. Rice. An interactive symbolic-numeric interface to parallel ELLPACK for building general PDE solvers. Technical report TR-1054, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, December 1990. 23 pp. URL http://www.cs.purdue.

edu/research/technical_reports/1990/TR%2090-1054.pdf.

Weerawarana:1994:SPI

[WHR94a] S. Weerawarana, Elias N. Houstis, and John R. Rice. A software platform for integrating symbolic computation with a PDE solving environment. Technical report TR-94-031, Department of Computer Science, Purdue University, West Lafayette, IN 47907-

2107, USA, May 1994. 5 pp. URL http://www.cs.purdue.edu/
research/technical_reports/1994/TR%2094-031.pdf.

Weerawarana:1994:POO

[WHR^{+}94b] S. Weerawarana, Elias N. Houstis, John R. Rice, Ann C. Catlin, C. L. Crabill, C. C. Chui, and S. Markus. PDE-Lab: An object-oriented framework for building problem solving environments for PDE based applications. Technical report TR-94-021, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, March 1994. 15

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1994/TR%2094-021.pdf.

Wu:1994:EPP

[WHR94c] P. Wu, E. N. Houstis, and J. R. Rice. EPPOD: a parallel problem solving environment for the electronic prototyping of physical objects design. In Makedon [Mak94], pages 135–151. ISBN ???? LCCN ????

Wu:1994:EPS

[WHR94d] P. Wu, Elias N. Houstis, and John R. Rice. EPPOD: a problem solving environment for parallel electronic prototyping of physical object design. Technical report TR-94-043, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, June 1994. 19 pp. URL http://www.cs.purdue.edu/

research/technical_reports/1994/TR%2094-043.pdf.

Weerawarana:1995:PKB

[WHR^{+}95] S. Weerawarana, E. H. Houstis, John R. Rice, A. Joshi, and C. E. Houstis. PYTHIA: a knowledge based system for intelligent scientific computing. Technical report TR-95-044, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, June 1995. 30 pp. URL http://www.cs.purdue.edu/

research/technical_reports/1995/TR%2095-044.pdf.

Weerawarana:1996:PPK

[WHR^{+}96a] S. Weerawarana, Elias N. Houstis, John R. Rice, Ann C. Catlin, M. G. Gaitatzes, C. L. Crabill, S. Markus, and T. T. Drashansky. The Purdue PSE Kernel: Towards a kernel for building PSEs. Technical report TR-96-082, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, Decem-

ber 1996. 12 pp. URL http://www.cs.purdue.edu/research/
technical_reports/1996/TR%2096-082.pdf.

Weerawarana:1996:WEN

[WHR^{+}96b] S. Weerawarana, Elias N. Houstis, John R. Rice, M. G. Gaitatzes, S. Markus, and A. Joshi. Web//ELLPACK: a networked computing service on the World Wide Web. Technical report TR-96-011, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, February 1996. 19

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1996/TR%2096-011.pdf.

Weerawarana:1996:PKB

[WHR^{+}96c] Sanjiva Weerawarana, Elias N. Houstis, John R. Rice, Anupam Joshi, and Catherine E. Houstis. PYTHIA: a knowledge based system for intelligent scientific computing. ACM Transactions on Mathematical Software, 22(4):447–468, December 1996. CODEN ACMSCU. ISSN 0098-3500 (print), 1557-7295 (elec-

tronic). URL http://www.acm.org/pubs/citations/journals/
toms/1996-22-4/p447-weerawarana/.

Weerawarana:2000:PTK

[WHR^{+}00] S. Weerawarana, E. N. Houstis, J. R. Rice, A. C. Catlin, M. Gaitatzes, C. Crabill, S. Markus, and T. Drashansky. PPK: Towards a kernel for building PSEs. In Houstis et al. [HGRB00], chapter 5, pages 53–63. ISBN 0-7923-7809-1. LCCN QA76 .E548 2000.

Weerawarana:1995:NIN

[WJH^{+}95] S. Weerawarana, A. Joshi, Elias N. Houstis, J. R. Rice, and Ann C. Catlin. Notebook interfaces for networked scientific computing: Design and WWW implementation. Technical report TR-95-048, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, July 1995. 21

pp. URL http://www.cs.purdue.edu/research/technical_
reports/1995/TR%2095-048.pdf.

Weerawarana:1997:NIN

[WJH^{+}97] Sanjiva Weerawarana, Anupam Joshi, Elias N. Houstis, John R. Rice, and Ann C. Catlin. Notebook interfaces for networked scientific computing: design and WWW implementation. Concurrency: Practice and Experience, 9(7):675–695, July 1997. CODEN CPEXEI. ISSN 1040-3108 (print), 1096-

9128 (electronic). URL http://www3.interscience.wiley.com/
cgi-bin/abstract?ID=13884;   http://www3.interscience.
wiley.com/cgi-bin/fulltext?ID=13884&PLACEBO=IE.pdf.

Wright:1989:ACA

[Wri89] M. Wright, editor. Aspects of Computation on Asynchronous Parallel Processors. Proceedings of the IFIP WG 2.5 Working Conference, Stanford, CA, USA, 22–26 August, 1988. North-Holland, Amsterdam, The Netherlands, 1989. ISBN 0-444-87310-4. LCCN QA76.5 .I2775 1988.

Zahar:1994:ACF

[Zah94] R. V. M. (Ramsay Vincent Michael) Zahar, editor. Approximation and computation: a festschrift in honor of Walter Gautschi: proceedings of the Purdue conference, December 2–5, 1993, volume 119 of International series of numerical mathematics. Birkhäuser, Cambridge, MA, USA; Berlin, Germany; Basel, Switzerland, 1994. ISBN 0-8176-3753-2. LCCN QA221 .A634 1994.

Zelkowitz:1998:ACE

[Zel98] Marvin V. Zelkowitz, editor. Advances in Computers: The engineering of large systems, volume 46. Academic Press, New York, NY, USA, 1998. ISBN 0-12-012146-8. xvii + 465 pp. LCCN QA76 .A346 1998eb.

Zelkowitz:2001:AC

[Zel01] Marvin V. Zelkowitz, editor. Advances in Computers, volume 55. Academic Press, New York, NY, USA, 2001. ISBN 0-12-012155-7. xvii + 301 pp. LCCN QA76 .A3 vol. 55.

Zhou:1999:GTS

[ZFHR99] C. Zhou, S. Fleeter, Elias N. Houstis, and J. R. Rice. Gas turbine spray dynamics and combustion simulation design. Technical report TR-99-024, Department of Computer Science, Purdue University, West Lafayette, IN 47907-2107, USA, August 1999.

10 pp. URL http://www.cs.purdue.edu/research/technical_
reports/1999/TR%2099-024.pdf.

Zhang:2004:BA

[ZJ04] David Zhang and Anil K. Jain, editors. Biometric Authentication: First International Conference, ICBA 2004, Hong Kong, China, July 15–17, 2004, Proceedings, volume 3072 of Lecture Notes in Computer Science. Springer-Verlag, Berlin, Germany / Heidelberg, Germany / London, UK / etc., 2004. CO-DEN LNCSD9. ISBN 3-540-22146-8. ISSN 0302-9743 (print), 1611-3349 (electronic). LCCN QA76 A1 L43 3072 (LC). URL

http://link.springer-ny.com/link/service/series/0558/
tocs/t3072.htm;   http://www.springerlink.com/openurl.
asp?genre=issue&issn=0302-9743&volume=3072;   http://
www.springerlink.com/openurl.asp?genre=volume&id=doi:
10.1007/b98225.